### v1

# Fractional binomial distributions induced by the generalized binomial theorem and their applications

Abstract.
A generalized binomial theorem was established by Hara and Hino [Bull. London Math. Soc. 42 (2010), 467–477] for proving the neo-classical inequality. We introduce a new fractional analogue of the binomial distribution on the basis of the generalized binomial theorem and study its properties. They include the law of large numbers, the central limit theorem, and the law of small numbers, the last of which leads to a fractional analogue of the Poisson distribution. As an application of the fractional binomial distribution, a fractional analogue of the Bernstein operator is proposed, which acts on the space of continuous functions on the unit interval. We show that the fractional Bernstein operator uniformly approximates every continuous function on the unit interval. Some limit theorems for its iterates are discussed as well. In particular, a generalization of the Wright–Fisher diffusion process arising in the study of population genetics is captured as a certain limit of the iterates.
Key words and phrases:
fractional binomial distribution, neo-classical inequality, fractional Taylor expansion, fractional Bernstein operator
2020 Mathematics Subject Classification:
Primary 60E05; Secondary 26A33, 60F05, 41A36.
1. Introduction
1.1. The neo-classical inequality and the generalized binomial theorem
Lyons [Lyo98] initiated the theory of rough paths and introduced the neo-classical inequality in order to show a fundamental theorem of rough path theory (see [Lyo98, Theorem 2.2.1]). This asserts that
$\alpha^{2}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}$ (1.1)
for $\alpha\in(0,1]$, $n\in\mathbb{N}$ and $x,y\geq 0$. Here, we define
$\binom{w}{z}:=\frac{\Gamma(w+1)}{\Gamma(z+1)\Gamma(w-z+1)}$
for $w\in\mathbb{C}\setminus\{-k\mid k\in\mathbb{N}\}$ and $z\in\mathbb{C}$ in general, where $\Gamma(\cdot)$ denotes the gamma function. If $\infty$ appears in the denominator of the right-hand side of the above, then $\binom{w}{z}$ is regarded as zero. Also, $0^{0}$ is defined as 1. The equality holds in (1.1) if $\alpha=1$, which is nothing but the usual binomial theorem. Lyons also conjectured that the coefficient $\alpha^{2}$ in (1.1) can be replaced by $\alpha$. Afterwards, Hara and Hino [HH10] gave an affirmative answer to Lyons’ conjecture.
Proposition 1.1 (Neo-classical inequality, [HH10, Theorem 1.2]).
Let $\alpha\in(0,1]$. For $n\in\mathbb{N}$ and $x,y\geq 0$, we have
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}.$ (1.2)
The equality holds if and only if $\alpha=1$ or $x=y=0$.
When $\alpha\in(0,1)$ and only one of $x$ and $y$ is zero, the inequality (1.2) is trivial with strict inequality. Therefore, in order to show (1.2), it suffices to consider the case where $\alpha\in(0,1)$ and $x,y>0$. Moreover, we may assume that $x\leq y$ by the symmetry of the roles of $x$ and $y$ in (1.2). By dividing both sides by $y^{\alpha n}$ and letting $\lambda=x/y$, (1.2) is equivalent to
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}\leq(1+\lambda)^{\alpha n},\qquad\alpha\in(0,1),\,n\in\mathbb{N},\,\lambda\in(0,1].$ (1.3)
For proving (1.3), Hara and Hino established a generalization of the binomial theorem [HH10, Theorem 1.3], which is stated roughly as follows:
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}=(1+\lambda)^{\alpha n}+R_{\alpha}^{(n)}(\lambda),\qquad\alpha\in(0,2),\,n\in\mathbb{N},\,\lambda\in(0,1],$ (1.4)
where the remainder term $R_{\alpha}^{(n)}(\lambda)$ is given explicitly. Indeed, by the expression, the term $R_{\alpha}^{(n)}(\lambda)$ is strictly negative when $\alpha\in(0,1)$, which readily implies (1.3).
Furthermore, they extended the generalized binomial theorem (1.4) to the case where $\alpha$ is an arbitrary positive number. Let us now give the precise statement. In general, for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq 0\}$ and $\gamma\in\mathbb{R}$, the power $z^{\gamma}$ is defined as $\exp(\gamma\,\mathrm{Log}\,z)$, where $\mathrm{Log}\,z$ is the principal value of $\log z$ so that Log $\mathrm{Log}\,1=0$. For $\alpha>0$, we put
$K_{\alpha}:=\{e^{i\theta}\in\mathbb{C}\mid-\pi<\theta\leq\pi,\,e^{i\theta\alpha}=1\}.$ (1.5)
We note that $K_{\alpha}=\{1\}$ when $\alpha\in(0,2)$ and that $-1\in K_{\alpha}$ if and only if $\alpha/2\in\mathbb{N}$. Then, the generalized binomial theorem is stated as follows.
Proposition 1.2 ([HH10, Theorem 3.2]).
For $\alpha>0$, $n\in\mathbb{N}$, and $\lambda\in(0,1]$, we have
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\left\{\frac{1}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{\lambda^{\alpha n}}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.6)
Here, the second term on the right-hand side is regarded as zero if $\alpha/2\in\mathbb{N}$.
It should be noted that the first term on the right-hand side of (1.2) is a real number because $\omega\in K_{\alpha}$ implies $\overline{\omega}\in K_{\alpha}$. For proving Proposition 1.2, the authors provided in [HH10, Theorem 3.1] the fractional-order Taylor-like series expansions with residual terms by using basic complex analysis. See Proposition 2.1 for further details.
1.2. Fractional binomial distribution
Suppose that $n\in\mathbb{N}$ and $x\in[0,1]$. The binomial distribution $\mu_{x}^{(n)}$ defined as
$\mu_{x}^{(n)}(\{j\})=\binom{n}{j}x^{j}(1-x)^{n-j},\qquad j=0,1,2,\dots,n$
on $\{0,1,2,\dots,n\}$ is one of the most important discrete probability distributions. The equality $\sum_{j=0}^{n}\mu_{x}^{(n)}(\{j\})=1$ is derived from the usual binomial theorem. Because Proposition 1.2 is regarded as a generalization of the binomial theorem, it is quite natural to ask for a fractional analogue of the binomial distribution in accordance with Proposition 1.2.
For $\alpha>0$, $n\in\mathbb{N}$ and $x\in(0,1)$, we have
$Z_{\alpha,x}^{(n)}:\!$ $=\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$
$=1+\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}-\frac{\alpha x^{\alpha}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\left\{\frac{(1-x)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{x^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\right\}\,\mathrm{d}t$ (1.7)
by putting $\lambda=x/(1-x)$ in (1.2) and multiplying both sides by $(1-x)^{\alpha n}$. Keeping in mind that the right-hand side of (1.7) is close to 1 when $n$ is sufficiently large, this identity gives rise to a discrete probability distribution on $\{0,1,2,\dots,n\}$ as follows.
Definition 1.3 (Fractional binomial distribution).
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Then, a probability measure $\mu_{\alpha,x}^{(n)}$ on $\{0,1,2,\dots,n\}$, called the $\alpha$-fractional binomial distribution, is defined as
$\mu_{\alpha,x}^{(n)}(\{j\})=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)},\qquad j=0,1,2,\dots,n.$
Moreover, we define $Z_{\alpha,0}^{(n)}=Z_{\alpha,1}^{(n)}=\alpha$, $\mu_{\alpha,0}^{(n)}=\delta_{0}$, and $\mu_{\alpha,1}^{(n)}=\delta_{1}$, where $\delta_{x}$ is the Dirac measure at $x$.
It is clear that $\mu_{1,x}^{(n)}$ coincides with the binomial distribution $\mu_{x}^{(n)}$. The probability distribution $\mu_{\alpha,x}^{(n)}$ is regarded as a fractional variant of $\mu_{x}^{(n)}$. As seen in Corollary 2.8, one of the features is that the mean of $\mu_{\alpha,x}^{(n)}$ is asymptotically equal to that of $\mu_{x}^{(n)}$ as $n\to\infty$, while the variance of $\mu_{\alpha,x}^{(n)}$ is asymptotically $1/\alpha$ times that of $\mu_{x}^{(n)}$.
One of the aims of this study is to study some basic properties of the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, and all such properties are given with their proofs in Section 2. In Section 2.1, we review the fractional-order Taylor-like series of holomorphic functions $f$ with residual terms [HH10, Theorem 2.1]. See Proposition 2.1 for its precise statement. In Section 2.2, we provide explicit expressions for moments of $\mu_{\alpha,x}^{(n)}$ by making effective use of the fractional-order Taylor-like series expansion. See Theorem 2.5 and Corollary 2.8. In particular, we see that the ($c$-)Whitney numbers of the second kind—which in combinatorics are a certain generalization of the Stirling numbers of the second kind—appear in the expression for each moment of $\mu_{\alpha,x}^{(n)}$. Section 2.3 discusses a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to both $n\in\mathbb{N}$ and $x\in[0,1]$. Furthermore, in Section 2.4, a uniform estimate for the moments of $\mu_{\alpha,x}^{(n)}$ is also obtained. Section 2.5 focuses on the characteristic function $\varphi_{\alpha,x}^{(n)}(\xi)$ of $\mu_{\alpha,x}^{(n)}$, which plays a crucial role in obtaining the central limit theorem for $\mu_{\alpha,x}^{(n)}$. Lastly, in Section 2.6, a multivariate version of the neo-classical inequality is deduced. Similarly to Definition 1.3, this inequality also induces a fractional analogue of the multinomial distribution. Being different from the binomial case, however, it seems quite difficult to obtain even basic characteristics of the fractional multinomial distribution such as its mean vector and its variance-covariance matrix, because the multivariate version of the fractional Taylor series expansion is yet to be obtained. The discussion of the multivariate cases is left to the future.
Here, note that the terminology fractional in probability theory is used in several different contexts. For instance, a typical use can be seen in the notion of fractional Brownian motion. See [Lee21] (for example) for the fractional binomial distribution in this sense. Yet another use is derived from fractional calculus. We refer to [CP12] for the fractional binomial process defined in terms of Caputo’s fractional derivatives. However, to the best of our knowledge, no paper to date has discussed the fractional binomial distribution in our sense, which means that the present paper gives a new insight into the study of probability theory and other areas of mathematics.
Based on the properties given in Section 2, we give analogues of three classical limit theorems in Section 3. Sections 3.1 and 3.2 are devoted to the proofs of the weak law of large numbers and the central limit theorem for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$. We also obtain the law of small numbers for $\mu_{\alpha,x}^{(n)}$ in Section 3.3. Namely, we show that $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to a certain probability distribution with masses on $\{0,1,2,\dots\}$ as $n\to\infty$ together with $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ for some $\alpha>0$ and $\lambda>0$. The limiting probability distribution $\nu_{\alpha,\lambda}$ is regarded as a fractional analogue of the Poisson distribution. See Definition 3.3 for the precise definition. Obviously, $\mu_{1,\lambda}$ coincides with the usual Poisson distribution. Note that $\nu_{\alpha,\lambda}$ is defined in terms of a special function called the Mittag-Leffler function $E_{\alpha}(z)$, $\alpha>0$, which is a generalization of the exponential function. The distribution $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions discussed in [Her16, CO17]. We also mention that the Mittag-Leffler function has already been used to introduce some probability distributions in different situations. We refer to [Pil90, PJ95] (for example) for the notion of the Mittag-Leffler distribution, although it is far from the Poisson distribution. A few basic properties of the fractional Poisson distribution such as a recurrence formula for its moments are discussed in Section 3.4 as well. Because the Mittag-Leffler function is used frequently in fractional calculus, further investigations of $\nu_{\alpha,\lambda}$ from the point of view of fractional calculus are expected.
1.3. Fractional Bernstein operator
There exist various applications of the binomial distribution to other branches of mathematics. One of them is seen in the study of positive linear operators acting on $C([0,1])$, the Banach space of all continuous functions on $[0,1]$ endowed with the supremum norm $\|\cdot\|_{\infty}$. Bernstein introduced in his short paper [Ber12] the following positive linear operator acting on $C([0,1])$ defined in terms of the binomial distribution.
Definition 1.4 (Bernstein operator, cf. [Ber12]).
For $n\in\mathbb{N}$, a positive linear operator $B_{n}$ acting on $C([0,1])$ is defined by
$B_{n}f(x)=\sum^{n}_{k=0}\binom{n}{k}x^{k}(1-x)^{n-k}f\left(\frac{k}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$ (1.8)
The positive linear operator $B_{n}$ is called the Bernstein operator.
The right-hand side of (1.8) is called the Bernstein polynomial. Bernstein proved in [Ber12] that $B_{n}$ uniformly approximates every continuous function $f\in C([0,1])$.
Proposition 1.5 (cf. [Ber12]).
For any $f\in C([0,1])$, we have
$\lim_{n\to\infty}\|B_{n}f-f\|_{\infty}=0.$
It is clear that Proposition 1.5 gives a constructive proof of the celebrated Weierstrass approximation theorem.
It is also known that the Bernstein operator has a number of properties of interest. We refer to [Bus17] for comprehensive information on the Bernstein operator with extensive references therein, and to [Alt10, AC94] for surveys on approximating properties of positive linear operators acting on some function spaces, including the Bernstein operator. On the other hand, it is natural to consider limiting behaviors of $B_{n}^{k}$, the $k$-times iterations of $B_{n}$, as $k\to\infty$. One such limiting behavior was first investigated in [KR67].
Proposition 1.6 (cf. [KR67, Theorem 1]).
Let $n\in\mathbb{N}$ be fixed. Then, we have
$\lim_{k\to\infty}\max_{x\in[0,1]}|B_{n}^{k}f(x)-\{f(1)x+f(0)(1-x)\}|=0,\qquad f\in C([0,1]).$
Since the publication of [KR67], several authors have tried to make the proof of Proposition 1.6 more sophisticated. We refer to [AI09, Rus04] (for example) for its short proofs. Moreover, Konstantopoulos, Yuan and Zazanis [KYZ18] discussed a limit theorem for iterates of $B_{n}$ with scales different from Proposition 1.6. In particular, they gave a probabilistic proof of the following limit theorem.
Proposition 1.7 (cf. [KYZ18, Theorem 3]).
For $f\in C([0,1])$ and $t\geq 0$, we have
$\lim_{n\to\infty}\|B_{n}^{\lfloor nt\rfloor}f-\mathbb{E}[f(X_{\bullet}(t))]\|_{\infty}=0,$
where $(X_{x}(t))_{t\geq 0}$ is the Wright–Fisher diffusion process which solves
$\mathrm{d}X_{x}(t)=\sqrt{X_{x}(t)(1-X_{x}(t))}\,\mathrm{d}W(t),\qquad X_{x}(0)=x\in[0,1].$
Here, $(W(t))_{t\geq 0}$ is standard one-dimensional Brownian motion.
This theorem asserts that the sequence of discrete semigroups $(B_{n}^{\lfloor nt\rfloor})_{t\geq 0}$, $n\in\mathbb{N}$ converges to the Wright–Fisher diffusion semigroup as $n\to\infty$. In other words, the properly scaled Markov chain $X_{x}^{(n)}(\lfloor nt\rfloor)$ whose one-step transition operator is given by $B_{n}$ converges in law to $X_{x}(t)$ as $n\to\infty$ for each $t\geq 0$. We also refer to [KZ70] for early works on limit theorems for iterates of some positive linear operators. For further details on the Wright–Fisher diffusion, see [EK86, Section 10] (for example).
In Section 4, we introduce a fractional analogue of the Bernstein operator with the binomial distribution replaced by the fractional one defined in Definition 1.3, and we investigate some of its properties including limit theorems for its iterates. The following is the definition of the fractional Bernstein operator.
Definition 1.8 (Fractional Bernstein operator).
Let $\alpha>0$ and $n\in\mathbb{N}$. The $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ is defined by
$B_{\alpha,n}f(x)$ $:=\sum_{j=0}^{n}\mu_{\alpha,x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
$=\sum_{j=0}^{n}\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$
We note that $B_{1,n}$ coincides with the Bernstein operator $B_{n}$.
In Section 4.1, we show in Theorem 4.1 that the fractional Bernstein operator $B_{\alpha,n}$ also uniformly approximates every continuous function in $C([0,1])$ similarly to Proposition 1.5. Moreover, Sections 4.2 and 4.3 discuss limiting behaviors of the iterates $B_{\alpha,n}^{k}f$, $f\in C([0,1])$ as $k\to\infty$. In the former section, we obtain the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$. In the latter section, we concentrate on a fractional analogue of Proposition 1.7. We first show that the sequence $\{n(B_{\alpha,n}-I)\}_{n=1}^{\infty}$ converges to a second-order differential operator $\mathcal{L}_{\alpha}=(2\alpha)^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$ with Neumann boundary condition. This means the convergence of the infinitesimal generator of the discrete Markov semigroup generated by $B_{\alpha,n}$ to that of a diffusion semigroup. Note that $\mathcal{L}_{\alpha}$ is the generator of the Wright–Fisher diffusion with diffusion coefficient different from that of $(X_{x}(t))_{t\geq 0}$ up to a multiple constant. Next, we show the uniform convergence of the iterates $B_{\alpha,n}^{\lfloor nt\rfloor}f$, $f\in C([0,1])$ to the Wright–Fisher semigroup corresponding to $\mathcal{L}_{\alpha}$ as $n\to\infty$, by employing the celebrated Trotter approximation theorem (cf. [Tro58]). See Lemma 4.5 and Theorem 4.4 for precise statements. We also discuss in Section 4.4 that Theorem 4.4 holds even though we replace the parameter $\alpha>0$ with a continuous function $\alpha(x)\colon[0,1]\to(0,\infty)$. In this case, the corresponding infinitesimal generator is given by $\mathcal{L}_{\alpha(\cdot)}=(2\alpha(x))^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$. Hence, it turns out that the stochastic processes captured by the limit of $B_{\alpha(\cdot),n}^{\lfloor nt\rfloor}$ as $n\to\infty$ are not only Wright–Fisher diffusions but also more-general diffusions on $[0,1]$.
Here, we mention some existing notions of linear operators named the fractional Bernstein operator. In [Man15], the author introduced the fractional Bernstein operator $(1-B_{n})^{\alpha}$, $\alpha>0$ in order to apply density estimations of probability distributions in statistics and data science, which differs completely from ours. Another different formulation of the fractional Bernstein operator from ours can be found in [Yu21]. Our fractional Bernstein operator $B_{\alpha,n}$ seems new in that we define it via the generalized binomial theorem deriving the neo-classical inequality. We expect to develop further studies of $B_{\alpha,n}$ in the near future.
1.4. Notation
We use the following notation.
  • We set =$\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
  • $\mathrm{Re}(\omega)$ denotes the real part of a complex number $\omega$.
  • The indicator function of a set $A$ is denoted by $\bm{1}_{A}$.
  • The cardinality of a set $A$ is denoted by #$\#(A)$.
  • For $x,y\in\mathbb{R}$, we write $x\vee y=\max\{x,y\}$ and $x\wedge y=\min\{x,y\}$.
  • For $x\in\mathbb{R}$, we denote by $\lfloor x\rfloor$ and $\lceil x\rceil$ the greatest integer less than or equal to $x$, and the smallest integer greater than or equal to $x$, respectively.
  • For $a\in\mathbb{R}$, we denote by $\mathrm{sgn}\,a$ the sign of $a$, that is,
    $\mathrm{sgn}\,a=\begin{cases}1&\text{if }a>0,\\ 0&\text{if }a=0,\\ -1&\text{if }a<0.\end{cases}$
  • Let $f(n)$ and $g(n)$ be functions defined on $\mathbb{N}$. Then, $f(n)=o(g(n))$ and $f(n)=O(g(n))$ as $n\to\infty$ mean the usual little-$o$ and big-$O$ notations, respectively. $f(n)=\Omega(g(n))$ means $g(n)=O(f(n))$. Moreover, we use $f(n)\sim g(n)$ as $n\to\infty$ to indicate $f(n)/g(n)\to 1$ as $n\to\infty$.
  • Let $X$ and $Y$ be two Banach spaces. The domain and range of a linear operator $L\colon X\to Y$ are denoted by $\mathrm{Dom}(L)\subset X$ and $\mathrm{Range}(L)\subset Y$, respectively.
  • A positive constant whose concrete expression is not important is denoted by $C_{i,j}$ when it is the $j$-th positive constant in Section $i$
2. The fractional binomial distribution and its properties
2.1. A fractional-order Taylor series
Let $D=\{z\in\mathbb{C}\,:\,|z|<1\}$ be the unit open disc and $\overline{D}$ its closure in $\mathbb{C}$. Consider a continuous function $f$ on $\overline{D}$ such that $f$ is holomorphic in $D$. For each $\xi\in\mathbb{R}$, we define
$f^{\#}(\xi):=\int_{-1/2}^{1/2}f(e^{2\pi ix})e^{-2\pi ix\xi}\,\mathrm{d}x=\frac{1}{2\pi i}\int_{C}\frac{f(z)}{z^{\xi+1}}\,\mathrm{d}z,$
where $C$ denotes the oriented contour $(-1,1)\ni t\longmapsto e^{i\pi t}\in\mathbb{C}$.
The following is the fractional-order Taylor series with residual terms. Recall that the set $K_{\alpha}$ was defined in (1.5).
Proposition 2.1 (cf. [HH10, Theorem 3.1]).
Suppose that $\alpha>0$ and $\lambda\in(0,1)$. If $\alpha/2\notin\mathbb{N}$, then we have
$\alpha\sum_{j=0}^{\infty}f^{\#}(\alpha j)\lambda^{\alpha j}$ $=\sum_{\omega\in K_{\alpha}}f(\lambda\omega)-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}f(-t)}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t,$ (2.1)
$\alpha\sum_{j=-\infty}^{-1}f^{\#}(\alpha j)\lambda^{-\alpha j}$ $=\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}f(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.2)
If $\alpha/2\in\mathbb{N}$, (2.1) and (2.2) still hold by regarding the second term of the right-hand side of (2.1) and the right-hand side of (2.2) as 00.
We note that these two equalities in Proposition 2.1 turn out to be true when $\lambda=1$ through the limit as $\lambda\nearrow 1$, provided that
$\sum_{j=-\infty}^{\infty}|f^{\#}(\alpha j)|<\infty.$
Since it holds that
$f^{\#}(\xi)=\begin{cases}\displaystyle\frac{1}{\xi!}\frac{\mathrm{d}^{\xi}f}{\mathrm{d}z^{\xi}}(0)&\text{if }\xi\geq 0,\\ 0&\text{if }\xi<0,\end{cases}\qquad\xi\in\mathbb{Z},$
the identity (2.1) is identical to the usual Taylor series expansion of $f$ at $z=0$ when $\alpha=1$. On the other hand, if $\xi\notin\mathbb{Z}$, we see that $\Gamma(\xi+1)f^{\#}(\xi)$ is regarded as the $\xi$-order fractional derivative of $f$ at $z=0$. We also recall the following result.
Proposition 2.2 (cf. [HH10, Proposition 2.4]).
Let $T>0$ and define $f(z)=(1+z)^{T}$ on $\overline{D}$. Then, we have
$f^{\#}(\xi)=\binom{T}{\xi},\qquad\xi\in\mathbb{R}.$
2.2. Moments of the fractional binomial distribution
Suppose that $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. We aim to find concrete expressions of moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ by applying the fractional Taylor series (2.1) and (2.2) to a certain holomorphic function. Let $c\in\mathbb{R}$ and define a differential operator $\mathscr{D}_{c}$ acting on holomorphic functions $f$ on a domain of $\mathbb{C}$ by
$\mathscr{D}_{c}f(z):=zf^{\prime}(z)+cf(z).$
We also define
$h_{c}^{(m)}(z):=\mathscr{D}_{c}^{m}\{(1+z)^{\alpha n}\},\qquad z\in\begin{cases}\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq-1\}&\text{if }\alpha n\notin\mathbb{N},\\ \mathbb{C}&\text{if }\alpha n\in\mathbb{N}\end{cases}$ (2.3)
for $m\in\mathbb{N}_{0}$. Moreover, we put
$(s)_{0}=1,\qquad(s)_{k}=s(s-1)\cdots(s-k+1),\qquad k\in\mathbb{N}$
for $s>0$. Then, we have the following explicit representation of $h_{c}^{(m)}$.
Lemma 2.3.
For $c\in\mathbb{R}$, we define a sequence $\{W_{c}(m,k)\}_{0\leq k\leq m}$, $m\in\mathbb{N}_{0}$, by
$W_{c}(m,0)$ $=c^{m},$ (2.4)
$W_{c}(m,k)$ $=(k+c)W_{c}(m-1,k)+W_{c}(m-1,k-1),\qquad k=1,2,\dots,m-1,$
$W_{c}(m,m)$ $=1.$
Then, it holds that
$h_{c}^{(m)}(z)=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$ (2.5)
for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq-1\}$ if $\alpha n\notin\mathbb{N}$ and $z\in\mathbb{C}$ if $\alpha n\in\mathbb{N}$. Moreover, each $W_{c}(m,k)$ is a polynomial in $c$ of degree $m-k$ with integer coefficients (see also Table 1).
$m$ $k$ 0 1 2 3 4
0 1
1 $c$ 1
2 $c^{2}$ $12c+$ 1 1
3 $c^{3}$ $13c^{2}+3c+$ 1 $33c+$ 3 1
4 $c^{4}$ $14c^{3}+6c^{2}+4c+$ 1 $76c^{2}+12c+$ 7 $64c+$ 6 1
Table 1. Values of $W_{c}(m,k)$ for $m=0,1,2,3,4$.
Proof.
The proof is by induction. When $m=0$, we have $h_{c}^{(0)}(z)=(1+z)^{\alpha n}$ and (2.5) holds. Suppose that (2.5) holds for some $m\in\mathbb{N}_{0}$. Then, we have
$h_{c}^{(m+1)}(z)$ $=\mathscr{D}_{c}h_{c}^{(m)}(z)$
$=z(h_{c}^{(m)})^{\prime}(z)+ch_{c}^{(m)}(z)$
$=z\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\{kz^{k-1}(1+z)^{\alpha n-k}+(\alpha n-k)z^{k}(1+z)^{\alpha n-k-1}\}$
$\hskip 28.45274pt+c\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
$=\sum_{k=1}^{m}\{(k+c)W_{c}(m,k)+W_{c}(m,k-1)\}(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
$\hskip 28.45274pt+W_{c}(m,m)(\alpha n)_{m+1}z^{m+1}(1+z)^{\alpha n-m-1}+cW_{c}(m,0)(1+z)^{\alpha n}.$
Therefore, it follows from (2.4) that the right-hand side coincides with
$\sum_{k=0}^{m+1}W_{c}(m+1,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k},$
which implies that (2.5) is also true for the case of $m+1$. The latter assertion can also be shown immediately by induction, noting the recurrence formula (2.4). ∎
Remark 2.4.
Let $c=0$. Then, the recurrence formula (2.4) gives the Stirling number of the second kind $W_{0}(m,k)=S(m,k)$, which is the number of ways to partition a set of $m$ objects into $k$ non-empty subsets. Therefore, the sequence $\{W_{c}(m,k)\}$ is regarded as an extension of the Stirling numbers of the second kind, which in combinatorics are referred to as the ($c$-)Whitney numbers of the second kind. To the best of our knowledge, this notion was first introduced in [Mez10], and indeed the theorem below tells us that $W_{c}(m,k)$ appears when we express the moments of the fractional binomial distribution.
By applying Lemma 2.3, we obtain the following key identity to deduce the moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ in Definition 1.3.
Theorem 2.5.
Let $\alpha>0$, $n\in\mathbb{N}$, $x\in(0,1)$, $c\in\mathbb{R}$, and $m\in\mathbb{N}_{0}$. If $\alpha n\notin\mathbb{N}$, suppose further that $m\leq\alpha n$. Then, we have
$Z_{\alpha,x}^{(n)}\sum_{j=0}^{n}(\alpha j+c)^{m}\mu_{\alpha,x}^{(n)}(\{j\})$
$=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}x^{k}+\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\left\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right\}$
$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.6)
Here, if $\alpha\in\mathbb{N}$, then the third and fourth terms of the right-hand side are regarded as 00.
Proof.
The proof is split into two parts.
Step 1. Let $f$ be a continuous function on $\overline{D}$ such that $f$ is holomorphic on $D$ and $f(-1)=0$. Suppose that a function $g$ defined by
$g(z):=\mathscr{D}_{c}f(z)=zf^{\prime}(z)+cf(z),\qquad z\in D$
is extended as a continuous function on $\overline{D}$ and the function $(-1/2,1/2)\ni x\longmapsto g(e^{2\pi ix})$ is of class $C^{1}$. Then, we have
$g^{\#}(\xi)$ $=\int_{-1/2}^{1/2}e^{2\pi ix}f^{\prime}(e^{2\pi ix})e^{-2\pi ix\xi}\,\mathrm{d}x+cf^{\#}(\xi)$
$=\lim_{\varepsilon\searrow 0}\frac{1}{2\pi i}\int_{-1/2+\varepsilon}^{1/2-\varepsilon}\left(\frac{\mathrm{d}}{\mathrm{d}x}f(e^{2\pi ix})\right)e^{-2\pi ix\xi}\,\mathrm{d}x+cf^{\#}(\xi)$
$=\lim_{\varepsilon\searrow 0}\left(\left[\frac{1}{2\pi i}f(e^{2\pi ix})e^{-2\pi ix\xi}\right]_{x=-1/2+\varepsilon}^{x=1/2-\varepsilon}+\int_{-1/2+\varepsilon}^{1/2-\varepsilon}f(e^{2\pi ix})\xi e^{-2\pi ix\xi}\,\mathrm{d}x\right)+cf^{\#}(\xi)$
$=(\xi+c)f^{\#}(\xi),\qquad\xi\in\mathbb{R}.$ (2.7)
In the last equality, we used $f(-1)=0$. Because $m\leq\alpha n$ or $\alpha n\in\mathbb{N}$ holds, we can take the function $h_{c}^{(k)}$, $k=0,1,\dots,m-1$ defined by (2.3) as $f$. Therefore, we obtain
$(h_{c}^{(m)})^{\#}(\xi)=(\xi+c)(h_{c}^{(m-1)})^{\#}(\xi)=\cdots=(\xi+c)^{m}(h_{c}^{(0)})^{\#}(\xi)=(\xi+c)^{m}\binom{\alpha n}{\xi}$ (2.8)
for $\xi\in\mathbb{R}$, from (2.7) and Proposition 2.2.
Step 2. We apply Proposition 2.1 to the function $f(z)=h_{c}^{(m)}(z)$. Then, it follows from (2.8) that
$\alpha\sum_{j=0}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}(\lambda\omega)-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t,$ (2.9)
$\alpha\sum_{j=-\infty}^{-1}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{-\alpha j}$ $=\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.10)
Here, if $\alpha\in\mathbb{N}$, then the second term of the right-hand side of (2.9) and the right-hand side of (2.10) are regarded as 0. Because it holds that
$\alpha\sum_{j=n+1}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\alpha\sum_{l=-\infty}^{-1}\{\alpha(n-l)+c\}^{m}\binom{\alpha n}{\alpha(n-l)}\lambda^{\alpha(n-l)}$
$=(-1)^{m}\lambda^{\alpha n}\alpha\sum_{l=-\infty}^{-1}\{\alpha l-(c+\alpha n)\}^{m}\binom{\alpha n}{\alpha l}\lambda^{-\alpha l}$
$=(-1)^{m}\lambda^{\alpha n}\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t$
by using (2.10) with $c$ replaced by $-(c+\alpha n)$, we apply it to (2.9) to obtain
$\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\alpha\sum_{j=0}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}-\alpha\sum_{j=n+1}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$
$=\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}(\lambda\omega)-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha\lambda^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.11)
We put $\lambda=x/(1-x)$, $20<x<1/$ 2 and multiply both sides in (2.11) by $(1-x)^{\alpha n}$. Then, we obtain
$\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$
$=(1-x)^{\alpha n}\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}\left(\frac{x\omega}{1-x}\right)$
$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.12)
Because
$|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}=\{(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i\alpha\pi}\}$
and
$|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}=\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}\}$
are extended to holomorphic functions in $x$ on a complex domain containing the set $\{x\in\mathbb{R}\mid 0<x<1\}$ and do not vanish if $\alpha\notin\mathbb{N}$, (2.12) also holds for $10<x<$ 1 by analytic continuation. Moreover, by applying (2.5) with $z=x\omega/(1-x)$ and $\omega\in K_{\alpha}$ to the first term of the right-hand side of (2.12), we have
$(1-x)^{\alpha n}\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}\left(\frac{x\omega}{1-x}\right)$
$=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\left\{\sum_{\omega\in K_{\alpha}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right\}$
$=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}x^{k}+\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\left\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right\}.$
We remark that if $\omega=-1\in K_{\alpha}$, then $\alpha/2\in\mathbb{N}$ so that $h_{c}^{(m)}(x\omega/(1-x))$ is well-defined. Thus, (2.12) implies (2.6). ∎
As seen from (1.7), the normalized constant $Z_{\alpha,x}^{(n)}$ satisfies
$Z_{\alpha,x}^{(n)}=1+O(e^{-\delta n}),\qquad\frac{1}{Z_{\alpha,x}^{(n)}}=1+O(e^{-\delta n})$ (2.13)
as $n\to\infty$ for some $\delta>0$, where $\delta$ can be chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\alpha\in(0,\infty)$. In what follows, we use the notation $O(e^{-\delta n})$ in this sense. Then, we also have the following.
Proposition 2.6.
For each $m\in\mathbb{N}_{0}$, there exists some $\delta>0$ such that
$\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}x^{k}+(|c|^{m}+1)O(e^{-\delta n})$ (2.14)
as $n\to\infty$.
In order to prove Proposition 2.6, we need to give the following estimates of the third term on the right-hand side of (1.7).
Lemma 2.7.
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Then, we have
$\Bigg{|}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}(1-t)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t\Bigg{|}$ $\leq(1-x)^{\alpha n}\psi(\alpha)$ (2.15)
and
$\Bigg{|}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}(1-t)^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg{|}$ $\leq x^{\alpha n}\psi(\alpha),$ (2.16)
where $\psi(\alpha)$ is the periodic function of period $2$ such that $\psi(\alpha)=1-|\alpha|$ for $|\alpha|\leq 1$.
Proof.
Because the left-hand sides of (2.15) and (2.16) are regarded as 00 if $\alpha\in\mathbb{N}$, we may suppose that $\alpha$ is not an integer. It is sufficient to show (2.15) because (2.16) is immediately obtained by replacing $x$ in (2.15) with $1-x$. We have
$\Bigg{|}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}(1-t)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t\Bigg{|}$
$\leq\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}}{\pi}|\sin\alpha\pi|\int_{0}^{1}\frac{t^{\alpha-1}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$ (2.17)
By the change of variable $s=((1-x)t/x)^{\alpha}$ and [GR15, Equation 12, page 329], the right-hand side of (2.17) becomes
$\frac{(1-x)^{\alpha n}}{\pi}|\sin\alpha\pi|\int_{0}^{(\frac{1-x}{x})^{\alpha}}\frac{\mathrm{d}s}{|s-e^{-i\alpha\pi}|^{2}}$
$\leq\frac{(1-x)^{\alpha n}}{\pi}|\sin\alpha\pi|\int_{0}^{\infty}\frac{\mathrm{d}s}{(s-\cos\alpha\pi)^{2}+\sin^{2}\alpha\pi}$
$=\frac{(1-x)^{\alpha n}}{\pi}|\sin\alpha\pi|\times\frac{\pi\psi(\alpha)}{|\sin\alpha\pi|}=(1-x)^{\alpha n}\psi(\alpha),$
which is the desired estimate. ∎
We are now in a position to prove Proposition 2.6.
Proof of Proposition 2.6.
For $\omega\in K_{\alpha}\setminus\{1\}$, that is, $\omega=e^{i\theta}$ for some $\theta\in(-\pi,\pi]\setminus\{0\}$, we observe that
$|1-x+x\omega|^{2}$ $=(1-x)^{2}+x^{2}+2x(1-x)\mathrm{Re}(\omega)$
$=1-2x(1-x)(1-\mathrm{Re}(\omega))<1,$ (2.18)
which implies that for each $k\in\mathbb{N}_{0}$,
$\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}=O(e^{-\delta n})$ (2.19)
as $n\to\infty$ for some $\delta>0$. Furthermore, Lemma 2.3 implies that
$|W_{c}(m,k)|\leq C_{2,1}(|c|^{m-k}+1),\qquad m\in\mathbb{N}_{0},\,\,k=0,1,\dots,m$ (2.20)
for some $C_{2,1}>0$ depending only on $m\in\mathbb{N}_{0}$. Then, for large $n$ such that $\alpha n\geq m$, we have
$|h_{c}^{(m)}(-t)|$ $\leq\sum_{k=0}^{m}C_{2,2}(|c|^{m-k}+1)n^{k},$ (2.21)
$|h_{-(c+\alpha n)}^{(m)}(-t)|$ $\leq\sum_{k=0}^{m}C_{2,3}(|c+\alpha n|^{m-k}+1)n^{k}$ (2.22)
for $t\in(0,1)$. Hence, the second term on the right-hand side of (2.6) is estimated as
$\left|\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\left\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right\}\right|$
$\leq\sum_{k=0}^{m}C_{2,4}(|c|^{m-k}+1)n^{k}\left|\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right|=(|c|^{m}+1)O(e^{-\delta n})$
as $n\to\infty$ for some $\delta>0$, by using (2.19) and (2.20). Then, (2.21), (2.22) and Lemma 2.7 imply that the sum of the third and fourth terms on the right-hand side of (2.6) is dominated by
$(1-x)^{\alpha n}\psi(\alpha)\sum_{k=0}^{m}C_{2,2}(|c|^{m-k}+1)n^{k}+x^{\alpha n}\psi(\alpha)\sum_{k=0}^{m}C_{2,3}(|c+\alpha n|^{m-k}+1)n^{k},$
which is $(|c|^{m}+1)O(e^{-\delta n})$ as well. Note that these items are regarded as 0 if $\alpha\in\mathbb{N}$. ∎
Note that Proposition 2.6 with $c=-\alpha nx$ is often used in the rest of this paper. When $c=-\alpha nx$, the second term in the right-hand side of (2.14) is $O(e^{-\delta n})$ with possibly different $\delta>0$. A direct calculation also gives
$\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}=\begin{cases}0&\text{if }m=1,\\ \alpha nx(1-x)&\text{if }m=2,\\ \alpha nx(1-x)(1-2x)&\text{if }m=3,\\ \alpha nx(1-x)\{3(\alpha n-2)x(1-x)+1\}&\text{if }m=4.\end{cases}$ (2.23)
The mean and variance of $\mu_{\alpha,x}^{(n)}$ are obtained as follows.
Corollary 2.8.
Let $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is given by $\mu_{\alpha,x}^{(n)}$. Then, for some $\delta>0$, we have
$\mathbb{E}[S_{\alpha,x}^{(n)}]=\sum_{j=0}^{n}j\mu_{\alpha,x}^{(n)}(\{j\})=nx+O(e^{-\delta n})$ (2.24)
and
$\mathrm{Var}(S_{\alpha,x}^{(n)})=\sum_{j=0}^{n}\left(j-\mathbb{E}[S_{\alpha,x}^{(n)}]\right)^{2}\mu_{\alpha,x}^{(n)}(\{j\})=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$
as $n\to\infty$.
Proof.
Applying Proposition 2.6 with $c=0$ and $m=1$, we have
$\alpha\sum_{j=0}^{n}(\alpha j)\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$ $=W_{0}(1,0)+W_{0}(1,1)\alpha nx+O(e^{-\delta n})$
$=\alpha nx+O(e^{-\delta n})$
as $n\to\infty$, which implies (2.24). We also apply (2.13), (2.23), and Proposition 2.6 with $c=-\alpha nx$ and $m=2$ to obtain
$\sum_{j=0}^{n}(j-nx)^{2}\mu_{\alpha,x}^{(n)}(\{j\})=\frac{1}{Z_{\alpha,x}^{(n)}}\left\{\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})\right\}=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$
as $n\to\infty$. Hence, (2.24) leads to
$\mathrm{Var}(S_{\alpha,x}^{(n)})$ $=\sum_{j=0}^{n}(j-nx)^{2}\mu_{\alpha,x}^{(n)}(\{j\})-\left(nx-\mathbb{E}[S_{\alpha,x}^{(n)}]\right)^{2}$
$=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$
as $n\to\infty$. ∎
2.3. Uniform bound of $Z_{\alpha,x}^{(n)}$
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. We will give a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to $x$ and $n$, and again we note that $Z_{1,x}^{(n)}=1$.
Theorem 2.9 (Uniform bound of $Z_{\alpha,x}^{(n)}$).
(1) Let $\alpha\in(0,2)$. Then, we have
$\alpha\wedge 1\leq Z_{\alpha,x}^{(n)}\leq\alpha\vee 1$ (2.25)
for $x\in[0,1]$ and $n\in\mathbb{N}$.
(2) Let $\alpha\in[2,\infty)$. Then, we have
$\alpha^{1-\alpha^{2}}\leq Z_{\alpha,x}^{(n)}\leq\alpha$ (2.26)
for $x\in[0,1]$ and $n\in\mathbb{N}$.
Proof.
If $\alpha\in(0,1)$, then Proposition 1.1 implies that $Z_{\alpha,x}^{(n)}\leq 1$ for $x\in[0,1]$ and $n\in\mathbb{N}$. On the other hand, Equation (1.7) in [GS23, Theorem 1.3] leads to $Z_{\alpha,x}^{(n)}\leq\alpha$ for $x\in[0,1]$ and $n\in\mathbb{N}$ when $\alpha\geq 1$. Hence, the uniform upper bounds of (2.25) and (2.26) have been obtained.
We next discuss the lower bounds of $Z_{\alpha,x}^{(n)}$. Because it holds that $Z_{\alpha,x}^{(n)}=Z_{\alpha,1-x}^{(n)}$ and $Z_{\alpha,0}^{(n)}=\alpha$, it suffices to consider only the case where $x\in(0,1/2]$. It follows from [HH10, Proposition 2.6] that
$R_{\alpha}^{(n)}(\lambda)\in\begin{cases}(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ (1-\alpha,0)&\text{if }\alpha\in(1,2)\end{cases}$
for $\lambda\in(0,1]$, where $R_{\alpha}^{(n)}(\lambda)$ is defined by
$R_{\alpha}^{(n)}(\lambda)=-\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}+(1+\lambda)^{\alpha n}.$ (2.27)
Letting $\lambda=x/(1-x)$ and multiplying both sides of (2.27) by $(1-x)^{\alpha n}$, we have
$-\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}+1\in\begin{cases}\left(0,(1-\alpha)(1-x)^{\alpha n}\right)\subset(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ \left((1-\alpha)(1-x)^{\alpha n},0\right)&\text{if }\alpha\in(1,2).\end{cases}$
This implies that $Z_{\alpha,x}^{(n)}>\alpha$ if $\alpha\in(0,1)$ and $Z_{\alpha,x}^{(n)}>1$ if $\alpha\in(1,2)$. Hence, all that remains is the case $\alpha\geq 2$. We have
$Z_{\alpha,x}^{(n)}$ $\geq\alpha\binom{\alpha n}{0}(1-x)^{\alpha n}$
$=\alpha(1-x)^{\frac{1}{x}\times\alpha nx}$
$\geq\alpha\left(1-\frac{1}{2}\right)^{2\alpha nx}=\alpha\times 2^{-2\alpha nx},\qquad x\in(0,1/2],\,n\in\mathbb{N}$ (2.28)
by noting the fact that the function $(0,1/2]\ni x\longmapsto(1-x)^{1/x}\in\mathbb{R}$ is monotonically decreasing. Let $\gamma>0$ be a constant that is specified later. Then, (2.28) implies that
$Z_{\alpha,x}^{(n)}\geq\alpha\times 2^{-2\gamma},\qquad x\in(0,1/2],\,n\in\mathbb{N}\,\,\text{with $\alpha nx\leq\gamma$.}$
We now deal with the upper bound of $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$ for $x\in(0,1/2]$ and $n\in\mathbb{N}$ with $\alpha nx>\gamma$. From (2.18), we have
$\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$
$\leq\sum_{\omega\in K_{\alpha}\setminus\{1\}}\{1-2x(1-x)(1-\mathrm{Re}(\omega))\}^{\alpha n/2}$
$\leq\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left[\exp\{-2x(1-x)(1-\mathrm{Re}(\omega))\}\right]^{\alpha n/2}$
$\leq\alpha\exp\left\{-\alpha nx(1-x)\left(1-\max_{\omega\in K_{\alpha}\setminus\{1\}}\mathrm{Re}(\omega)\right)\right\}$
$\leq\alpha\exp\left(-\alpha nx\times\frac{1}{2}\times\frac{8}{\alpha^{2}}\right)=\alpha\exp\left(-\alpha nx\times\frac{4}{\alpha^{2}}\right)$ (2.29)
for $x\in(0,1/2]$ and $n\in\mathbb{N}$, where we used #$\#(K_{\alpha})=\lfloor\alpha/2\rfloor+\lceil\alpha/2\rceil\leq\alpha+1$ in the fourth line and
$\max_{\omega\in K_{\alpha}\setminus\{1\}}\mathrm{Re}(\omega)=\cos\frac{2\pi}{\alpha}\leq 1-\frac{8}{\alpha^{2}}$
in the last line. Here, the estimate $\cos(2\pi/\alpha)\leq 1-8/\alpha^{2}$ follows from the inequality
$\cos t\leq 1-\frac{2t^{2}}{\pi^{2}},\qquad t\in[0,\pi],$
which is straightforward by elementary calculus. Therefore, we obtain
$Z_{\alpha,x}^{(n)}$ $\geq 1-\alpha\exp\left(-\alpha nx\times\frac{4}{\alpha^{2}}\right)-\{(1-x)^{\alpha n}+x^{\alpha n}\}\psi(\alpha)$
$\geq 1-\alpha\exp\left(-\alpha nx\times\frac{4}{\alpha^{2}}\right)-2e^{-\alpha nx}$
by applying (2.29) and Lemma 2.7 to (1.7). In particular, we see that
$Z_{\alpha,x}^{(n)}\geq 1-\alpha\exp\left(-\frac{4\gamma}{\alpha^{2}}\right)-2e^{-\gamma}$
when $\alpha nx>\gamma$. Bringing everything together, it holds that
$Z_{\alpha,x}^{(n)}\geq\begin{cases}\alpha\times 2^{-2\gamma}&\text{if }\alpha nx\leq\gamma,\\ \displaystyle 1-\alpha\exp\left(-\frac{4\gamma}{\alpha^{2}}\right)-2e^{-\gamma}&\text{if }\alpha nx>\gamma.\end{cases}$
Letting $\gamma=(\alpha^{2}\log\alpha)/(2\log 2)$, we have
$\alpha\times 2^{-2\gamma}=\alpha\times 2^{-\alpha^{2}\log_{2}\alpha}=\alpha^{1-\alpha^{2}}\leq 2^{1-2^{2}}=0.125$
and
$1-\alpha\exp\left(-\frac{4\gamma}{\alpha^{2}}\right)-2e^{-\gamma}$ $=1-\alpha e^{-2\log_{2}\alpha}-2\alpha^{-\alpha^{2}/(2\log 2)}$
$=1-\alpha^{1-2/\log 2}-2\alpha^{-\alpha^{2}/(2\log 2)}$
$\geq 1-2^{1-2/\log 2}-2\times 2^{-2^{2}/(2\log 2)}$
$=1-4e^{-2}=0.458\cdots.$
Therefore, we obtain
$Z_{\alpha,x}^{(n)}\geq\min\left\{\alpha\times 2^{-2\gamma},1-\alpha\exp\left(-\frac{4\gamma}{\alpha^{2}}\right)-2e^{-\gamma}\right\}\Bigg{|}_{\gamma=(\alpha^{2}\log\alpha)/(2\log 2)}=\alpha^{1-\alpha^{2}}$
for $x\in(0,1/2]$ and $n\in\mathbb{N}$. ∎
2.4. Uniform estimates for moments of the fractional binomial distribution
In Theorem 2.5, we discussed the explicit representations of moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$. Here, we discuss their uniform estimates with respect to $n\in\mathbb{N}$ and $x\in[0,1]$. In order to deduce them, the following lemma is useful.
Lemma 2.10.
Let $p>0$ and $q>0$. Then, we have
$\max_{t\in[0,1]}t^{p}(1-t)^{q}\leq\left(\frac{p}{eq}\right)^{p}.$
Proof.
Let $g(t)=t^{p}(1-t)^{q}$ for $t\in[0,1]$. Then, we have
$g^{\prime}(t)=t^{p-1}(1-t)^{q-1}\{p(1-t)-qt\}.$
Therefore, we see that the function $g$ attains its maximum at $t=p/(p+q)$, which is
$g\left(\frac{p}{p+q}\right)=\left(\frac{p}{q}\right)^{p}\left\{\left(1-\frac{p}{p+q}\right)^{\frac{p+q}{p}}\right\}^{p}\leq\left(\frac{p}{eq}\right)^{p}.\qed$
The purpose of this section is to show the following.
Theorem 2.11.
Let $m\in\mathbb{N}$ and $K$ be a compact subset of $(0,\infty)$. Then, we have
$\sup_{\alpha\in K}\max_{x\in[0,1]}\Bigg{|}\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}$
$\hskip 28.45274pt-(\alpha n)^{-m}\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}\Bigg{|}=O\left(\frac{1}{n^{m}}\right)$ (2.30)
as $n\to\infty$.
Proof.
Let us set
$M_{\alpha,m}^{(n)}(x):=\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}.$
By continuity, it suffices to estimate the supremum over $x\in(0,1)$. The proof is split into three parts.
Step 1. By applying Theorem 2.5 with $c=-\alpha nx$, we have
$(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)$
$=\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}$
$\hskip 28.45274pt+\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}\Bigg{\{}\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\Bigg{\}}$
$\hskip 28.45274pt-\Bigg{\{}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt+(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg{\}}$
$=:\mathrm{I}_{\alpha,m}^{(n)}(x)+\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x).$ (2.31)
We note here that
$M_{\alpha,m}^{(n)}(1-x)$ $=(-1)^{m}\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}(1-x)^{\alpha j}x^{\alpha(n-j)}\left(\frac{n-j}{n}-x\right)^{m}$
$=(-1)^{m}\alpha\sum_{k=0}^{n}\binom{\alpha n}{\alpha n-\alpha k}(1-x)^{\alpha(n-k)}x^{\alpha k}\left(\frac{k}{n}-x\right)^{m}$
$=(-1)^{m}M_{\alpha,m}^{(n)}(x)$
and
$\mathrm{III}_{\alpha,m}^{(n)}(1-x)$ $=\frac{\alpha(1-x)^{\alpha}x^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(tx)^{\alpha}-(1-x)^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt+(-1)^{m}\frac{\alpha(1-x)^{\alpha(n+1)}x^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|x^{\alpha}e^{-i\alpha\pi}-(t(1-x))^{\alpha}|^{2}}\,\mathrm{d}t$
$=(-1)^{m}\mathrm{III}_{\alpha,m}^{(n)}(x)$
for $x\in(0,1)$. Furthermore, $\mathrm{II}_{\alpha,m}(x)=0$ if $\alpha\in(0,2)$ because $K_{\alpha}=\{1\}$. Therefore, it follows from (2.31) that for $\alpha\in(0,2)$,
$\mathrm{I}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{I}_{\alpha,m}^{(n)}(x).$ (2.32)
Because $\mathrm{I}_{\alpha,m}^{(n)}(x)$ is a polynomial in $\alpha$ thanks to Lemma 2.3, (2.32) holds true for arbitrary $\alpha>0$, which in turn implies that
$\mathrm{II}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{II}_{\alpha,m}^{(n)}(x),\qquad\alpha>0.$ (2.33)
In what follows, the integer $n\in\mathbb{N}$ is taken to be sufficiently large so that $\alpha n\geq m+1$.
Step 2. We give a uniform estimate of $\mathrm{II}_{\alpha,m}^{(n)}(x)$. Suppose that $\omega\in K_{\alpha}\setminus\{1\}$ and $x\in(0,1/2]$. For fixed $k\in\mathbb{N}_{0}$ with $0\leq k\leq m$, we have
$|W_{-\alpha nx}(m,k)(\alpha n)_{k}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}|$
$\leq C_{2,1}\{(\alpha nx)^{m-k}+1\}(\alpha n)^{k}x^{k}\{1-2x(1-x)(1-\mathrm{Re}(\omega))\}^{(\alpha n-k)/2}$
$\leq C_{2,1}\{(\alpha nx)^{m}+(\alpha nx)^{k}\}\{1-x(1-\mathrm{Re}(\omega))\}^{(\alpha n-k)/2}$
$\leq C_{2,1}\left(\frac{m}{e(\alpha n-k)/2}\right)^{m}\left(\frac{\alpha n}{1-\mathrm{Re}(\omega)}\right)^{m}+C_{2,1}\left(\frac{k}{e(\alpha n-k)/2}\right)^{k}\left(\frac{\alpha n}{1-\mathrm{Re}(\omega)}\right)^{k}$
$\leq C_{2,5}$
by using (2.18), (2.20), and Lemma 2.10. Thus, we have
$\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq(m+1)(\#(K_{\alpha})-1)C_{2,5}\leq C_{2,6},$
where $C_{2,6}>0$ depends on only $m\in\mathbb{N}$ and a compact set $K$ in which $\alpha$ takes values. By (2.33), we obtain that
$\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq C_{2,6}.$ (2.34)
Step 3. We next give a uniform estimate of $\mathrm{III}_{\alpha,m}^{(n)}(x)$. We write
$\mathrm{III}_{\alpha,m}^{(n)}(x)=\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)+(-1)^{m}\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(1-x)$
with
$\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)=\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
It follows from (2.20) that
$|h_{-\alpha nx}^{(m)}(-t)|$ $=\left|\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(-\alpha nx)(\alpha n)_{k}(-t)^{k}(1-t)^{\alpha n-k}\right|$
$\leq C_{2,7}\sum_{k=0}^{m}\{(\alpha nx)^{m-k}+1\}(\alpha n)^{k}t^{k}(1-t)^{\alpha n-k}.$
Thus, we have
$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,8}\sum_{k=0}^{m}x^{\alpha}(1-x)^{\alpha(n+1)}|\sin\alpha\pi|$
×$\hskip 28.45274pt\times\{(nx)^{m-k}+1\}\int_{0}^{1}\frac{t^{\alpha-1}n^{k}t^{k}(1-t)^{\alpha n-k}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
By noting
$n^{k}t^{k}(1-t)^{\alpha n-k}\leq n^{k}\left(\frac{k}{e(\alpha n-k)}\right)^{k}\leq C_{2,9}$
and Lemma 2.7, we have
$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,10}\sum_{k=0}^{m}(1-x)^{\alpha n}\{(nx)^{m-k}+1\}$
$\leq C_{2,10}\sum_{k=0}^{m}\{(1-x)^{\alpha n}(nx)^{m-k}+(1-x)^{\alpha n}\}$
$\leq C_{2,10}\sum_{k=0}^{m}\left\{\left(\frac{m-k}{e\alpha n}\right)^{m-k}n^{m-k}+1\right\}\leq C_{2,11},$
from which we conclude that
$\sup_{x\in(0,1)}|\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,12},$ (2.35)
where $C_{2,12}>0$ depends on only $m\in\mathbb{N}$ and $K$ in which $\alpha$ takes values.
By combining (2.31) with (2.34) and (2.35), we obtain that
$\sup_{x\in(0,1)}|(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)-\mathrm{I}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,13}.$
This implies (2.11) once we divide both sides of the above inequality by $(\alpha n)^{m}$ and take the supremum over $\alpha\in K$. ∎
We also have the following by applying Theorem 2.11 and (2.23).
Corollary 2.12.
Let $K$ be a compact subset of $(0,\infty)$. As $n\to\infty$,
$\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}\right|=\begin{cases}\displaystyle O\left(\frac{1}{n}\right)&\text{if }m=1,2,\vspace{2mm}\\ \displaystyle O\left(\frac{1}{n^{2}}\right)&\text{if }m=3,4.\end{cases}$
2.5. Characteristic function of the fractional binomial distribution
In this section, we obtain a quantitative estimate of the characteristic function of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, that is,
$\varphi_{\alpha,x}^{(n)}(\xi):=\sum_{j=0}^{n}e^{i\xi j}\mu_{\alpha,x}^{(n)}(\{j\}),\qquad\xi\in\mathbb{R}.$
We define $\underline{\alpha}:=2\lfloor\alpha/2\rfloor$ and $\overline{\alpha}:=2\lceil\alpha/2\rceil$ for $\alpha>0$. Moreover, we put
$\theta_{\alpha}:=\begin{cases}\{(\alpha-\underline{\alpha})\wedge(\overline{\alpha}-\alpha)\}\pi&\text{if $\alpha\notin\mathbb{N}$},\\ 2\pi&\text{if $\alpha\in\mathbb{N}$}.\end{cases}$
We then have the following lemma.
Lemma 2.13.
Let $\alpha>0$.
(1) If $\alpha\in\mathbb{N}$, we have
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}=\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}$ (2.36)
for $\lambda\in\mathbb{C}$.
(2) If $\alpha\notin\mathbb{N}$, we have
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$
$=\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 14.22636pt\times\Bigg{\{}\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\Bigg{\}}\,\mathrm{d}t$ (2.37)
for $\lambda\in D_{\alpha}:=\{re^{i\theta}\mid r>0,\,|\theta|<\theta_{\alpha}/\alpha\}$.
Proof.
We assume $\alpha\in\mathbb{N}$. Then, (2.36) holds for all $\lambda\in(0,1]$ by Proposition 1.2. Because both sides of (2.36) are holomorphic on $\mathbb{C}$, (2.36) holds for all $\lambda\in\mathbb{C}$ by analytic continuation.
We next assume $\alpha\notin\mathbb{N}$. It then follows from Proposition 1.2 that (2.37) holds for $\lambda\in(0,1]$. We now find a domain of $\lambda\in\mathbb{C}$ including $(0,1)$ in which the functions defined by both sides of (2.37) are holomorphic. Let us put $\lambda=re^{i\theta}$ with $r>0$ and $|\theta|<\pi$. Then, it suffices to confirm that $\lambda\in\mathbb{C}\setminus(-\infty,0]$, $\lambda\omega\in\mathbb{C}\setminus(-\infty,0]$ for all $\omega\in K_{\alpha}$, $\lambda^{\alpha}e^{\pm i\alpha\pi}\notin[0,\infty)$, and $\lambda^{-\alpha}e^{\pm i\alpha\pi}\notin[0,\infty)$. The condition $\lambda\in\mathbb{C}\setminus(-\infty,0]$ merely says that $|\theta|<\pi$. By noting that
$K_{\alpha}=\left\{\exp\left(i\frac{2k\pi}{\alpha}\right)\,:\,k\in\mathbb{Z},\,-\left\lfloor\frac{\alpha}{2}\right\rfloor\leq k\leq\left\lfloor\frac{\alpha}{2}\right\rfloor\right\},$
the condition $\lambda\omega\in\mathbb{C}\setminus(-\infty,0]$ for all $\omega\in K_{\alpha}$ is satisfied if
$\theta\in(-\pi,\pi)\cap\bigcap_{k\in\mathbb{Z},-\lfloor\alpha/2\rfloor\leq k\leq\lfloor\alpha/2\rfloor}\left(-\pi-\frac{2k\pi}{\alpha},\pi-\frac{2k\pi}{\alpha}\right)=\left(-\frac{\alpha-\underline{\alpha}}{\alpha}\pi,\frac{\alpha-\underline{\alpha}}{\alpha}\pi\right).$
Moreover, the conditions $\lambda^{\alpha}e^{\pm i\alpha\pi}\notin[0,\infty)$ and $\lambda^{-\alpha}e^{\pm i\alpha\pi}\notin[0,\infty)$ are satisfied if
$\arg(e^{\mp i\alpha\pi})-2\pi<\alpha\theta<\arg(e^{\mp i\alpha\pi})\quad\text{and}\quad\arg(e^{\mp i\alpha\pi})-2\pi<-\alpha\theta<\arg(e^{\mp i\alpha\pi}),$ (2.38)
where $\arg(e^{\mp i\alpha\pi})$ is taken from $(0,2\pi)$. Because $\arg(e^{-i\alpha\pi})=(\overline{\alpha}-\alpha)\pi$ and $\arg(e^{i\alpha\pi})=(\alpha-\underline{\alpha})\pi$, (2.38) is rewritten as $-\eta_{\alpha}<\theta<\eta_{\alpha}$, where
$\eta_{\alpha}=-\big{[}\{2-(\alpha-\underline{\alpha})\}\wedge\{2-(\overline{\alpha}-\alpha)\}\wedge(\overline{\alpha}-\alpha)\wedge(\alpha-\underline{\alpha})\big{]}\frac{\pi}{\alpha}.$
Noting that $\eta_{\alpha}=\theta_{\alpha}/\alpha$ because $\overline{\alpha}-\underline{\alpha}=2$, both sides of (2.37) turn out to be holmorphic in the domain $D_{\alpha}$. Thus, (2.37) also holds for all $\lambda\in D_{\alpha}$ by analytic continuation. ∎
We provide a simple lemma for the proof of Theorem 2.15 below.
Lemma 2.14.
For $a\geq 0$, $b\geq 0$ and $\theta\in\mathbb{R}$, we have
$|a-be^{i\theta}|\geq\sqrt{a^{2}+b^{2}}\left|\sin\frac{\theta}{2}\right|.$
Proof.
Because $|a-be^{i\theta}|=|b-ae^{-i\theta}|$, we may assume $a\leq b$. First, suppose $\cos\theta\leq 0$. Then, we have
$|a-be^{i\theta}|^{2}=a^{2}+b^{2}-2ab\cos\theta\geq a^{2}+b^{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.$
Next, suppose $\cos\theta>0$. Then, it holds that $\cos^{2}(\theta/2)=(1+\cos\theta)/2>1/2$. Therefore, we have
$|a-be^{i\theta}|^{2}$ $=(a-b\cos\theta)^{2}+b^{2}\sin^{2}\theta$
$=(a-b\cos\theta)^{2}+4b^{2}\sin^{2}\frac{\theta}{2}\cos^{2}\frac{\theta}{2}$
$\geq 2b^{2}\sin^{2}\frac{\theta}{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.\qed$
A quantitative estimate of $\varphi_{\alpha,x}^{(n)}$ is now stated as follows.
Theorem 2.15.
Let $\alpha>0$, $n\in\mathbb{N}$, $x\in(0,1)$, and $\xi\in\mathbb{R}$ satisfy that $|\xi|<\theta_{\alpha}$. Then, we have
$\varphi_{\alpha,x}^{(n)}(\xi)$ $=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega e^{i\xi/\alpha}\right)^{\alpha n}$
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{[}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)}}{\{(1-x)^{\alpha}e^{-i(\xi+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{]}\,\mathrm{d}t$ (2.39)
$=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+O(e^{-\delta n})$ (2.40)
as $n\to\infty$ for some $\delta>0$ that is chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\xi\in(-\theta_{\alpha},\theta_{\alpha})$. Here, if $\alpha\in\mathbb{N}$, then the last term of (2.39) is regarded as 00.
Proof.
We have
$\varphi_{\alpha,x}^{(n)}(\xi)$ $=\sum_{j=0}^{n}e^{i\xi j}\times\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$
$=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\left(\frac{xe^{i\xi/\alpha}}{1-x}\right)^{\alpha j}(1-x)^{\alpha n}.$
We put $\lambda=xe^{i\xi/\alpha}/(1-x)$. We first consider the case where $\alpha\notin\mathbb{N}$. From Lemma 2.13,
$\varphi_{\alpha,x}^{(n)}(\xi)$ $=\frac{1}{Z_{\alpha,x}^{(n)}}(1-x)^{\alpha n}\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$
$=\frac{1}{Z_{\alpha,x}^{(n)}}(1-x)^{\alpha n}\Bigg{[}\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\left\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\right\}\,\mathrm{d}t\Bigg{]}$
$=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega e^{i\xi/\alpha}\right)^{\alpha n}$
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n-1)}e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{[}\frac{(1-x)^{2\alpha}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha n}(1-x)^{-\alpha(n-2)}e^{i\xi n}}{\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}}\Bigg{]}\,\mathrm{d}t,$
which implies (2.39). Next, we prove (2.40). We have
$\left|1-x+x\omega e^{i\xi/\alpha}\right|^{2}$ $=(1-x)^{2}+2x(1-x)\mathrm{Re}\left(\omega e^{i\xi/\alpha}\right)+x^{2}$
$=1-2x(1-x)\left\{1-\mathrm{Re}\left(\omega e^{i\xi/\alpha}\right)\right\}<1$
for $\omega\in K_{\alpha}\setminus\{1\}$ because $\mathrm{Re}(\omega e^{i\xi/\alpha})<1$ from the assumption of $\xi$. Here, we also note that $K_{\alpha}\setminus\{1\}\neq\varnothing$ only if $\alpha\geq 2$. Moreover, from Lemma 2.14, we have
$|(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi\pm\alpha\pi)}|\geq x^{\alpha}\left|\sin\frac{\xi\pm\alpha\pi}{2}\right|$
and
$|(1-x)^{\alpha}e^{-i(\xi\pm\alpha\pi)}-(tx)^{\alpha}|\geq(1-x)^{\alpha}\left|\sin\frac{\xi\pm\alpha\pi}{2}\right|.$
Note that $|\xi|<\theta_{\alpha}$ implies $\sin(\xi\pm\alpha\pi)/2\neq 0$. From these estimates and (2.13), we obtain (2.40). The case for $\alpha\in\mathbb{N}$ is simpler to prove. ∎
2.6. The fractional multinomial distribution
Let $d\geq 2$ be an integer. For $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\mathbb{R}^{d}$, we write
$|\bm{x}|:=x_{1}+x_{2}+\cdots+x_{d},\qquad\|\bm{x}\|:=(x_{1}^{2}+x_{2}^{2}+\cdots+x_{d}^{2})^{1/2}.$
We define the $(d-1)$-simplex in $\mathbb{R}^{d}$ by
$\Delta_{d-1}:=\left\{\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\mathbb{R}^{d}\,|\,x_{i}\geq 0,i=1,2,\dots,d,\,|\bm{x}|=1\right\}.$
For $n\in\mathbb{N}$, we also set
$J_{d}^{(n)}=\{\bm{j}=(j_{1},j_{2},\dots,j_{d})\mid j_{k}\in\mathbb{N}_{0},\,k=1,2,\dots,d,\,|\bm{j}|=n\}.$
According to the preliminary version [FR11, Lemma 1] of the paper [FR13], we can generalize Proposition 1.1 to the multivariate cases, which is stated as follows.
Theorem 2.16 (Multivariate neo-classical inequality; see also [LCL07, Exercise 3.9]).
Let $\alpha\in(0,2)$, $n\in\mathbb{N}$ and $x_{1},x_{2},\dots,x_{d}\geq 0$. Then, we have
$\alpha^{d}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
$\begin{cases}\leq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(0,1],\\ \geq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(1,2).\\ \end{cases}$ (2.41)
The equality holds if and only if $\alpha=1$ or $x_{1}=x_{2}=\cdots=x_{d}=0$.
Proof.
The proof is by induction on $d\geq 2$. For $d=2$, (2.41) is nothing but the neo-classical inequality and its converse. Next, we assume that (2.41) holds for some $d>2$. Then for $\alpha\in(0,1]$, we have
$\alpha^{d+1}\sum_{\bm{j}\in J_{d+1}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d+1}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d+1}^{\alpha j_{d+1}}$
$=\alpha\sum_{j_{d+1}=0}^{n}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha(n-j_{d+1})+1)\Gamma(\alpha j_{d+1}+1)}x_{d+1}^{\alpha j_{d+1}}$
×$\hskip 28.45274pt\times\alpha^{d}\sum_{(j_{1},j_{2},\dots,j_{d})\in J_{d}^{(n-j_{d+1})}}\frac{\Gamma(\alpha(n-j_{d+1})+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
$\leq\alpha\sum_{j_{d+1}=0}^{n}\binom{\alpha n}{\alpha j_{d+1}}(x_{1}+x_{2}+\cdots+x_{d})^{\alpha(n-j_{d+1})}x_{d+1}^{\alpha j_{d+1}}$
$\leq(x_{1}+x_{2}+\cdots+x_{d+1})^{\alpha n}.$
The proof for $\alpha\in(1,2)$ is similar to that for $\alpha\in(0,1]$. ∎
Needless to say, if $\alpha=1$, the (2.41) reads as the usual multinomial theorem. As with Definition 1.3, this multivariate neo-classical inequality allows us to define the fractional analogue of the multinomial distribution.
Definition 2.17 (Fractional multinomial distribution).
Let $\alpha>0$, $n\in\mathbb{N}$ and $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\Delta_{d-1}$. Then, a probability measure $\mu_{\alpha,\bm{x}}^{(n)}$ on $J_{d}^{(n)}$ is called the $\alpha$-fractional multinomial distribution if
$\mu_{\alpha,\bm{x}}^{(n)}(\{\bm{j}\})=\frac{\alpha^{d}}{Z_{\alpha,\bm{x}}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$ (2.42)
for $\bm{j}\in J_{d}^{(n)}$, where $Z_{\alpha,\bm{x}}^{(n)}$ is the normalized constant given by
$Z_{\alpha,\bm{x}}^{(n)}=\alpha^{d}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}\,(>0).$
If $\alpha=1$, then (2.42) becomes the usual definition of the multinomial distribution. However, we have no explicit formula for the remainder term $R^{(n)}_{\alpha}(x_{1},x_{2},\dots,x_{d})$ satisfying
$\alpha^{d}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
$=(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}+R^{(n)}_{\alpha}(x_{1},x_{2},\dots,x_{d})$
for $d>2$. Obtaining such a formula explicitly might require a multivariate version of the fractional Taylor series, which although seeming rather difficult might nevertheless be worth investigating further.
3. Limit theorems for the fractional binomial distribution
3.1. The weak law of large numbers
First, we prove the weak law of large numbers for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
Theorem 3.1 (Weak law of large numbers for $\mu_{\alpha,x}^{(n)}$).
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. Let $S_{\alpha,x}^{(n)}$ denote the random variable whose law is $\mu_{\alpha,x}^{(n)}$. Then, for every $\varepsilon>0$ and a compact subset $K$ of $(0,\infty)$, we have
$\lim_{n\to\infty}\sup_{\alpha\in K}\sup_{x\in[0,1]}\mathbb{P}\left(\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\varepsilon\right)=0.$ (3.1)
Proof.
It follows from the Chebyshev inequality, Theorem 2.9, and Corollary 2.12 that
$\mathbb{P}\left(\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\varepsilon\right)$
$\leq\frac{1}{\varepsilon^{2}}\mathbb{E}\left[\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\right]$
$=\frac{1}{\varepsilon^{2}}\sum_{j=0}^{n}\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}$
$\leq\frac{1}{\varepsilon^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|=O\left(\frac{1}{n}\right)$
as $n\to\infty$, which implies (3.1). ∎
3.2. The central limit theorem
We show the central limit theorem for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, which is stated as follows.
Theorem 3.2 (Central limit theorem for $\mu_{\alpha,x}^{(n)}$).
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Let $\widetilde{\mu}_{\alpha,x}^{(n)}$ denote the probability law of the random variable
$\widetilde{S}_{\alpha,x}^{(n)}:=\frac{S_{\alpha,x}^{(n)}-\mathbb{E}[S_{\alpha,x}^{(n)}]}{\sqrt{\mathrm{Var}(S_{\alpha,x}^{(n)})}}.$
Then, $\widetilde{\mu}_{\alpha,x}^{(n)}$ converges weakly to the standard normal distribution $N(0,1)$ as $n\to\infty$.
Proof.
We write $m_{n}=\mathbb{E}[S_{\alpha,x}^{(n)}]$ and $v_{n}=\mathrm{Var}(S_{\alpha,x}^{(n)})$. Let $\widetilde{\varphi}_{\alpha,x}^{(n)}$ denote the characteristic function of $\widetilde{\mu}_{\alpha,x}^{(n)}$, and let $\xi\in\mathbb{R}$. We note that $v_{n}$ diverges as $n\to\infty$ from Corollary 2.8. Hence, it holds that $|\xi/\sqrt{v_{n}}|<\theta_{\alpha}$ for sufficiently large $n$, and we consider only such $n$ in the following. By applying Theorem 2.15, we have
$\widetilde{\varphi}_{\alpha,x}^{(n)}(\xi)$
$=\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\varphi_{\alpha,x}^{(n)}\left(\frac{\xi}{\sqrt{v_{n}}}\right)$
$=\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\Bigg{[}\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
$\hskip 28.45274pt+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha\sin\alpha\pi}{\pi}\exp\left(\frac{i\xi}{\sqrt{v_{n}}}\right)\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{\{}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)/\sqrt{v_{n}}}}{\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{\}}\,\mathrm{d}t\Bigg{]}.$ (3.2)
From Corollary 2.8, we have
$\frac{m_{n}}{n}=x+O(e^{-\delta n}),\qquad\sqrt{v_{n}}=\sqrt{\frac{1}{\alpha}nx(1-x)}+O(e^{-\delta n})=\Omega(\sqrt{n})$
as $n\to\infty$ for some $\delta>0$. Because
$e^{i\theta}=1+i\theta-\frac{\theta^{2}}{2}+O(\theta^{3})$
as $\theta\to 0$, we obtain
$\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\left(1-x+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
=$=\left\{(1-x)\exp\left(-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}\right)+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)\right)\right\}^{\alpha n}$
$=\Bigg{[}(1-x)\left\{1-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}-\frac{\xi^{2}m_{n}^{2}}{2(\alpha n\sqrt{v_{n}})^{2}}+O(n^{-3/2})\right\}$
$\hskip 28.45274pt+x\left\{1+\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)-\frac{\xi^{2}}{2(\alpha\sqrt{v_{n}})^{2}}\left(1-\frac{m_{n}}{n}\right)^{2}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
$=\Bigg{[}(1-x)\left\{1-\frac{i\xi x}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}x^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}$
$\hskip 28.45274pt+x\left\{1+\frac{i\xi(1-x)}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}(1-x)^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
=$=\left\{1-\frac{\xi^{2}}{2\alpha n}+O(n^{-3/2})\right\}^{\alpha n}\to\exp\left(-\frac{\xi^{2}}{2}\right)$ (3.3)
as $n\to\infty$. For each $\omega\in K_{\alpha}\setminus\{1\}$, we also have
$\left|\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}\right|=\left|1-x+x\omega+O(n^{-1/2})\right|^{\alpha n}\to 0$ (3.4)
as $n\to\infty$, by noting (2.18). To estimate the last term of (3.2), we may assume $\alpha\notin\mathbb{N}$; otherwise this term vanishes. Because it follows from Lemma 2.14 that
$|(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}\mp\alpha\pi)}|\geq x^{\alpha}\left|\sin\frac{1}{2}\left(\frac{\xi}{\sqrt{v_{n}}}\mp\alpha\pi\right)\right|\to x^{\alpha}\left|\sin\frac{\alpha\pi}{2}\right|>0$
and
$|(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}\pm\alpha\pi)}-(tx)^{\alpha}|\geq(1-x)^{\alpha}\left|\sin\frac{1}{2}\left(\frac{\xi}{\sqrt{v_{n}}}\pm\alpha\pi\right)\right|\to(1-x)^{\alpha}\left|\sin\frac{\alpha\pi}{2}\right|>0$
as $n\to\infty$, (2.13) implies that the third term of the right-hand side of (3.2) converges to zero as $n\to\infty$. Therefore, (3.2), (3.3), and (3.4) lead to
$\lim_{n\to\infty}\widetilde{\varphi}_{\alpha,x}^{(n)}(\xi)=\exp\left(-\frac{\xi^{2}}{2}\right),\qquad\xi\in\mathbb{R}.$
Because the right-hand side of the above is nothing but the characteristic function of the normal distribution $N(0,1)$, Lévy’s continuity theorem (cf. [Kle08, Theorem 15.23]) concludes the desired weak convergence. ∎
3.3. The law of small numbers
Recall that the Mittag-Leffler function of parameter $\alpha>0$ is defined by
$E_{\alpha}(z)=\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+1)},\qquad z\in\mathbb{C},$
which is a special function appearing in the context of fractional calculus. In terms of this function, we introduce a fractional analogue of the Poisson distribution.
Definition 3.3 (Fractional Poisson distribution).
Let $\alpha>0$ and $\lambda>0$. The $\alpha$-fractional Poisson distribution with parameter $\lambda$ is a probability measure $\nu_{\alpha,\lambda}$ on $\mathbb{N}_{0}$ defined as
$\nu_{\alpha,\lambda}(\{j\})=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)},\qquad j\in\mathbb{N}_{0}.$
The $1$-fractional Poisson distribution coincides with the Poisson distribution. Some basic properties of this fractional Poisson distribution will be summarized in the subsequent section. We note that $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions defined via the generalized Mittag-Leffler function $E_{\alpha,\beta}(z)$ given by (3.9) below. See [Her16, CO17] for some properties as well as its definition.
Here, we show that $\nu_{\alpha,\lambda}$ is captured through the law of small numbers for the fractional binomial distribution.
Theorem 3.4 (Law of small numbers for $\mu_{\alpha,x}^{(n)}$).
Let $\alpha>0$, $\lambda>0$, and $x\in(0,1)$. Suppose that two sequences $\{\alpha_{n}\}_{n=1}^{\infty}\subset(0,\infty)$ and $\{x_{n}\}_{n=1}^{\infty}\subset(0,1)$ satisfy that $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ as $n\to\infty$. Then, $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to $\nu_{\alpha,\lambda}$ as $n\to\infty$.
Proof.
The proof is split into two parts.
Step 1. We put
$b_{\alpha_{n},x_{n}}^{(n)}(j)=\alpha_{n}\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}\bm{1}_{[0,n]}(j),\qquad j\in\mathbb{N}_{0}.$
Because $\alpha_{n}\to\alpha>0$ and $\alpha_{n}nx_{n}\to\lambda>0$, there exist four constants $C_{3,1},C_{3,2},C_{3,3},C_{3,4}>0$ such that
$C_{3,1}\leq\alpha_{n}\leq C_{3,2},\qquad C_{3,3}\leq nx_{n}\leq C_{3,4},\qquad n\in\mathbb{N}.$ (3.5)
Moreover, the Stirling formula implies the existence of two constants $C_{3,5}>0$ and $C_{3,6}>0$ satisfying
$C_{3,5}s^{s+1/2}e^{-s}\leq\Gamma(s+1)\leq C_{3,6}s^{s+1/2}e^{-s},\qquad s\geq\alpha>0.$ (3.6)
We now give an estimate of each $b_{\alpha_{n},x_{n}}^{(n)}(j)$. Let $j\in\mathbb{N}_{0}$ be fixed. If $j=0$, then we have
$b_{\alpha_{n},x_{n}}^{(n)}(j)=\alpha_{n}(1-x_{n})^{\alpha_{n}n}\leq C_{3,2}\left(1-\frac{C_{3,3}}{n}\right)^{C_{3,1}n}\leq C_{3,2}e^{-C_{3,1}C_{3,3}}.$
Suppose that $j\geq 1$. If $n<j$, then $b_{\alpha_{n},x_{n}}^{(n)}(j)=0$. If $n=j$, then it follows from (3.5) that
$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\alpha_{n}x_{n}^{\alpha_{n}n}\leq C_{3,2}x_{n}^{C_{3,1}n}\leq C_{3,2}\left(\frac{C_{3,4}}{n}\right)^{C_{3,1}n}=C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}.$
On the other hand, if $n>j$, then (3.5) and (3.6) allow us to obtain
(0(0( 0 $\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$
$=\frac{\alpha_{n}\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}(n-j)+1)\Gamma(\alpha_{n}j+1)}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}$
$\leq\frac{C_{3,2}C_{3,6}(\alpha_{n}n)^{\alpha_{n}n+1/2}e^{-\alpha_{n}n}}{C_{3,5}\{\alpha_{n}(n-j)\}^{\alpha_{n}(n-j)+1/2}e^{-\alpha_{n}(n-j)}\Gamma(\alpha_{n}j+1)}\left(\frac{C_{3,4}}{n}\right)^{\alpha_{n}j}$
$\leq\frac{C_{3,2}C_{3,6}(C_{3,4}\alpha_{n})^{\alpha_{n}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}e^{-\alpha_{n}j}\left(\frac{n}{n-j}\right)^{\alpha_{n}(n-j)+1/2}.$
By noting that the function $(0,\infty)\ni s\longmapsto(1+s)^{1/s}\in(0,\infty)$ is non-increasing, we have
$\left(\frac{n}{n-j}\right)^{\alpha_{n}(n-j)}=\left\{\left(1+\frac{j}{n-j}\right)^{(n-j)/j}\right\}^{\alpha_{n}j}\leq e^{\alpha_{n}j}.$
Moreover, we have
$\left(\frac{n}{n-j}\right)^{1/2}\leq\left(\frac{j+1}{j+1-j}\right)^{1/2}=(j+1)^{1/2}.$
Bringing everything together, we have
$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ =$=\begin{cases}\displaystyle C_{3,2}e^{-C_{3,1}C_{3,3}}&\text{if }j=0,\vspace{2mm}\\ \displaystyle\frac{C_{3,2}C_{3,6}(C_{3,2}C_{3,4}\vee 1)^{C_{3,2}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}(j+1)^{1/2}&\text{if }j=1,2,\dots,n-1,\vspace{2mm}\\ \displaystyle C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}&\text{if }j=n,\\ 0&\text{if }j>n\end{cases}$
$\leq C_{3,7}\left\{\frac{C_{3,8}^{C_{3,2}j}}{\Gamma(C_{3,1}j+1)}\vee\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}\right\}$ (3.7)
for some constants $C_{3,7},C_{3,8}>0$. Hence, (3.3) leads to
$\sum_{j=0}^{\infty}\sup_{n\in\mathbb{N}}b_{\alpha_{n},x_{n}}^{(n)}(j)<\infty.$ (3.8)
Step 2. By applying the Stirling formula, we have
$\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}$ $=\frac{\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}j+1)\Gamma(\alpha_{n}(n-j)+1)}x_{n}^{\alpha_{n}j}$
$\sim\frac{\sqrt{2\pi}(\alpha_{n}n)^{\alpha_{n}n+1/2}e^{-\alpha_{n}n}}{\Gamma(\alpha_{n}j+1)\sqrt{2\pi}\{\alpha_{n}(n-j)\}^{\alpha_{n}(n-j)+1/2}e^{-\alpha_{n}(n-j)}}\left(\frac{\lambda}{\alpha_{n}n}\right)^{\alpha_{n}j}$
$=\frac{\lambda^{\alpha_{n}j}}{\Gamma(\alpha_{n}j+1)}e^{-\alpha_{n}j}\left\{\left(1-\frac{j}{n}\right)^{n}\right\}^{-\alpha_{n}(1-j/n)-1/(2n)}$
$\to\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}e^{-\alpha j}e^{\alpha j}=\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}$
as $n\to\infty$. Therefore, for $j\in\mathbb{N}_{0}$, we obtain
$b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\alpha_{n}\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}\bm{1}_{[0,n]}(j)$
$=\alpha_{n}\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}\left(1-\frac{1}{\alpha_{n}n}(\lambda+o(1))\right)^{\alpha_{n}(n-j)}\bm{1}_{[0,n]}(j)$
$\to\frac{\alpha\lambda^{\alpha j}}{\Gamma(\alpha j+1)}e^{-\lambda}$
as $n\to\infty$. Because
$Z_{\alpha_{n},x_{n}}^{(n)}=\sum_{j=0}^{n}\alpha_{n}\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}=\sum_{j=0}^{\infty}b_{\alpha_{n},x_{n}}^{(n)}(j),$
(3.8) and the dominated convergence theorem imply that
$\lim_{n\to\infty}Z_{\alpha_{n},x_{n}}^{(n)}=\sum_{j=0}^{\infty}\left(\lim_{n\to\infty}b_{\alpha_{n},x_{n}}^{(n)}(j)\right)=\sum_{j=0}^{\infty}\frac{\alpha\lambda^{\alpha j}}{\Gamma(\alpha j+1)}e^{-\lambda}=\alpha e^{-\lambda}E_{\alpha}(\lambda^{\alpha}).$
Thus, we obtain
$\mu_{\alpha_{n},x_{n}}^{(n)}(\{j\})=\frac{1}{Z_{\alpha_{n},x_{n}}^{(n)}}b_{\alpha_{n},x_{n}}^{(n)}(j)\to\frac{1}{\alpha e^{-\lambda}E_{\alpha}(\lambda^{\alpha})}\frac{\alpha\lambda^{\alpha j}}{\Gamma(\alpha j+1)}e^{-\lambda}=\nu_{\alpha,\lambda}(\{j\})$
as $n\to\infty$ for all $j\in\mathbb{N}_{0}$. ∎
3.4. Some properties of the fractional Poisson distribution
First, we seek the mean and variance of the fractional Poisson distribution $\nu_{\alpha,\lambda}$ introduced in Definition 3.3.
Lemma 3.5.
Let $S_{\alpha,\lambda}$ be a random variable whose probability law is $\nu_{\alpha,\lambda}$. Then, we have
$\mathbb{E}[S_{\alpha,\lambda}]$ $=\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})},$
$\mathrm{Var}(S_{\alpha,\lambda})$ $=\frac{\lambda^{\alpha}}{\alpha^{2}E_{\alpha}(\lambda^{\alpha})^{2}}\Big{[}\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]},$
where in general $E_{\alpha,\beta}(z)$ is the Mittag-Leffler function of two parameters $\alpha>0$ and $\beta\in\mathbb{C}$ defined by
$E_{\alpha,\beta}(z):=\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+\beta)},\qquad z\in\mathbb{C}.$ (3.9)
Proof.
Let $\varphi_{\alpha,\lambda}(\xi)$, $\xi\in\mathbb{R}$ be the characteristic function of $\nu_{\alpha,\lambda}$. It is straightforward to check that
$\varphi_{\alpha,\lambda}(\xi)$ $=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\sum_{j=0}^{\infty}e^{i\xi j}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\sum_{j=0}^{\infty}\frac{(\lambda^{\alpha}e^{i\xi})^{j}}{\Gamma(\alpha j+1)}=\frac{E_{\alpha}(\lambda^{\alpha}e^{i\xi})}{E_{\alpha}(\lambda^{\alpha})}$
for $\xi\in\mathbb{R}$. Because
$\frac{\mathrm{d}}{\mathrm{d}z}E_{\alpha}(z)$ $=\sum_{j=1}^{\infty}j\frac{z^{j-1}}{\Gamma(\alpha j+1)}=\frac{1}{\alpha}\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+\alpha)}=\frac{1}{\alpha}E_{\alpha,\alpha}(z)$ (3.10)
and
$\frac{\mathrm{d}}{\mathrm{d}z}E_{\alpha,\alpha}(z)$ $=\sum_{j=1}^{\infty}j\frac{z^{j-1}}{\Gamma(\alpha j+\alpha)}$
$=\sum_{j=0}^{\infty}(j+1)\frac{z^{j}}{\Gamma(\alpha j+2\alpha)}$
$=\frac{1}{\alpha}\sum_{j=0}^{\infty}(\alpha j+2\alpha-1)\frac{z^{j}}{\Gamma(\alpha j+2\alpha)}-\frac{\alpha-1}{\alpha}\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+2\alpha)}$
$=\frac{1}{\alpha}\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+2\alpha-1)}-\frac{\alpha-1}{\alpha}\sum_{j=0}^{\infty}\frac{z^{j}}{\Gamma(\alpha j+2\alpha)}$
$=\frac{1}{\alpha}E_{\alpha,2\alpha-1}(z)-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(z),$ (3.11)
we have
$\frac{\mathrm{d}}{\mathrm{d}\xi}\varphi_{\alpha,\lambda}(\xi)$ $=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\frac{1}{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})i\lambda^{\alpha}e^{i\xi}=i\frac{\lambda^{\alpha}e^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})}{\alpha E_{\alpha}(\lambda^{\alpha})}$ (3.12)
and
$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\varphi_{\alpha,\lambda}(\xi)$ $=i\frac{\lambda^{\alpha}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}ie^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
$\hskip 28.45274pt+i\lambda^{\alpha}e^{2i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}$
$=-\frac{\lambda^{\alpha}e^{i\xi}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
$\hskip 28.45274pt+\lambda^{\alpha}e^{i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}.$ (3.13)
By substituting $\xi=0$ for (3.12) and (3.13), we obtain
$\mathbb{E}[S_{\alpha,\lambda}]$ $=\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})},$
$\mathbb{E}[S_{\alpha,\lambda}^{2}]$ $=\frac{\lambda^{\alpha}}{\alpha^{2}E_{\alpha}(\lambda^{\alpha})}\big{[}\alpha E_{\alpha,\alpha}(\lambda^{\alpha})+\lambda^{\alpha}\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}\big{]}.$
The variance of $S_{\alpha,\lambda}$ is obtained straightforwardly from the two equalities above. ∎
As with the Poisson distribution, it is quite cumbersome to express both the raw and central moments of the fractional Poisson distribution, but we can derive the following recurrence formula for its central moments.
Theorem 3.6.
For $m\in\mathbb{N}_{0}$, we define
$M_{\alpha,\lambda}^{(m)}:=\mathbb{E}\big{[}(S_{\alpha,\lambda}-\mathbb{E}[S_{\alpha,\lambda}])^{m}\big{]}=\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\nu_{\alpha,\lambda}(\{j\}).$ (3.14)
Then, we have
$M_{\alpha,\lambda}^{(m+1)}$ $=\frac{\lambda}{\alpha}\frac{\mathrm{d}}{\mathrm{d}\lambda}M_{\alpha,\lambda}^{(m)}+mM_{\alpha,\lambda}^{(2)}M_{\alpha,\lambda}^{(m-1)}.$ (3.15)
Proof.
Thanks to (3.10) and (3.11), we have
$\frac{\mathrm{d}}{\mathrm{d}\lambda}E_{\alpha}(\lambda^{\alpha})$ $=\lambda^{\alpha-1}E_{\alpha,\alpha}(\lambda^{\alpha}),$
$\frac{\mathrm{d}}{\mathrm{d}\lambda}E_{\alpha,\alpha}(\lambda^{\alpha})$ $=\lambda^{\alpha-1}\left\{E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\right\}.$
Moreover, we have
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\nu_{\alpha,\lambda}(\{j\})$ $=-\frac{\lambda^{\alpha-1}E_{\alpha,\alpha}(\lambda^{\alpha})}{E_{\alpha}(\lambda^{\alpha})^{2}}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}+\frac{\alpha j}{\lambda E_{\alpha}(\lambda^{\alpha})}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}$
$=\frac{\alpha}{\lambda E_{\alpha}(\lambda^{\alpha})}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}$
$=\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
for $j\in\mathbb{N}_{0}$. By differentiating both sides of (3.14) with respect to $\lambda$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}\lambda}M_{\alpha,\lambda}^{(m)}$ $=-m\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m-1}\nu_{\alpha,\lambda}(\{j\})\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$
$\hskip 28.45274pt+\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
$=-mM_{\alpha,\lambda}^{(m-1)}\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)+\frac{\alpha}{\lambda}M_{\alpha,\lambda}^{(m+1)}.$
Because it holds that
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$
$=\frac{\lambda^{\alpha-1}}{\alpha E_{\alpha}(\lambda^{\alpha})^{2}}\Big{[}\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]}$
$=\frac{\alpha}{\lambda}\mathrm{Var}(S_{\alpha,\lambda})=\frac{\alpha}{\lambda}M_{\alpha,\lambda}^{(2)},$
we conclude (3.15). ∎
4. The fractional Bernstein operator and limit theorems for its iterates
4.1. The fractional Bernstein operator and its approximating property
Let $\alpha>0$. In this section, we show that the fractional Bernstein operator $B_{\alpha,n}$ introduced in Definition 1.8 uniformly approximates all continuous functions on $[0,1]$, similarly to Proposition 1.5.
Theorem 4.1.
Let $K$ be a compact subset of $(0,\infty)$. For $f\in C([0,1])$, we have
$\lim_{n\to\infty}\sup_{\alpha\in K}\|B_{\alpha,n}f-f\|_{\infty}=0.$
Proof.
Let $\alpha>0$, $n\in\mathbb{N}$, $x\in[0,1]$, and $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is $\mu_{\alpha,x}^{(n)}$. Then, $B_{\alpha,n}f$, $f\in C([0,1])$ can be expressed as
$B_{\alpha,n}f(x)=\mathbb{E}\left[f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)\right],\qquad x\in[0,1].$ (4.1)
Because $f\in C([0,1])$ is uniformly continuous on $[0,1]$, for $\varepsilon>0$, there exists some $\rho=\rho(\varepsilon)>0$ such that $|x-y|<\rho$ implies that $|f(x)-f(y)|<\varepsilon$. Then, it follows from (4.1) and the Chebyshev inequality that
$|B_{\alpha,n}f(x)-f(x)|$ $=\left|\mathbb{E}\left[f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right]\right|$
$\leq\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|<\rho\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right]$
$\leq\varepsilon+2\|f\|_{\infty}\mathbb{P}\left(\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right).$
Therefore, we obtain from Theorem 3.1 that
$\sup_{\alpha\in K}\|B_{\alpha,n}f-f\|_{\infty}=\varepsilon+O\left(\frac{1}{n}\right)$
as $n\to\infty$, which concludes the desired uniform convergence because $\varepsilon>0$ is chosen arbitrarily. ∎
4.2. The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
It is natural to consider the behaviors of the $k$-fold iterates $B_{\alpha,n}^{k}$ of the $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ as $k\to\infty$. In this section, we prove the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$ based on an argument similar to that in the proof of [Rus04, Theorem 1].
Theorem 4.2.
Let $\alpha>0$ and $n\in\mathbb{N}$. For $f\in C([0,1])$, there exists a unique function $f_{\alpha,n}\in C([0,1])$ satisfying $B_{\alpha,n}f_{\alpha,n}=f_{\alpha,n}$, $f_{\alpha,n}(0)=f(0)$, $f_{\alpha,n}(1)=f(1)$, and
$\lim_{k\to\infty}\|B_{\alpha,n}^{k}f-f_{\alpha,n}\|_{\infty}=0.$ (4.2)
In particular, we have
$f_{1,n}(x)=f(1)x+f(0)(1-x),\qquad x\in[0,1].$
Proof.
Let $\alpha>0$ and $n\in\mathbb{N}$. For $a,b\in\mathbb{R}$, we define a closed subset $C_{a,b}([0,1])$ of $C([0,1])$ by
$C_{a,b}([0,1]):=\{f\in C([0,1])\mid f(0)=a,f(1)=b\}.$
Then, we see that $f\in C_{a,b}([0,1])$ implies $B_{\alpha,n}f\in C_{a,b}([0,1])$ because $B_{\alpha,n}f(0)=f(0)$ and $B_{\alpha,n}f(1)=f(1)$. For $f,g\in C_{a,b}([0,1])$, we have
$|B_{\alpha,n}f(x)-B_{\alpha,n}g(x)|$ $=\left|\sum_{j=0}^{n}\mu_{\alpha,x}^{(n)}(\{j\})\left\{f\left(\frac{j}{n}\right)-g\left(\frac{j}{n}\right)\right\}\right|$
$=\left|\sum_{j=1}^{n-1}\mu_{\alpha,x}^{(n)}(\{j\})\left\{f\left(\frac{j}{n}\right)-g\left(\frac{j}{n}\right)\right\}\right|$
$\leq\|f-g\|_{\infty}\left|\sum_{j=0}^{n}\mu_{\alpha,x}^{(n)}(\{j\})-\frac{\alpha x^{\alpha n}}{Z_{\alpha,x}^{(n)}}-\frac{\alpha(1-x)^{\alpha n}}{Z_{\alpha,x}^{(n)}}\right|$
$=\|f-g\|_{\infty}\left(1-\frac{\alpha}{Z_{\alpha,x}^{(n)}}\{x^{\alpha n}+(1-x)^{\alpha n}\}\right).$
We note that
$x^{\alpha n}+(1-x)^{\alpha n}\geq\left(\frac{1}{2}\right)^{\alpha n},\qquad x\in[0,1].$
Moreover, we have
$\frac{\alpha}{Z_{\alpha,x}^{(n)}}\geq 1\wedge\alpha,\qquad x\in[0,1]$
by applying Theorem 2.9. Therefore, we obtain
$1-\frac{\alpha}{Z_{\alpha,x}^{(n)}}\{x^{\alpha n}+(1-x)^{\alpha n}\}\leq 1-(1\wedge\alpha)\left(\frac{1}{2}\right)^{\alpha n}<1,\qquad x\in[0,1],$
which means that $B_{\alpha,n}$ is contractive on $C_{a,b}([0,1])$ for $a,b\in\mathbb{R}$. Then, the contraction principle allows us to find the unique function $f_{\alpha,n}\in C([0,1])$ satisfying $B_{\alpha,n}f_{\alpha,n}=f_{\alpha,n}$, $f_{\alpha,n}(0)=f(0)$, $f_{\alpha,n}(1)=f(1)$, and (4.2). The explicit representation of $f_{1,n}$ was obtained in Proposition 1.6. ∎
4.3. Limit theorems for iterates of the fractional Bernstein operator
Let $\alpha>0$ be fixed. In this section, we identify some limiting behaviors of the iterate of $B_{\alpha,n}$. Let $\mathcal{L}_{\alpha}$ be the second-order differential operator acting on $C^{2}([0,1])$ defined by
$\mathcal{L}_{\alpha}f(x):=\frac{1}{2\alpha}x(1-x)f^{\prime\prime}(x)$
for $f\in C^{2}([0,1])$ and $x\in[0,1]$. The following property is well known.
Lemma 4.3 (cf. [EK86, Theorem 8.2.8] and [AC94, Theorem 6.2.6]).
The linear operator $(\mathcal{L}_{\alpha},C^{2}([0,1]))$ is closable on $C([0,1])$ and its closure $(\mathcal{L}_{\alpha},\mathrm{Dom}(\mathcal{L}_{\alpha}))$ generates the Feller semigroup $(T_{\alpha}(t))_{t\geq 0}$ on $C([0,1])$.
We prove that the discrete semigroup defined by the iterates of $B_{\alpha,n}$ converges to $(T_{\alpha}(t))_{t\geq 0}$ as follows.
Theorem 4.4.
Let $\alpha>0$. Then, we have
$\lim_{n\to\infty}\|B_{\alpha,n}^{\lfloor nt\rfloor}f-T_{\alpha}(t)f\|_{\infty}=0,\qquad f\in C([0,1]),\,t\geq 0.$ (4.3)
In order to prove Theorem 4.4, we prove the Voronovskaya-type theorem which asserts the convergence of the sequence $\{n(B_{\alpha,n}-I)\}_{n=1}^{\infty}$ to $\mathcal{L}_{\alpha}$, where $I$ denotes the identity operator on $C([0,1])$. Let us put
$\mathcal{D}_{0}:=\{f\in C^{2}([0,1])\mid f^{\prime}(0)=f^{\prime}(1)=0\}.$
Lemma 4.5.
Let $K$ be a compact subset of $(0,\infty)$. Then, we have
$\lim_{n\to\infty}\sup_{\alpha\in K}\|n(B_{\alpha,n}-I)f-\mathcal{L}_{\alpha}f\|_{\infty}=0,\qquad f\in\mathcal{D}_{0}.$ (4.4)
Proof.
Let $f\in\mathcal{D}_{0}$. We apply (4.1) and Taylor’s formula to obtain
$n(B_{\alpha,n}-I)f(x)-\mathcal{L}_{\alpha}f(x)$
$=n\mathbb{E}\left[f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right]-\mathcal{L}_{\alpha}f(x)$
$=nf^{\prime}(x)\mathbb{E}\left[\frac{1}{n}S_{\alpha,x}^{(n)}-x\right]$
$\hskip 28.45274pt+n\mathbb{E}\Bigg{[}\int_{0}^{1}(1-t)\left\{f^{\prime\prime}\left(x+t\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)\right)-f^{\prime\prime}(x)\right\}\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\,\mathrm{d}t\Bigg{]}$
$\hskip 28.45274pt+\frac{n}{2Z_{\alpha,x}^{(n)}}\left\{\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}-\frac{1}{\alpha n}x(1-x)\right\}f^{\prime\prime}(x)$
$\hskip 28.45274pt+\frac{1}{2\alpha}\left(\frac{1}{Z_{\alpha,x}^{(n)}}-1\right)x(1-x)f^{\prime\prime}(x)$
$=:\mathrm{I}_{\alpha}^{(n)}(x)+\mathrm{II}_{\alpha}^{(n)}(x)+\mathrm{III}_{\alpha}^{(n)}(x)+\mathrm{IV}_{\alpha}^{(n)}(x).$ (4.5)
Let $\varepsilon>0$. By assumption, there exists some $r>0$ such that
$|f^{\prime}(x)|<\varepsilon,\qquad x\in[0,r]\cup[1-r,1].$
Then, it follows from Proposition 2.6, Theorem 2.9, and Corollary 2.12 that
$\left|\mathrm{I}_{\alpha}^{(n)}(x)\right|$ $\leq n\max_{x\in[0,r]\cup[1-r,1]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right||f^{\prime}(x)|$
$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\max_{x\in[r,1-r]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|$
$\leq n\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|\varepsilon$
$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)O(e^{-\delta n})$
$=\varepsilon O(1)+nO(e^{-\delta n})=\varepsilon O(1)+o(1)$ (4.6)
as $n\to\infty$. Because $f^{\prime\prime}$ is uniformly continuous on $[0,1]$, there exists some $\rho=\rho(\varepsilon)\in(0,1)$ such that $|x-y|<\rho$ implies $|f^{\prime\prime}(x)-f^{\prime\prime}(y)|<\varepsilon$. Hence, Theorem 2.9 and Corollary 2.12 allow us to obtain that
$\left|\mathrm{II}_{\alpha}^{(n)}(x)\right|$
$\leq\frac{n\varepsilon}{2}\max_{x\in[0,1]}\mathbb{E}\left[\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\right]+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\max_{x\in[0,1]}\mathbb{E}\left[\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{4}\right]$
$\leq\frac{n\varepsilon}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|$
$\hskip 14.22636pt+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{4}\right|$
$=n\varepsilon O\left(\frac{1}{n}\right)+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}O\left(\frac{1}{n^{2}}\right)$
$=\varepsilon O(1)+O\left(\frac{1}{n}\right)$ (4.7)
as $n\to\infty$. Furthermore, by applying Theorems 2.9 and 2.11, we have
$\left|\mathrm{III}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{n}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\|f^{\prime\prime}\|_{\infty}O\left(\frac{1}{n^{2}}\right)=O\left(\frac{1}{n}\right)$ (4.8)
as $n\to\infty$. On the other hand, for $\varepsilon>0$, there exists some sufficiently small $u=u(\varepsilon)>0$ satisfying $x(1-x)<\varepsilon$ for $x\in[0,u]\cup[1-u,1]$. Then, we have
$\left|\mathrm{IV}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\max_{x\in[0,u]\cup[1-u,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
$\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\sup_{\alpha\in K}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
$=\varepsilon O(1)+o(1)$ (4.9)
as $n\to\infty$, by virtue of (2.13) and Theorem 2.9. Therefore, by combining (4.5) with (4.3), (4.7), (4.8), and (4.9), we obtain
$\sup_{\alpha\in K}\|n(B_{\alpha,n}-I)f-\mathcal{L}_{\alpha}f\|_{\infty}=\varepsilon O(1)+o(1)$
as $n\to\infty$. Because $\varepsilon>0$ is arbitrary, we conclude (4.4) after letting $\varepsilon\searrow 0$. ∎
Lemma 4.6.
$\mathcal{D}_{0}$ is a core for $(\mathcal{L}_{\alpha},\mathrm{Dom}(\mathcal{L}_{\alpha}))$.
Proof.
Because $C^{2}([0,1])$ is a core for $(\mathcal{L}_{\alpha},\mathrm{Dom}(\mathcal{L}_{\alpha}))$, it is sufficient to show that for $f\in C^{2}([0,1])$, there exists a family $(f_{\varepsilon})_{\varepsilon>0}\subset\mathcal{D}_{0}$ such that
$\lim_{\varepsilon\searrow 0}(\|\mathcal{L}_{\alpha}(f_{\varepsilon}-f)\|_{\infty}+\|f_{\varepsilon}-f\|_{\infty})=0.$ (4.10)
In the sequel, we assume that $f\in C^{2}([0,1])$ satisfies $f^{\prime}(0)\neq 0$ and $f^{\prime}(1)\neq 0$. If $f^{\prime}(0)=0$ or $f^{\prime}(1)=0$, then the proof is also valid by minor modification. Let $\varepsilon>0$ be sufficiently small so that $\varepsilon<1/2$ and
$\varepsilon\left(\log 2-\frac{1}{2}\right)<\min\{|f^{\prime}(0)|,|f^{\prime}(1)|\}.$ (4.11)
For $\delta\in(0,\varepsilon]$, we define
$g_{\delta}(x)$ $:=\frac{\varepsilon}{x+\delta}-\frac{\varepsilon}{\varepsilon+\delta},\qquad x\in[0,\varepsilon],$
$\widehat{g}_{\delta}(x)$ $:=\frac{\varepsilon}{1-x+\delta}-\frac{\varepsilon}{\varepsilon+\delta},\qquad x\in[1-\varepsilon,1].$
The function $g_{\delta}$ (resp. $\widehat{g}_{\delta}$) is non-negative and continuous on $[0,\varepsilon]$ (resp. $[1-\varepsilon,1]$) with $g_{\delta}(\varepsilon)=0$ (resp. $\widehat{g}_{\delta}(1-\varepsilon)=0$). We also define
$G(\delta)$ $:=\int_{0}^{\varepsilon}g_{\delta}(x)\,\mathrm{d}x=\varepsilon\log\frac{\varepsilon+\delta}{\delta}-\frac{\varepsilon^{2}}{\varepsilon+\delta},\qquad\delta\in(0,\varepsilon],$
$\widehat{G}(\delta)$ $:=\int_{1-\varepsilon}^{1}\widehat{g}_{\delta}(x)\,\mathrm{d}x=\varepsilon\log\frac{\varepsilon+\delta}{\delta}-\frac{\varepsilon^{2}}{\varepsilon+\delta},\qquad\delta\in(0,\varepsilon].$
Then, both $G$ and $\widehat{G}$ are continuous on $(0,\varepsilon]$ and satisfy that
$G(0+)$ =+=$\infty$,= + ∞ , $\qquad G(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(0)|,$
$\widehat{G}(0+)$ =+=$\infty$,= + ∞ , $\qquad\widehat{G}(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(1)|$
due to (4.11). Hence, the mean value theorem implies the existences of $\delta_{0},\delta_{1}\in(0,\varepsilon)$ such that $G(\delta_{0})=|f^{\prime}(0)|$ and $\widehat{G}(\delta_{1})=|f^{\prime}(1)|$. We now define a family of continuous functions $(f_{\varepsilon})_{\varepsilon>0}\subset C([0,1])$ by
$f_{\varepsilon}(x):=\begin{cases}f(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}\int_{y}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in[0,\varepsilon),\\ f(x)&\text{if }x\in[\varepsilon,1-\varepsilon],\\ f(x)+\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\int_{1-\varepsilon}^{y}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
Then, we have
$f_{\varepsilon}^{\prime}(x):=\begin{cases}f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s&\text{if }x\in[0,\varepsilon),\\ f^{\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
In particular, we see that
$f^{\prime}_{\varepsilon}(0)$ $=f^{\prime}(0)-\mathrm{sgn}\{f^{\prime}(0)\}|f^{\prime}(0)|=0,$
$f^{\prime}_{\varepsilon}(1)$ $=f^{\prime}(1)-\mathrm{sgn}\{f^{\prime}(1)\}|f^{\prime}(1)|=0,$
$f^{\prime}_{\varepsilon}(\varepsilon\pm)$ $=f^{\prime}(\varepsilon),\qquad f^{\prime}_{\varepsilon}((1-\varepsilon)\pm)=f^{\prime}(1-\varepsilon).$
We also have
$f_{\varepsilon}^{\prime\prime}(x):=\begin{cases}f^{\prime\prime}(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle g_{\delta_{0}}(x)&\text{if }x\in[0,\varepsilon),\\ f^{\prime\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\widehat{g}_{\delta_{1}}(x)&\text{if }x\in(1-\varepsilon,1]\\ \end{cases}$
and
$f_{\varepsilon}^{\prime\prime}(\varepsilon\pm)=f^{\prime\prime}(\varepsilon),\qquad f_{\varepsilon}^{\prime\prime}((1-\varepsilon)\pm)=f^{\prime\prime}(1-\varepsilon).$
These mean that $f_{\varepsilon}\in\mathcal{D}_{0}$ for $\varepsilon>0$. Furthermore, we obtain that
$|\mathcal{L}_{\alpha}(f_{\varepsilon}-f)(x)|$ $=\left|\frac{1}{\alpha}x(1-x)g_{\delta_{0}}(x)\right|$
$=\frac{1}{\alpha}x(1-x)\frac{\varepsilon(\varepsilon-x)}{(x+\delta_{0})(\varepsilon+\delta_{0})}\leq\frac{(1-x)(\varepsilon-x)}{\alpha}\leq\frac{\varepsilon}{\alpha}$
for $x\in[0,\varepsilon)$ and
$|\mathcal{L}_{\alpha}(f_{\varepsilon}-f)(x)|$ $=\left|\frac{1}{\alpha}x(1-x)\widehat{g}_{\delta_{1}}(x)\right|$
$=\frac{1}{\alpha}x(1-x)\frac{\varepsilon\{\varepsilon-(1-x)\}}{(1-x+\delta_{1})(\varepsilon+\delta_{1})}\leq\frac{x\{\varepsilon-(1-x)\}}{\alpha}\leq\frac{\varepsilon}{\alpha}$
for $x\in(1-\varepsilon,1]$. On the other hand, it holds that $|\mathcal{L}_{\alpha}(f_{\varepsilon}-f)(x)|=0$ for $x\in[\varepsilon,1-\varepsilon]$. Hence, we have
$\|\mathcal{L}_{\alpha}(f_{\varepsilon}-f)\|_{\infty}\leq\frac{\varepsilon}{\alpha},\qquad\varepsilon>0.$ (4.12)
We also observe that
$|f_{\varepsilon}(x)-f(x)|$ $=\int_{x}^{\varepsilon}\int_{y}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s\,\mathrm{d}y$
$\leq\int_{0}^{\varepsilon}\int_{0}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s\,\mathrm{d}y=\int_{0}^{\varepsilon}G(\delta_{0})\,\mathrm{d}y=\varepsilon|f^{\prime}(0)|$
for $x\in[0,\varepsilon)$ and
$|f_{\varepsilon}(x)-f(x)|$ $=\int_{1-\varepsilon}^{x}\int_{1-\varepsilon}^{y}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s\,\mathrm{d}y$
$\leq\int_{1-\varepsilon}^{1}\int_{1-\varepsilon}^{1}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s\,\mathrm{d}y=\int_{1-\varepsilon}^{1}\widehat{G}(\delta_{1})\,\mathrm{d}y=\varepsilon|f^{\prime}(1)|$
for $x\in(1-\varepsilon,1]$. Because $|f_{\varepsilon}(x)-f(x)|=0$ for $x\in[\varepsilon,1-\varepsilon]$ by definition, we obtain
$\|f_{\varepsilon}-f\|_{\infty}\leq\varepsilon\max\{|f^{\prime}(0)|,|f^{\prime}(1)|\}.$ (4.13)
Therefore, (4.10) follows from (4.12) and (4.13). ∎
We can now prove Theorem 4.4.
Proof of Theorem 4.4.
It follows from Lemma 4.6 and [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, by combining this fact with Lemma 4.5, we can apply Trotter’s approximation theorem (cf. [Tro58]) to conclude (4.3). ∎
Finally, we mention that the limiting semigroup $(T_{\alpha}(t))_{t\geq 0}$ obtained in Theorem 4.4 has the following probabilistic interpretation.
Proposition 4.7.
For $\alpha>0$ and $x\in[0,1]$, let $(X_{\alpha,x}(t))_{t\geq 0}$ be the one-dimensional diffusion process that is a unique strong solution to the stochastic differential equation
$\mathrm{d}X_{\alpha,x}(t)=\sqrt{\frac{1}{\alpha}X_{\alpha,x}(t)(1-X_{\alpha,x}(t))}\,\mathrm{d}W(t),\qquad X_{\alpha,x}(0)=x,$ (4.14)
where $(W(t))_{t\geq 0}$ is standard one-dimensional Brownian motion. Then, we have
$T_{\alpha}(t)f(x)=\mathbb{E}\big{[}f\big{(}X_{\alpha,x}(t)\big{)}\big{]},\qquad f\in C([0,1]),\,x\in[0,1],\,t\geq 0.$
Proof.
Because (4.14) has a unique strong solution, $(X_{\alpha,x}(t))_{t\geq 0}$ is the unique solution to the martingale problem for $(\mathcal{L}_{\alpha},\delta_{x})$ for every $x\in[0,1]$. Moreover, we note that the diffusion coefficient $(2\alpha)^{-1}x(1-x)$ of $\mathcal{L}_{\alpha}$ is continuous and bounded on $[0,1]$. Then, we can apply [Kal21, Theorem 32.11] to conclude that
$\widetilde{T}_{\alpha}(t)f(x):=\mathbb{E}[f(X_{\alpha,x}(t))],\qquad f\in C([0,1]),\,x\in[0,1],\,t\geq 0$
is a Feller semigroup on $C([0,1])$ and $\mathcal{L}_{\alpha}$ uniquely extends to the associated generator. Because a Feller semigroup is uniquely determined by its generator (cf. [Kal21, Lemma 17.5]), we have $T_{\alpha}(t)=\widetilde{T}_{\alpha}(t)$ for all $t\geq 0$. ∎
4.4. The fractional Bernstein operator with $\alpha=\alpha(x)$ being a function of $x$
In this section, we consider the fractional Bernstein operator whose parameter $\alpha>0$ is replaced with a function $\alpha=\alpha(x)\colon[0,1]\to(0,\infty)$. In what follows, we assume that there exist some positive constants $\alpha_{L}>0$ and $\alpha_{U}>0$ such that
$(0<)\,\alpha_{L}\leq\alpha(x)\leq\alpha_{U},\qquad x\in[0,1].$ (4.15)
Let $n\in\mathbb{N}$ and $x\in[0,1]$. Then, the $\alpha(x)$-fractional binomial distribution
$\mu_{\alpha(x),x}^{(n)}(\{j\})=\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)},\qquad x\in[0,1],\,j=0,1,2,\dots,n$
with (4.15) turns out to have almost the same properties as those of $\mu_{\alpha,x}^{(n)}$ discussed in Section 2. Indeed, the fractional-order Taylor series (Proposition 2.1) is still valid if we replace $\alpha$ with $\alpha(x)$ satisfying (4.15). Hence, Theorems 2.5 and 2.11, Proposition 2.6, and Corollaries 2.8 and 2.12 also hold for the $\alpha(x)$-fractional binomial distribution. In particular, Theorem 2.9 and (4.15) imply the existence of two constants $C_{4,1}>0$ and $C_{4,2}>0$ independent of $n\in\mathbb{N}$ such that
$C_{4,1}\leq Z_{\alpha(x),x}^{(n)}\leq C_{4,2},\qquad x\in[0,1].$
We now assume that $\alpha(x)$ is continuous in $x$. Let us consider the $\alpha(x)$-fractional Bernstein operator $B_{\alpha(\cdot),n}$ given by
$B_{\alpha(\cdot),n}f(x)$ $=\sum_{j=0}^{n}\mu_{\alpha(x),x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
$=\sum_{j=0}^{n}\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)}f\left(\frac{j}{n}\right)$
for $f\in C([0,1])$ and $x\in[0,1]$. We also consider the second-order differential operator
$\mathcal{L}_{\alpha(\cdot)}f(x)=\frac{1}{2\alpha(x)}x(1-x)f^{\prime\prime}(x)$
for $f\in C^{2}([0,1])$ and $x\in[0,1]$. Then, the operator $(\mathcal{L}_{\alpha(\cdot)},C^{2}([0,1]))$ is also closable on $C([0,1])$, and its closure $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$ generates the Feller semigroup $(T_{\alpha(\cdot)}(t))_{t\geq 0}$ on $C([0,1])$, which is no longer the Wright–Fisher diffusion semigroup in general (see [AC94, Theorems 1.6.11 and 6.2.6]).
By following the proof of Theorem 4.4, we also obtain that the discrete semigroup defined by the iterates of $B_{\alpha(\cdot),n}$ converges to $(T_{\alpha(\cdot)}(t))_{t\geq 0}$.
Theorem 4.8.
Suppose that a function $\alpha=\alpha(x)\colon[0,1]\to(0,\infty)$ is continuous. Then, we have
$\lim_{n\to\infty}\|B_{\alpha(\cdot),n}^{\lfloor nt\rfloor}f-T_{\alpha(\cdot)}(t)f\|_{\infty}=0,\qquad f\in C([0,1]),\,t\geq 0.$ (4.16)
Proof.
From Lemma 4.5, it holds that
$\lim_{n\to\infty}\|n(B_{\alpha(\cdot),n}-I)f-\mathcal{L}_{\alpha(\cdot)}f\|_{\infty}=0$
for $f\in C^{2}([0,1])$, by noting (4.15). On the other hand, Lemma 4.6 tells us that for $f\in C^{2}([0,1])$, there exists a family $(f_{\varepsilon})_{\varepsilon>0}\subset\mathcal{D}_{0}$ satisfying (4.10) with $\alpha$ replaced by $\alpha_{L}$. Then, (4.15) leads to
$\|\mathcal{L}_{\alpha(\cdot)}(f_{\varepsilon}-f)\|_{\infty}+\|f_{\varepsilon}-f\|_{\infty}\leq\|\mathcal{L}_{\alpha_{L}}(f_{\varepsilon}-f)\|_{\infty}+\|f_{\varepsilon}-f\|_{\infty}\to 0$
as $\varepsilon\searrow 0$, which implies that $\mathcal{D}_{0}$ is also a core for $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$. Therefore, it follows from [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha(\cdot)}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, Trotter’s approximation theorem allows us to obtain (4.16). ∎
Furthermore, we obtain a probabilistic interpretation of Theorem 4.8 in terms of a diffusion process corresponding to $\mathcal{L}_{\alpha(\cdot)}$. Consequently, we can capture more-general diffusion processes on $[0,1]$ via Theorem 4.8 and Proposition 4.9 as follows.
Proposition 4.9.
For $x\in[0,1]$, let $(X_{\alpha(\cdot),x}(t))_{t\geq 0}$ be the one-dimensional diffusion process that solves the stochastic differential equation
$\mathrm{d}X_{\alpha(\cdot),x}(t)=\sqrt{\frac{1}{\alpha(X_{\alpha(\cdot),x}(t))}X_{\alpha(\cdot),x}(t)(1-X_{\alpha(\cdot),x}(t))}\,\mathrm{d}W(t),\qquad X_{\alpha(\cdot),x}(0)=x,$ (4.17)
where $(W(t))_{t\geq 0}$ is standard one-dimensional Brownian motion. Then, we have
$T_{\alpha(\cdot)}(t)f(x)=\mathbb{E}\big{[}f\big{(}T_{\alpha(\cdot),x}(t)\big{)}\big{]},\qquad f\in C([0,1]),\,x\in[0,1],\,t\geq 0.$
Here, we note that (4.17) has the unique strong solution by applying the Yamada–Watanabe theorem (cf. [YW71, Theorem 1]). We omit the proof of Proposition 4.9 because it is similar to that of Proposition 4.7.
Acknowledgements. The first-named author was supported by JSPS KAKENHI Grant Number 19H00643. The second-named author was supported by JSPS KAKENHI Grant Number 23K12986.

### v3

@@
 Fractional binomial distributions induced by the generalized binomial theorem and their applications
-Abstract.
-A generalized binomial theorem was established by Hara and Hino [Bull.
+Abstract
+We develop a fractional extension of the classical binomial distribution and the associated Bernstein operator, formulated within the framework of the generalized binomial theorem (Hara and Hino [Bull.
 London Math.
 Soc.
-42 (2010), 467–477] for proving the neo-classical inequality.
-We introduce a new fractional analogue of the binomial distribution on the basis of the generalized binomial theorem and study its properties.
-They include the law of large numbers, the central limit theorem, and the law of small numbers, the last of which leads to a fractional analogue of the Poisson distribution.
-As an application of the fractional binomial distribution, a fractional analogue of the Bernstein operator is proposed, which acts on the space of continuous functions on the unit interval.
-We show that the fractional Bernstein operator uniformly approximates every continuous function on the unit interval.
-Some limit theorems for its iterates are discussed as well.
-In particular, a generalization of the Wright–Fisher diffusion process arising in the study of population genetics is captured as a certain limit of the iterates.
-Key words and phrases:
-fractional binomial distribution, neo-classical inequality, fractional Taylor expansion, fractional Bernstein operator
-2020 Mathematics Subject Classification:
-Primary 60E05; Secondary 26A33, 60F05, 41A36.
-1. Introduction
-1.1. The neo-classical inequality and the generalized binomial theorem
-Lyons [Lyo98] initiated the theory of rough paths and introduced the neo-classical inequality in order to show a fundamental theorem of rough path theory (see [Lyo98, Theorem 2.2.1]).
-This asserts that
-$\alpha^{2}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}$ (1.1)
-for $\alpha\in(0,1]$, $n\in\mathbb{N}$ and $x,y\geq 0$. Here, we define
+42 (2010), 467–477]).
+This provides a new probabilistic structure not representable as the law of the sum of independent and identically distributed random variables.
+Despite this nonstandard nature, we establish several of its fundamental analytic and probabilistic properties, including limit theorems, through a unified framework based on the generalized binomial theorem.
+We further analyze the properties of the fractional Bernstein operator associated with the fractional binomial distribution.
+In particular, we prove that the iterates of the operator converge to a generalized Wright–Fisher diffusion semigroup after a proper diffusive rescaling.
+keywords:
+fractional binomial distribution, fractional Taylor expansion, fractional Bernstein operator
+pacs:
+MSC Classification]60E05, 26A33, 60F05, 41A36.
+1 Introduction
+1.1 Fractional binomial distribution and fractional Bernstein operator
+The binomial distribution and the associated Bernstein operator form one of the classical bridges between probability and approximation theory. The classical binomial distribution
+$\mu_{x}^{(n)}(\mathrm{d}z):=\sum_{j=0}^{n}\binom{n}{j}x^{j}(1-x)^{n-j}\delta_{j}(\mathrm{d}z),\qquad n\in\mathbb{N},\ x\in(0,1),$
+arises naturally from the classical binomial theorem $(x+y)^{n}=\sum_{j=0}^{n}\binom{n}{j}x^{j}y^{n-j}$.
+Here, $\delta_{j}$ denotes the Dirac measure at $j$.
+It corresponds to the law of the sum of independent and identically distributed (i.i.d.
+in short) Bernoulli random variables, and induces the celebrated Bernstein operator, acting on the set $C([0,1])$ consisting of continuous functions on $[0,1]$, defined as
+$B_{n}f(x):=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{x}^{(n)}(\mathrm{d}z)=\sum_{j=0}^{n}\binom{n}{j}x^{j}(1-x)^{n-j}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]).$
+Bernstein proved in [3] that $\|B_{n}f-f\|_{\infty}\to 0$ holds as $n\to\infty$ for $f\in C([0,1])$, where $\|\cdot\|_{\infty}$ denotes the supremum norm.
+We refer to e.g., [4, 1, 2] for further details.
+Moreover,
+$B_{n}^{k}:=\underbrace{B_{n}\circ\cdots\circ B_{n}}_{k},\qquad k\in\mathbb{N},$
+forms a discrete semigroup, whose certain scaling limit as $n\to\infty$ yields the Wright–Fisher diffusion semigroup studied in population genetics. See [15, 7, 18].
+While such a classical theory has been studied intensively and extensively in probability theory, the fractional analogues of such structures remain largely unexplored.
+It is worth noting that the generalized binomial theorem obtained by Hara–Hino [12] provides a natural framework for such an extension.
+In general, for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq 0\}$ and $\gamma\in\mathbb{R}$, the power $z^{\gamma}$ is defined as $\exp(\gamma\,\mathrm{Log}\,z)$, where $\mathrm{Log}\,z$ is the principal value of $\log z$ so that $\mathrm{Log}\,1=0$.
+For $\alpha>0$, put
+$K_{\alpha}:=\{e^{i\theta}\in\mathbb{C}\mid-\pi<\theta\leq\pi,\,e^{i\theta\alpha}=1\}.$ (1.1)
+Note that $K_{\alpha}=\{1\}$ when $\alpha\in(0,2)$ and that $1\in K_{\alpha}$ if and only if $\alpha/2\in\mathbb{N}$. We define
 $\binom{w}{z}:=\frac{\Gamma(w+1)}{\Gamma(z+1)\Gamma(w-z+1)}$
@@
 Also, $0^{0}$ is defined as 1.
-The equality holds in (1.1) if $\alpha=1$, which is nothing but the usual binomial theorem.
-Lyons also conjectured that the coefficient $\alpha^{2}$ in (1.1) can be replaced by $\alpha$.
-Afterwards, Hara and Hino [HH10] gave an affirmative answer to Lyons’ conjecture.
-Proposition 1.1 (Neo-classical inequality, [HH10, Theorem 1.2]).
-Let $\alpha\in(0,1]$. For $n\in\mathbb{N}$ and $x,y\geq 0$, we have
-$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}.$ (1.2)
-The equality holds if and only if $\alpha=1$ or $x=y=0$.
-When $\alpha\in(0,1)$ and only one of $x$ and $y$ is zero, the inequality (1.2) is trivial with strict inequality.
-Therefore, in order to show (1.2), it suffices to consider the case where $\alpha\in(0,1)$ and $x,y>0$.
-Moreover, we may assume that $x\leq y$ by the symmetry of the roles of $x$ and $y$ in (1.2).
-By dividing both sides by $y^{\alpha n}$ and letting $\lambda=x/y$, (1.2) is equivalent to
-$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}\leq(1+\lambda)^{\alpha n},\qquad\alpha\in(0,1),\,n\in\mathbb{N},\,\lambda\in(0,1].$ (1.3)
-For proving (1.3), Hara and Hino established a generalization of the binomial theorem [HH10, Theorem 1.3], which is stated roughly as follows:
-$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}=(1+\lambda)^{\alpha n}+R_{\alpha}^{(n)}(\lambda),\qquad\alpha\in(0,2),\,n\in\mathbb{N},\,\lambda\in(0,1],$ (1.4)
-where the remainder term $R_{\alpha}^{(n)}(\lambda)$ is given explicitly.
-Indeed, by the expression, the term $R_{\alpha}^{(n)}(\lambda)$ is strictly negative when $\alpha\in(0,1)$, which readily implies (1.3).
-Furthermore, they extended the generalized binomial theorem (1.4) to the case where $\alpha$ is an arbitrary positive number.
-Let us now give the precise statement.
-In general, for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq 0\}$ and $\gamma\in\mathbb{R}$, the power $z^{\gamma}$ is defined as $\exp(\gamma\,\mathrm{Log}\,z)$, where $\mathrm{Log}\,z$ is the principal value of $\log z$ so that Log $\mathrm{Log}\,1=0$.
-For $\alpha>0$, we put
-$K_{\alpha}:=\{e^{i\theta}\in\mathbb{C}\mid-\pi<\theta\leq\pi,\,e^{i\theta\alpha}=1\}.$ (1.5)
-We note that $K_{\alpha}=\{1\}$ when $\alpha\in(0,2)$ and that $-1\in K_{\alpha}$ if and only if $\alpha/2\in\mathbb{N}$. Then, the generalized binomial theorem is stated as follows.
-Proposition 1.2 ([HH10, Theorem 3.2]).
+The generalized binomial theorem is stated as follows.
+Proposition 1.1 ([12, Theorem 3.2]).
 For $\alpha>0$, $n\in\mathbb{N}$, and $\lambda\in(0,1]$, we have
 $\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\left\{\frac{1}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{\lambda^{\alpha n}}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.6)
+$\quad\times\left\{\frac{1}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{\lambda^{\alpha n}}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.2)
 Here, the second term on the right-hand side is regarded as zero if $\alpha/2\in\mathbb{N}$.
-It should be noted that the first term on the right-hand side of (1.2) is a real number because $\omega\in K_{\alpha}$ implies $\overline{\omega}\in K_{\alpha}$.
-For proving Proposition 1.2, the authors provided in [HH10, Theorem 3.1] the fractional-order Taylor-like series expansions with residual terms by using basic complex analysis.
+We note that Proposition 1.1 also holds for $\lambda>0$ by analytic continuation in the parameter $\lambda$.
+For proving Proposition 1.1, the authors established the fractional-order Taylor-like series expansions with residual terms in [12, Theorem 3.1], which has the fractional calculus as its theoretical background.
 See Proposition 2.1 for further details.
-1.2. Fractional binomial distribution
-Suppose that $n\in\mathbb{N}$ and $x\in[0,1]$. The binomial distribution $\mu_{x}^{(n)}$ defined as
-$\mu_{x}^{(n)}(\{j\})=\binom{n}{j}x^{j}(1-x)^{n-j},\qquad j=0,1,2,\dots,n$
-on $\{0,1,2,\dots,n\}$ is one of the most important discrete probability distributions.
-The equality $\sum_{j=0}^{n}\mu_{x}^{(n)}(\{j\})=1$ is derived from the usual binomial theorem.
-Because Proposition 1.2 is regarded as a generalization of the binomial theorem, it is quite natural to ask for a fractional analogue of the binomial distribution in accordance with Proposition 1.2.
-For $\alpha>0$, $n\in\mathbb{N}$ and $x\in(0,1)$, we have
+Incidentally, Proposition 1.1 is used to answer Lyons’ conjecture relating to the neo-classical inequality, which plays a crucial role in the proof of a fundamental theorem of rough path theory (see [22, Theorem 2.2.1]).
+Let $\alpha>0$ and $n\in\mathbb{N}$. By putting $\lambda=x/(1-x)$, $x\in(0,1)$, in (1.1) and multiplying both sides by $(1-x)^{\alpha n}$, we have
 $Z_{\alpha,x}^{(n)}:\!$ $=\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$
 $=1+\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}-\frac{\alpha x^{\alpha}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\left\{\frac{(1-x)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{x^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\right\}\,\mathrm{d}t$ (1.7)
-by putting $\lambda=x/(1-x)$ in (1.2) and multiplying both sides by $(1-x)^{\alpha n}$.
-Keeping in mind that the right-hand side of (1.7) is close to 1 when $n$ is sufficiently large, this identity gives rise to a discrete probability distribution on $\{0,1,2,\dots,n\}$ as follows.
+$\quad\times\left\{\frac{(1-x)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{x^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.3)
+Keeping in mind that the right-hand side of (1.3) is close to 1 when $n$ is sufficiently large, we undertake a systematic study of the fractional binomial distribution and its associated fractional Bernstein operator, which are defined as follows.
 Definition 1.2 (Fractional binomial distribution).
-Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Then, a probability measure $\mu_{\alpha,x}^{(n)}$ on $\{0,1,2,\dots,n\}$, called the $\alpha$-fractional binomial distribution, is defined as
-$\mu_{\alpha,x}^{(n)}(\{j\})=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)},\qquad j=0,1,2,\dots,n.$
-Moreover, we define $Z_{\alpha,0}^{(n)}=Z_{\alpha,1}^{(n)}=\alpha$, $\mu_{\alpha,0}^{(n)}=\delta_{0}$, and $\mu_{\alpha,1}^{(n)}=\delta_{1}$, where $\delta_{x}$ is the Dirac measure at $x$.
-It is clear that $\mu_{1,x}^{(n)}$ coincides with the binomial distribution $\mu_{x}^{(n)}$.
-The probability distribution $\mu_{\alpha,x}^{(n)}$ is regarded as a fractional variant of $\mu_{x}^{(n)}$.
-As seen in Corollary 2.8, one of the features is that the mean of $\mu_{\alpha,x}^{(n)}$ is asymptotically equal to that of $\mu_{x}^{(n)}$ as $n\to\infty$, while the variance of $\mu_{\alpha,x}^{(n)}$ is asymptotically $1/\alpha$ times that of $\mu_{x}^{(n)}$.
-One of the aims of this study is to study some basic properties of the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, and all such properties are given with their proofs in Section 2.
-In Section 2.1, we review the fractional-order Taylor-like series of holomorphic functions $f$ with residual terms [HH10, Theorem 2.1].
-See Proposition 2.1 for its precise statement.
-In Section 2.2, we provide explicit expressions for moments of $\mu_{\alpha,x}^{(n)}$ by making effective use of the fractional-order Taylor-like series expansion.
-See Theorem 2.5 and Corollary 2.8.
-In particular, we see that the ($c$-)Whitney numbers of the second kind—which in combinatorics are a certain generalization of the Stirling numbers of the second kind—appear in the expression for each moment of $\mu_{\alpha,x}^{(n)}$.
-Section 2.3 discusses a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to both $n\in\mathbb{N}$ and $x\in[0,1]$.
-Furthermore, in Section 2.4, a uniform estimate for the moments of $\mu_{\alpha,x}^{(n)}$ is also obtained.
-Section 2.5 focuses on the characteristic function $\varphi_{\alpha,x}^{(n)}(\xi)$ of $\mu_{\alpha,x}^{(n)}$, which plays a crucial role in obtaining the central limit theorem for $\mu_{\alpha,x}^{(n)}$.
-Lastly, in Section 2.6, a multivariate version of the neo-classical inequality is deduced.
-Similarly to Definition 1.3, this inequality also induces a fractional analogue of the multinomial distribution.
-Being different from the binomial case, however, it seems quite difficult to obtain even basic characteristics of the fractional multinomial distribution such as its mean vector and its variance-covariance matrix, because the multivariate version of the fractional Taylor series expansion is yet to be obtained.
-The discussion of the multivariate cases is left to the future.
-Here, note that the terminology fractional in probability theory is used in several different contexts.
+Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. Then, a probability measure $\mu_{\alpha,x}^{(n)}$ defined as
+$\mu_{\alpha,x}^{(n)}(\mathrm{d}z):=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\delta_{j}(\mathrm{d}z)$
+is called the $\alpha$-fractional binomial distribution.
+Note that $Z_{\alpha,0}^{(n)}=Z_{\alpha,1}^{(n)}=\alpha$, $\mu_{\alpha,0}^{(n)}=\delta_{0}$, and $\mu_{\alpha,1}^{(n)}=\delta_{1}$.
+Definition 1.3 (Fractional Bernstein operator).
+Let $\alpha>0$ and $n\in\mathbb{N}$. The $\alpha$-fractional Bernstein operator $B_{\alpha,n}$, acting on $C([0,1])$, is defined by
+$B_{\alpha,n}f(x):\!$ $=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{\alpha,x}^{(n)}(\mathrm{d}z)$
+$=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$
+It is clear that $\mu_{1,x}^{(n)}$ and $B_{1,n}$ coincide with $\mu_{x}^{(n)}$ and $B_{n}$, respectively.
+In this sense, the fractional binomial distribution (resp.
+fractional Bernstein operator) is a natural generalization of the binomial distribution (resp.
+Bernstein operator) with additional parameter $\alpha$.
+This family interpolates smoothly between the classical and the newly introduced fractional structures; the probability mass function of $\mu_{\alpha,x}^{(n)}$ depends smoothly on both $\alpha$ and $x$.
+The purpose of this paper is to develop a fractional framework that encompasses both $\mu_{\alpha,x}^{(n)}$ and $B_{\alpha,n}$ in a unified manner based on the generalized binomial theorem.
+Although many of the resulting statements may appear elementary at first glance, their proofs often require techniques quite different from those used in the classical setting.
+These differences arise because, in general, $\mu_{\alpha,x}^{(n)}$ is not expressed as the law of the sum of i.i.d.
+random variables.
+Our main contribution is to clarify, in a systematic and unified way, the probabilistic and analytic features of $\mu_{\alpha,x}^{(n)}$ and $B_{\alpha,n}$.
+To the best of our knowledge, no previous work has constructed fractional analogues of the binomial distribution and the Bernstein operator in the unified way developed here.
+1.2 Outline of the paper
+The remainder of the paper is organized as follows.
+The first part (Section 2). We establish explicit expressions and asymptotics for the moments of $\mu_{\alpha,x}^{(n)}$. In particular, we have
+$\mathbb{E}[S_{\alpha,x}^{(n)}]=nx+O(e^{-\delta n})\quad\text{and}\quad\mathrm{Var}(S_{\alpha,x}^{(n)})=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$ (1.4)
+as $n\to\infty$ for some $\delta>0$, where $S_{\alpha,x}^{(n)}$ is a random variable obeying $\mu_{\alpha,x}^{(n)}$.
+This reveals that the variance scales inversely with $\alpha$.
+See Figure 1.
+We also obtain a uniform bound of $Z_{\alpha,x}^{(n)}$ with respect to both $n\in\mathbb{N}$ and $x\in[0,1]$, and an explicit expression for the characteristic function of $\mu_{\alpha,x}^{(n)}$ with its quantitative estimate.
+Figure 1: Variation of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ with respect to $\alpha$.
+Shown are the probability mass functions for $n=20$ and $x=0.35$, where the parameter $\alpha$ takes the values $0.2$, $0.6$, 11, $1.4$, $1.8$, and $2.2$.
+The second part (Section 3).
+Based on the properties given in Section 2, we prove several limit theorems: the weak law of large numbers, the central limit theorem, and the law of small numbers for $\mu_{\alpha,x}^{(n)}$.
+In particular, the limiting distribution in the latter case is a fractional variant of the Poisson distribution defined in terms of the Mittag-Leffler function (see e.g., [13, 6]).
+The third part (Section 4).
+We investigate several properties of the fractional Bernstein operator $B_{\alpha,n}$ and its iterates.
+We prove the uniform approximation result and the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$.
+We also prove the convergence of the appropriately scaled operators $(B_{\alpha,n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, to the Wright–Fisher diffusion semigroup with generator $\mathcal{L}_{\alpha}f(x)=(2\alpha)^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$ as $n\to\infty$ in some sense.
+Furthermore, we extend this convergence to the case where the fractional order is a continuous function $\alpha(x)\colon[0,1]\to(0,\infty)$, and the limiting generator is given by $\mathcal{L}_{\alpha(\cdot)}f(x)=(2\alpha(x))^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$, which leads to more-general diffusion processes on $[0,1]$.
+1.3 Future perspectives
+We systematically generalize known results for the classical binomial distribution and Bernstein operator into their fractional counterparts by means of a unified framework based on the generalized binomial theorem.
+This approach enables us to make substantial contributions to the study of probability distributions and related fields.
+We expect that this framework will shed light on the analytical and geometric aspects of such structures, and deepen the understanding of various mathematical models, including stochastic processes, applied statistics, and information geometry.
+For example, the fractional binomial distributions may provide flexible models for count data which can describe both overdispersion and underdispersion, in view of (1.4).
+Furthermore, the study of the geometric properties of the statistical manifold formed by these distributions may lead to new developments in statistical theory and its applications.
+1.4 Remarks
+The terminology fractional in probability theory is used in several different contexts.
 For instance, a typical use can be seen in the notion of fractional Brownian motion.
-See [Lee21] (for example) for the fractional binomial distribution in this sense.
+See [20] for the fractional binomial distribution in this sense.
 Yet another use is derived from fractional calculus.
-We refer to [CP12] for the fractional binomial process defined in terms of Caputo’s fractional derivatives.
-However, to the best of our knowledge, no paper to date has discussed the fractional binomial distribution in our sense, which means that the present paper gives a new insight into the study of probability theory and other areas of mathematics.
-Based on the properties given in Section 2, we give analogues of three classical limit theorems in Section 3.
-Sections 3.1 and 3.2 are devoted to the proofs of the weak law of large numbers and the central limit theorem for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
-We also obtain the law of small numbers for $\mu_{\alpha,x}^{(n)}$ in Section 3.3.
-Namely, we show that $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to a certain probability distribution with masses on $\{0,1,2,\dots\}$ as $n\to\infty$ together with $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ for some $\alpha>0$ and $\lambda>0$.
-The limiting probability distribution $\nu_{\alpha,\lambda}$ is regarded as a fractional analogue of the Poisson distribution.
-See Definition 3.3 for the precise definition.
-Obviously, $\mu_{1,\lambda}$ coincides with the usual Poisson distribution.
-Note that $\nu_{\alpha,\lambda}$ is defined in terms of a special function called the Mittag-Leffler function $E_{\alpha}(z)$, $\alpha>0$, which is a generalization of the exponential function.
-The distribution $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions discussed in [Her16, CO17].
-We also mention that the Mittag-Leffler function has already been used to introduce some probability distributions in different situations.
-We refer to [Pil90, PJ95] (for example) for the notion of the Mittag-Leffler distribution, although it is far from the Poisson distribution.
-A few basic properties of the fractional Poisson distribution such as a recurrence formula for its moments are discussed in Section 3.4 as well.
-Because the Mittag-Leffler function is used frequently in fractional calculus, further investigations of $\nu_{\alpha,\lambda}$ from the point of view of fractional calculus are expected.
-1.3. Fractional Bernstein operator
-There exist various applications of the binomial distribution to other branches of mathematics.
-One of them is seen in the study of positive linear operators acting on $C([0,1])$, the Banach space of all continuous functions on $[0,1]$ endowed with the supremum norm $\|\cdot\|_{\infty}$.
-Bernstein introduced in his short paper [Ber12] the following positive linear operator acting on $C([0,1])$ defined in terms of the binomial distribution.
-Definition 1.4 (Bernstein operator, cf. [Ber12]).
-For $n\in\mathbb{N}$, a positive linear operator $B_{n}$ acting on $C([0,1])$ is defined by
-$B_{n}f(x)=\sum^{n}_{k=0}\binom{n}{k}x^{k}(1-x)^{n-k}f\left(\frac{k}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$ (1.8)
-The positive linear operator $B_{n}$ is called the Bernstein operator.
-The right-hand side of (1.8) is called the Bernstein polynomial. Bernstein proved in [Ber12] that $B_{n}$ uniformly approximates every continuous function $f\in C([0,1])$.
-Proposition 1.5 (cf. [Ber12]).
-For any $f\in C([0,1])$, we have
-$\lim_{n\to\infty}\|B_{n}f-f\|_{\infty}=0.$
-It is clear that Proposition 1.5 gives a constructive proof of the celebrated Weierstrass approximation theorem.
-It is also known that the Bernstein operator has a number of properties of interest.
-We refer to [Bus17] for comprehensive information on the Bernstein operator with extensive references therein, and to [Alt10, AC94] for surveys on approximating properties of positive linear operators acting on some function spaces, including the Bernstein operator.
-On the other hand, it is natural to consider limiting behaviors of $B_{n}^{k}$, the $k$-times iterations of $B_{n}$, as $k\to\infty$.
-One such limiting behavior was first investigated in [KR67].
-Proposition 1.6 (cf. [KR67, Theorem 1]).
-Let $n\in\mathbb{N}$ be fixed. Then, we have
-$\lim_{k\to\infty}\max_{x\in[0,1]}|B_{n}^{k}f(x)-\{f(1)x+f(0)(1-x)\}|=0,\qquad f\in C([0,1]).$
-Since the publication of [KR67], several authors have tried to make the proof of Proposition 1.6 more sophisticated.
-We refer to [AI09, Rus04] (for example) for its short proofs.
-Moreover, Konstantopoulos, Yuan and Zazanis [KYZ18] discussed a limit theorem for iterates of $B_{n}$ with scales different from Proposition 1.6.
-In particular, they gave a probabilistic proof of the following limit theorem.
-Proposition 1.7 (cf. [KYZ18, Theorem 3]).
-For $f\in C([0,1])$ and $t\geq 0$, we have
-$\lim_{n\to\infty}\|B_{n}^{\lfloor nt\rfloor}f-\mathbb{E}[f(X_{\bullet}(t))]\|_{\infty}=0,$
-where $(X_{x}(t))_{t\geq 0}$ is the Wright–Fisher diffusion process which solves
-$\mathrm{d}X_{x}(t)=\sqrt{X_{x}(t)(1-X_{x}(t))}\,\mathrm{d}W(t),\qquad X_{x}(0)=x\in[0,1].$
-Here, $(W(t))_{t\geq 0}$ is standard one-dimensional Brownian motion.
-This theorem asserts that the sequence of discrete semigroups $(B_{n}^{\lfloor nt\rfloor})_{t\geq 0}$, $n\in\mathbb{N}$ converges to the Wright–Fisher diffusion semigroup as $n\to\infty$.
-In other words, the properly scaled Markov chain $X_{x}^{(n)}(\lfloor nt\rfloor)$ whose one-step transition operator is given by $B_{n}$ converges in law to $X_{x}(t)$ as $n\to\infty$ for each $t\geq 0$.
-We also refer to [KZ70] for early works on limit theorems for iterates of some positive linear operators.
-For further details on the Wright–Fisher diffusion, see [EK86, Section 10] (for example).
-In Section 4, we introduce a fractional analogue of the Bernstein operator with the binomial distribution replaced by the fractional one defined in Definition 1.3, and we investigate some of its properties including limit theorems for its iterates.
-The following is the definition of the fractional Bernstein operator.
-Definition 1.8 (Fractional Bernstein operator).
-Let $\alpha>0$ and $n\in\mathbb{N}$. The $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ is defined by
-$B_{\alpha,n}f(x)$ $:=\sum_{j=0}^{n}\mu_{\alpha,x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
-$=\sum_{j=0}^{n}\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$
-We note that $B_{1,n}$ coincides with the Bernstein operator $B_{n}$.
-In Section 4.1, we show in Theorem 4.1 that the fractional Bernstein operator $B_{\alpha,n}$ also uniformly approximates every continuous function in $C([0,1])$ similarly to Proposition 1.5.
-Moreover, Sections 4.2 and 4.3 discuss limiting behaviors of the iterates $B_{\alpha,n}^{k}f$, $f\in C([0,1])$ as $k\to\infty$.
-In the former section, we obtain the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$.
-In the latter section, we concentrate on a fractional analogue of Proposition 1.7.
-We first show that the sequence $\{n(B_{\alpha,n}-I)\}_{n=1}^{\infty}$ converges to a second-order differential operator $\mathcal{L}_{\alpha}=(2\alpha)^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$ with Neumann boundary condition.
-This means the convergence of the infinitesimal generator of the discrete Markov semigroup generated by $B_{\alpha,n}$ to that of a diffusion semigroup.
-Note that $\mathcal{L}_{\alpha}$ is the generator of the Wright–Fisher diffusion with diffusion coefficient different from that of $(X_{x}(t))_{t\geq 0}$ up to a multiple constant.
-Next, we show the uniform convergence of the iterates $B_{\alpha,n}^{\lfloor nt\rfloor}f$, $f\in C([0,1])$ to the Wright–Fisher semigroup corresponding to $\mathcal{L}_{\alpha}$ as $n\to\infty$, by employing the celebrated Trotter approximation theorem (cf.
-[Tro58]).
-See Lemma 4.5 and Theorem 4.4 for precise statements.
-We also discuss in Section 4.4 that Theorem 4.4 holds even though we replace the parameter $\alpha>0$ with a continuous function $\alpha(x)\colon[0,1]\to(0,\infty)$.
-In this case, the corresponding infinitesimal generator is given by $\mathcal{L}_{\alpha(\cdot)}=(2\alpha(x))^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$.
-Hence, it turns out that the stochastic processes captured by the limit of $B_{\alpha(\cdot),n}^{\lfloor nt\rfloor}$ as $n\to\infty$ are not only Wright–Fisher diffusions but also more-general diffusions on $[0,1]$.
-Here, we mention some existing notions of linear operators named the fractional Bernstein operator.
-In [Man15], the author introduced the fractional Bernstein operator $(1-B_{n})^{\alpha}$, $\alpha>0$ in order to apply density estimations of probability distributions in statistics and data science, which differs completely from ours.
-Another different formulation of the fractional Bernstein operator from ours can be found in [Yu21].
-Our fractional Bernstein operator $B_{\alpha,n}$ seems new in that we define it via the generalized binomial theorem deriving the neo-classical inequality.
-We expect to develop further studies of $B_{\alpha,n}$ in the near future.
-1.4. Notation
+We refer to [5] for the fractional binomial process defined in terms of Caputo’s fractional derivatives.
+We also mention some existing notions of linear operators named the fractional Bernstein operator.
+In [24], the author studies the fractional Bernstein operator $(1-B_{n})^{\alpha}$, $\alpha>0$, which differs completely from ours.
+Another different formulation of the fractional Bernstein operator from ours can be found in [31].
+Our fractional Bernstein operator $B_{\alpha,n}$ appears to be new, being directly derived from the generalized binomial theorem.
+1.5 Notation
 We use the following notation.
-  • We set =$\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
+  • We set $\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
   • $\mathrm{Re}(\omega)$ denotes the real part of a complex number $\omega$.
   • The indicator function of a set $A$ is denoted by $\bm{1}_{A}$.
-  • The cardinality of a set $A$ is denoted by #$\#(A)$.
+  • The cardinality of a set $A$ is denoted by #(A)\#(A).
   • For $x,y\in\mathbb{R}$, we write $x\vee y=\max\{x,y\}$ and $x\wedge y=\min\{x,y\}$.
-  • For $x\in\mathbb{R}$, we denote by $\lfloor x\rfloor$ and $\lceil x\rceil$ the greatest integer less than or equal to $x$, and the smallest integer greater than or equal to $x$, respectively.
-  • For $a\in\mathbb{R}$, we denote by $\mathrm{sgn}\,a$ the sign of $a$, that is,
+  • For $\in\mathbb{R}, x\lfloor x\rfloor$ and $\lceil x\rceil$ denote the greatest integer less than or equal to $x$, and the smallest integer greater than or equal to $x$, respectively.
+  • For $a\in\mathbb{R}$, $\mathrm{sgn}\,a$ denotes the sign of $a$, that is,
     $\mathrm{sgn}\,a=\begin{cases}1&\text{if }a>0,\\ 0&\text{if }a=0,\\ -1&\text{if }a<0.\end{cases}$
@@
   • Let $X$ and $Y$ be two Banach spaces. The domain and range of a linear operator $L\colon X\to Y$ are denoted by $\mathrm{Dom}(L)\subset X$ and $\mathrm{Range}(L)\subset Y$, respectively.
-  • A positive constant whose concrete expression is not important is denoted by $C_{i,j}$ when it is the $j$-th positive constant in Section $i$
-2. The fractional binomial distribution and its properties
-2.1. A fractional-order Taylor series
+  • A positive constant whose concrete expression is not important is denoted by $C_{i,j}$ when it is the $j$-th positive constant in Section $i$ .
+2 The fractional binomial distribution and its properties
+2.1 A fractional-order Taylor series
 Let $D=\{z\in\mathbb{C}\,:\,|z|<1\}$ be the unit open disc and $\overline{D}$ its closure in $\mathbb{C}$.
@@
 The following is the fractional-order Taylor series with residual terms. Recall that the set $K_{\alpha}$ was defined in (1.1).
-Proposition 2.1 (cf. [HH10, Theorem 3.1]).
+Proposition 2.1 (cf. [12, Theorem 3.1]).
 Suppose that $\alpha>0$ and $\lambda\in(0,1)$. If $\alpha/2\notin\mathbb{N}$, then we have
@@
 $\alpha\sum_{j=-\infty}^{-1}f^{\#}(\alpha j)\lambda^{-\alpha j}$ $=\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}f(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.2)
-If $\alpha/2\in\mathbb{N}$, (2.1) and (2.2) still hold by regarding the second term of the right-hand side of (2.1) and the right-hand side of (2.2) as 00.
+If $\alpha/2\in\mathbb{N}$, (2.1) and (2.2) still hold by regarding the second term of the right-hand side of (2.1) and the right-hand side of (2.2) as 0.
 We note that these two equalities in Proposition 2.1 turn out to be true when $\lambda=1$ through the limit as $\lambda\nearrow 1$, provided that
@@
 We also recall the following result.
-Proposition 2.2 (cf. [HH10, Proposition 2.4]).
+Proposition 2.2 (cf. [12, Proposition 2.4]).
 Let $T>0$ and define $f(z)=(1+z)^{T}$ on $\overline{D}$. Then, we have
 $f^{\#}(\xi)=\binom{T}{\xi},\qquad\xi\in\mathbb{R}.$
-2.2. Moments of the fractional binomial distribution
-Suppose that $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$.
-We aim to find concrete expressions of moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ by applying the fractional Taylor series (2.1) and (2.2) to a certain holomorphic function.
-Let $c\in\mathbb{R}$ and define a differential operator $\mathscr{D}_{c}$ acting on holomorphic functions $f$ on a domain of $\mathbb{C}$ by
+2.2 Moments of the fractional binomial distribution
+Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$.
+We now derive the explicit expressions for the moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ by applying Proposition 2.1 to a certain holomorphic function.
+For $c\in\mathbb{R}$, we define a differential operator $\mathscr{D}_{c}$ acting on holomorphic functions $f$ on a domain of $\mathbb{C}$ by
 $\mathscr{D}_{c}f(z):=zf^{\prime}(z)+cf(z).$
@@
 for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq-1\}$ if $\alpha n\notin\mathbb{N}$ and $z\in\mathbb{C}$ if $\alpha n\in\mathbb{N}$.
-Moreover, each $W_{c}(m,k)$ is a polynomial in $c$ of degree $m-k$ with integer coefficients (see also Table 1).
+Moreover, each $W_{c}(m,k)$ is a polynomial in $c$ of degree mkm-k with integer coefficients.
+See Table 1 for concrete expressions of some of $W_{c}(m,k)$.
+Table 1: Expressions of $W_{c}(m,k)$ for $m=0,1,2,3,4$.
 $m$ $k$ 0 1 2 3 4
@@
 1 $c$ 1
-2 $c^{2}$ $12c+$ 1 1
-3 $c^{3}$ $13c^{2}+3c+$ 1 $33c+$ 3 1
-4 $c^{4}$ $14c^{3}+6c^{2}+4c+$ 1 $76c^{2}+12c+$ 7 $64c+$ 6 1
-Table 1. Values of $W_{c}(m,k)$ for $m=0,1,2,3,4$.
+2 $c^{2}$ $2c+1$ 1
+3 $c^{3}$ $3c^{2}+3c+1$ $3c+3$ 1
+4 $c^{4}$ $4c^{3}+6c^{2}+4c+1$ $6c^{2}+12c+7$ $4c+6$ 1
 Proof.
@@
 $=z\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\{kz^{k-1}(1+z)^{\alpha n-k}+(\alpha n-k)z^{k}(1+z)^{\alpha n-k-1}\}$
-$\hskip 28.45274pt+c\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
+$\quad+c\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
 $=\sum_{k=1}^{m}\{(k+c)W_{c}(m,k)+W_{c}(m,k-1)\}(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
-$\hskip 28.45274pt+W_{c}(m,m)(\alpha n)_{m+1}z^{m+1}(1+z)^{\alpha n-m-1}+cW_{c}(m,0)(1+z)^{\alpha n}.$
+$\quad+W_{c}(m,m)(\alpha n)_{m+1}z^{m+1}(1+z)^{\alpha n-m-1}+cW_{c}(m,0)(1+z)^{\alpha n}.$
 Therefore, it follows from (2.4) that the right-hand side coincides with
@@
 Therefore, the sequence $\{W_{c}(m,k)\}$ is regarded as an extension of the Stirling numbers of the second kind, which in combinatorics are referred to as the ($c$-)Whitney numbers of the second kind.
-To the best of our knowledge, this notion was first introduced in [Mez10], and indeed the theorem below tells us that $W_{c}(m,k)$ appears when we express the moments of the fractional binomial distribution.
+To the best of our knowledge, this notion was first introduced in [25], and indeed the theorem below tells us that $W_{c}(m,k)$ appears when we express the moments of the fractional binomial distribution.
 By applying Lemma 2.3, we obtain the following key identity to deduce the moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ in Definition 1.2.
@@
 $=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}x^{k}+\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}\left\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right\}$
-$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
-$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.6)
-Here, if $\alpha\in\mathbb{N}$, then the third and fourth terms of the right-hand side are regarded as 00.
+$\quad-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
+$\quad-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.6)
+Here, if $\alpha\in\mathbb{N}$, then the third and fourth terms of the right-hand side are regarded as 0.
 Proof.
 The proof is split into two parts.
-Step 1. Let $f$ be a continuous function on $\overline{D}$ such that $f$ is holomorphic on $D$ and $f(-1)=0$. Suppose that a function $g$ defined by
+Step 1. Let $f$ be a continuous function on $\overline{D}$ such that $f$ is holomorphic on $D$ and f(1)=0f(-1)=0. Suppose that a function $g$ defined by
 $g(z):=\mathscr{D}_{c}f(z)=zf^{\prime}(z)+cf(z),\qquad z\in D$
-is extended as a continuous function on $\overline{D}$ and the function $(-1/2,1/2)\ni x\longmapsto g(e^{2\pi ix})$ is of class $C^{1}$. Then, we have
+is extended as a continuous function on $\overline{D}$ and the function $(-1/2,1/2)\ni x\longmapsto f(e^{2\pi ix})$ is of class $C^{1}$. Then, we have
 $g^{\#}(\xi)$ $=\int_{-1/2}^{1/2}e^{2\pi ix}f^{\prime}(e^{2\pi ix})e^{-2\pi ix\xi}\,\mathrm{d}x+cf^{\#}(\xi)$
@@
 $=(\xi+c)f^{\#}(\xi),\qquad\xi\in\mathbb{R}.$ (2.7)
-In the last equality, we used $f(-1)=0$.
+In the last equality, we used f(1)=0f(-1)=0.
 Because $m\leq\alpha n$ or $\alpha n\in\mathbb{N}$ holds, we can take the function $h_{c}^{(k)}$, $k=0,1,\dots,m-1$ defined by (2.3) as $f$.
@@
 $(h_{c}^{(m)})^{\#}(\xi)=(\xi+c)(h_{c}^{(m-1)})^{\#}(\xi)=\cdots=(\xi+c)^{m}(h_{c}^{(0)})^{\#}(\xi)=(\xi+c)^{m}\binom{\alpha n}{\xi}$ (2.8)
-for $\xi\in\mathbb{R}$, from (2.7) and Proposition 2.2.
+for $\xi\in\mathbb{R}$, from (2.7) and Proposition $2.$
 Step 2. We apply Proposition 2.1 to the function $f(z)=h_{c}^{(m)}(z)$. Then, it follows from (2.8) that
@@
 $=(-1)^{m}\lambda^{\alpha n}\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t$
-by using (2.10) with $c$ replaced by $-(c+\alpha n)$, we apply it to (2.9) to obtain
+by using (2.10) with $c$ replaced by $(c+\alpha n)$, we apply it to (2.9) to obtain
 $\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$ $=\alpha\sum_{j=0}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}-\alpha\sum_{j=n+1}^{\infty}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}$
 $=\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}(\lambda\omega)-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
-$\hskip 28.45274pt-(-1)^{m}\frac{\alpha\lambda^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.11)
-We put $\lambda=x/(1-x)$, $20<x<1/$ 2 and multiply both sides in (2.11) by $(1-x)^{\alpha n}$. Then, we obtain
+$\quad-(-1)^{m}\frac{\alpha\lambda^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.11)
+We put $\lambda=x/(1-x)$, $0<x<1/2$ and multiply both sides in (2.11) by $(1-x)^{\alpha n}$. Then, we obtain
 $\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$
 $=(1-x)^{\alpha n}\sum_{\omega\in K_{\alpha}}h_{c}^{(m)}\left(\frac{x\omega}{1-x}\right)$
-$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
-$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.12)
+$\quad-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
+$\quad-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.12)
 Because
@@
 $|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}=\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}\}$
-are extended to holomorphic functions in $x$ on a complex domain containing the set $\{x\in\mathbb{R}\mid 0<x<1\}$ and do not vanish if $\alpha\notin\mathbb{N}$, (2.12) also holds for $10<x<$ 1 by analytic continuation.
+are extended to holomorphic functions in $x$ on a complex domain containing the set $\{x\in\mathbb{R}\mid 0<x<1\}$ and do not vanish if $\alpha\notin\mathbb{N}$, (2.12) also holds for $0<x<1$ by the identity theorem.
 Moreover, by applying (2.5) with $z=x\omega/(1-x)$ and $\omega\in K_{\alpha}$ to the first term of the right-hand side of (2.12), we have
@@
 We remark that if $\omega=-1\in K_{\alpha}$, then $\alpha/2\in\mathbb{N}$ so that $h_{c}^{(m)}(x\omega/(1-x))$ is well-defined. Thus, (2.12) implies (2.6). ∎
-As seen from (1.7), the normalized constant $Z_{\alpha,x}^{(n)}$ satisfies
-$Z_{\alpha,x}^{(n)}=1+O(e^{-\delta n}),\qquad\frac{1}{Z_{\alpha,x}^{(n)}}=1+O(e^{-\delta n})$ (2.13)
-as $n\to\infty$ for some $\delta>0$, where $\delta$ can be chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\alpha\in(0,\infty)$.
-In what follows, we use the notation $O(e^{-\delta n})$ in this sense.
-Then, we also have the following.
 Proposition 2.6.
-For each $m\in\mathbb{N}_{0}$, there exists some $\delta>0$ such that
+  1. (1)
+    As $n\to\infty$,
+    $Z_{\alpha,x}^{(n)}=1+O(e^{-\delta n}),\qquad\frac{1}{Z_{\alpha,x}^{(n)}}=1+O(e^{-\delta n}).$ (2.13)
+  2. (2)
+    For each $m\in\mathbb{N}_{0}$,
     $\alpha\sum_{j=0}^{n}(\alpha j+c)^{m}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}=\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}x^{k}+(|c|^{m}+1)O(e^{-\delta n})$ (2.14)
     as $n\to\infty$.
+Here, the symbol $O(e^{-\delta n})$ means that there exist positive constants $C$ and $\delta$ depending only on $x$, $\alpha$ (and $m$) that are chosen uniformly on any compact subset of $x\in(0,1)$ and any compact subset of $\alpha\in(0,\infty)$ such that the modulus of the corresponding term is dominated by $Ce^{-\delta n}$.
+In what follows, we use the notation $O(e^{-\delta n})$ in this sense; $\delta$ may change from line to line.
 In order to prove Proposition 2.6, we need to give the following estimates of the third term on the right-hand side of (1.3).
@@
 Proof.
-Because the left-hand sides of (2.15) and (2.16) are regarded as 00 if $\alpha\in\mathbb{N}$, we may suppose that $\alpha$ is not an integer.
-It is sufficient to show (2.15) because (2.16) is immediately obtained by replacing $x$ in (2.15) with $1-x$.
+Because the left-hand sides of (2.15) and (2.16) are regarded as 0 if $\alpha\in\mathbb{N}$, we may suppose that $\alpha$ is not an integer.
+It is sufficient to show (2.15) because (2.16) is immediately obtained by replacing $x$ in (2.15) with 1x1-x.
 We have
@@
 $\leq\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}}{\pi}|\sin\alpha\pi|\int_{0}^{1}\frac{t^{\alpha-1}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$ (2.17)
-By the change of variable $s=((1-x)t/x)^{\alpha}$ and [GR15, Equation 12, page 329], the right-hand side of (2.17) becomes
+By the change of variable $s=((1-x)t/x)^{\alpha}$ and [11, Equation 12, page 329], the right-hand side of (2.17) becomes
 $\frac{(1-x)^{\alpha n}}{\pi}|\sin\alpha\pi|\int_{0}^{(\frac{1-x}{x})^{\alpha}}\frac{\mathrm{d}s}{|s-e^{-i\alpha\pi}|^{2}}$
@@
 $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}=O(e^{-\delta n})$ (2.19)
-as $n\to\infty$ for some $\delta>0$. Furthermore, Lemma 2.3 implies that
+as $n\to\infty$ for some $\delta>0$. (2.13) follows from (1.3), (2.19) and Lemma 2.7. Furthermore, Lemma 2.3 implies that
 $|W_{c}(m,k)|\leq C_{2,1}(|c|^{m-k}+1),\qquad m\in\mathbb{N}_{0},\,\,k=0,1,\dots,m$ (2.20)
@@
 $\leq\sum_{k=0}^{m}C_{2,4}(|c|^{m-k}+1)n^{k}\left|\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\right|=(|c|^{m}+1)O(e^{-\delta n})$
-as $n\to\infty$ for some $\delta>0$, by using (2.19) and (2.20). Then, (2.21), (2.22) and Lemma 2.7 imply that the sum of the third and fourth terms on the right-hand side of (2.6) is dominated by
+as $n\to\infty$ for some $\delta>0$, by using (2.19) and (2.20).
+Moreover, (2.21), (2.22) and Lemma 2.7 imply that the modulus of the sum of the third and fourth terms on the right-hand side of (2.6) is dominated by
 $(1-x)^{\alpha n}\psi(\alpha)\sum_{k=0}^{m}C_{2,2}(|c|^{m-k}+1)n^{k}+x^{\alpha n}\psi(\alpha)\sum_{k=0}^{m}C_{2,3}(|c+\alpha n|^{m-k}+1)n^{k},$
-which is $(|c|^{m}+1)O(e^{-\delta n})$ as well. Note that these items are regarded as 0 if $\alpha\in\mathbb{N}$. ∎
-Note that Proposition 2.6 with $c=-\alpha nx$ is often used in the rest of this paper.
-When $c=-\alpha nx$, the second term in the right-hand side of (2.14) is $O(e^{-\delta n})$ with possibly different $\delta>0$.
+which is $(|c|^{m}+1)O(e^{-\delta n})$ as well. Thus, (2.14) holds. Note that these items are regarded as 0 if $\alpha\in\mathbb{N}$. ∎
+We often use Proposition 2.6 (2) with $c=-\alpha nx$ in the rest of this paper.
+When $c=-\alpha nx$, the second term on the right-hand side of (2.14) is $O(e^{-\delta n})$ with possibly different $\delta>0$.
 A direct calculation also gives
@@
 Corollary 2.8.
-Let $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is given by $\mu_{\alpha,x}^{(n)}$. Then, for some $\delta>0$, we have
+Let $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is $\mu_{\alpha,x}^{(n)}$. Then, for some $\delta>0$, we have
 $\mathbb{E}[S_{\alpha,x}^{(n)}]=\sum_{j=0}^{n}j\mu_{\alpha,x}^{(n)}(\{j\})=nx+O(e^{-\delta n})$ (2.24)
@@
 Proof.
-Applying Proposition 2.6 with $c=0$ and $m=1$, we have
+Applying Proposition 2.6 (2) with $c=0$ and $m=1$, we have
 $\alpha\sum_{j=0}^{n}(\alpha j)\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}$ $=W_{0}(1,0)+W_{0}(1,1)\alpha nx+O(e^{-\delta n})$
 $=\alpha nx+O(e^{-\delta n})$
-as $n\to\infty$, which implies (2.24). We also apply (2.13), (2.23), and Proposition 2.6 with $c=-\alpha nx$ and $m=2$ to obtain
+as $n\to\infty$. This and Proposition 2.6 (1) imply (2.24). We also apply (2.23) and Proposition 2.6 (2) with $c=-\alpha nx$ and $m=2$ to obtain
 $\sum_{j=0}^{n}(j-nx)^{2}\mu_{\alpha,x}^{(n)}(\{j\})=\frac{1}{Z_{\alpha,x}^{(n)}}\left\{\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})\right\}=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$
@@
 as $n\to\infty$. ∎
-2.3. Uniform bound of $Z_{\alpha,x}^{(n)}$
-Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. We will give a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to $x$ and $n$, and again we note that $Z_{1,x}^{(n)}=1$.
+2.3 Uniform bound of $Z_{\alpha,x}^{(n)}$
+Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. We establish a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to $x$ and $n$, and again we note that $Z_{1,x}^{(n)}=1$.
 Theorem 2.9 (Uniform bound of $Z_{\alpha,x}^{(n)}$).
@@
 $\alpha\wedge 1\leq Z_{\alpha,x}^{(n)}\leq\alpha\vee 1$ (2.25)
-for $x\in[0,1]$ and $n\in\mathbb{N}$.
+for $n\in\mathbb{N}$ and $x\in[0,1]$.
 (2) Let $\alpha\in[2,\infty)$. Then, we have
 $\alpha^{1-\alpha^{2}}\leq Z_{\alpha,x}^{(n)}\leq\alpha$ (2.26)
-for $x\in[0,1]$ and $n\in\mathbb{N}$.
+for $n\in\mathbb{N}$ and $x\in[0,1]$.
 Proof.
-If $\alpha\in(0,1)$, then Proposition 1.1 implies that $Z_{\alpha,x}^{(n)}\leq 1$ for $x\in[0,1]$ and $n\in\mathbb{N}$.
-On the other hand, Equation (1.7) in [GS23, Theorem 1.3] leads to $Z_{\alpha,x}^{(n)}\leq\alpha$ for $x\in[0,1]$ and $n\in\mathbb{N}$ when $\alpha\geq 1$.
+If $\alpha\in(0,1)$, then [12, Theorem 1.2] implies that $Z_{\alpha,x}^{(n)}\leq 1$ for $n\in\mathbb{N}$ and $x\in[0,1]$.
+On the other hand, Equation (1.7) in [10, Theorem 1.3] leads to $Z_{\alpha,x}^{(n)}\leq\alpha$ for $n\in\mathbb{N}$ and $x\in[0,1]$ when $\alpha\geq 1$.
 Hence, the uniform upper bounds of (2.25) and (2.26) have been obtained.
@@
 Because it holds that $Z_{\alpha,x}^{(n)}=Z_{\alpha,1-x}^{(n)}$ and $Z_{\alpha,0}^{(n)}=\alpha$, it suffices to consider only the case where $x\in(0,1/2]$.
-It follows from [HH10, Proposition 2.6] that
+It follows from [12, Proposition 2.6] that
 $R_{\alpha}^{(n)}(\lambda)\in\begin{cases}(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ (1-\alpha,0)&\text{if }\alpha\in(1,2)\end{cases}$
@@
 Letting $\lambda=x/(1-x)$ and multiplying both sides of (2.27) by $(1-x)^{\alpha n}$, we have
-$-\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}+1\in\begin{cases}\left(0,(1-\alpha)(1-x)^{\alpha n}\right)\subset(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ \left((1-\alpha)(1-x)^{\alpha n},0\right)&\text{if }\alpha\in(1,2).\end{cases}$
+$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}+1\in\begin{cases}\left(0,(1-\alpha)(1-x)^{\alpha n}\right)\subset(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ \left((1-\alpha)(1-x)^{\alpha n},0\right)&\text{if }\alpha\in(1,2).\end{cases}$
 This implies that $Z_{\alpha,x}^{(n)}>\alpha$ if $\alpha\in(0,1)$ and $Z_{\alpha,x}^{(n)}>1$ if $\alpha\in(1,2)$. Hence, all that remains is the case $\alpha\geq 2$. We have
@@
 $=\alpha(1-x)^{\frac{1}{x}\times\alpha nx}$
-$\geq\alpha\left(1-\frac{1}{2}\right)^{2\alpha nx}=\alpha\times 2^{-2\alpha nx},\qquad x\in(0,1/2],\,n\in\mathbb{N}$ (2.28)
+$\geq\alpha\left(1-\frac{1}{2}\right)^{2\alpha nx}=\alpha\times 2^{-2\alpha nx},\qquad n\in\mathbb{N},\,x\in(0,1/2]$ (2.28)
 by noting the fact that the function $(0,1/2]\ni x\longmapsto(1-x)^{1/x}\in\mathbb{R}$ is monotonically decreasing. Let $\gamma>0$ be a constant that is specified later. Then, (2.28) implies that
-$Z_{\alpha,x}^{(n)}\geq\alpha\times 2^{-2\gamma},\qquad x\in(0,1/2],\,n\in\mathbb{N}\,\,\text{with $\alpha nx\leq\gamma$.}$
-We now deal with the upper bound of $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$ for $x\in(0,1/2]$ and $n\in\mathbb{N}$ with $\alpha nx>\gamma$. From (2.18), we have
+$Z_{\alpha,x}^{(n)}\geq\alpha\times 2^{-2\gamma},\qquad n\in\mathbb{N},\,x\in(0,1/2]\,\,\text{with $\alpha nx\leq\gamma$.}$
+We now deal with the upper bound of $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$ for $n\in\mathbb{N}$ and $x\in(0,1/2]$ with $\alpha nx>\gamma$. From (2.18), we have
 $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$
@@
 $\leq\alpha\exp\left(-\alpha nx\times\frac{1}{2}\times\frac{8}{\alpha^{2}}\right)=\alpha\exp\left(-\alpha nx\times\frac{4}{\alpha^{2}}\right)$ (2.29)
-for $x\in(0,1/2]$ and $n\in\mathbb{N}$, where we used #$\#(K_{\alpha})=\lfloor\alpha/2\rfloor+\lceil\alpha/2\rceil\leq\alpha+1$ in the fourth line and
+for $n\in\mathbb{N}$ and $x\in(0,1/2]$, where we used $(K_{\alpha})=\lfloor\alpha/2\rfloor+\lceil\alpha/2\rceil\leq\alpha+1$ in the fourth line and
 $\max_{\omega\in K_{\alpha}\setminus\{1\}}\mathrm{Re}(\omega)=\cos\frac{2\pi}{\alpha}\leq 1-\frac{8}{\alpha^{2}}$
@@
 $Z_{\alpha,x}^{(n)}\geq\min\left\{\alpha\times 2^{-2\gamma},1-\alpha\exp\left(-\frac{4\gamma}{\alpha^{2}}\right)-2e^{-\gamma}\right\}\Bigg|_{\gamma=(\alpha^{2}\log\alpha)/(2\log 2)}=\alpha^{1-\alpha^{2}}$
-for $x\in(0,1/2]$ and $n\in\mathbb{N}$. ∎
-2.4. Uniform estimates for moments of the fractional binomial distribution
+for $n\in\mathbb{N}$ and $x\in(0,1/2]$. ∎
+2.4 Uniform estimates for moments of the fractional binomial distribution
 In Theorem 2.5, we discussed the explicit representations of moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
@@
 $\sup_{\alpha\in K}\max_{x\in[0,1]}\Bigg|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}$
-$\hskip 28.45274pt-(\alpha n)^{-m}\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}\Bigg{|}=O\left(\frac{1}{n^{m}}\right)$ (2.30)
+$\qquad\qquad-(\alpha n)^{-m}\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}\Bigg|=O\left(\frac{1}{n^{m}}\right)$ (2.30)
 as $n\to\infty$.
@@
 $=\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}$
-$\hskip 28.45274pt+\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}\Bigg{\{}\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\Bigg{\}}$
-$\hskip 28.45274pt-\Bigg{\{}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
-$\hskip 28.45274pt+(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg{\}}$
-$=:\mathrm{I}_{\alpha,m}^{(n)}(x)+\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x).$ (2.31)
+$\quad+\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}\Bigg\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\Bigg\}$
+$\quad-\Bigg\{\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
+$\quad+(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg\}$
+$=:\mathrm{I}_{\alpha,m}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)-\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x).$ (2.31)
 We note here that
@@
 and
-$\mathrm{III}_{\alpha,m}^{(n)}(1-x)$ $=\frac{\alpha(1-x)^{\alpha}x^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(tx)^{\alpha}-(1-x)^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
-$\hskip 28.45274pt+(-1)^{m}\frac{\alpha(1-x)^{\alpha(n+1)}x^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|x^{\alpha}e^{-i\alpha\pi}-(t(1-x))^{\alpha}|^{2}}\,\mathrm{d}t$
-$=(-1)^{m}\mathrm{III}_{\alpha,m}^{(n)}(x)$
-for $x\in(0,1)$. Furthermore, $\mathrm{II}_{\alpha,m}(x)=0$ if $\alpha\in(0,2)$ because $K_{\alpha}=\{1\}$. Therefore, it follows from (2.31) that for $\alpha\in(0,2)$,
+$\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(1-x)$ $=\frac{\alpha(1-x)^{\alpha}x^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(tx)^{\alpha}-(1-x)^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
+$\quad+(-1)^{m}\frac{\alpha(1-x)^{\alpha(n+1)}x^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|x^{\alpha}e^{-i\alpha\pi}-(t(1-x))^{\alpha}|^{2}}\,\mathrm{d}t$
+$=(-1)^{m}\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$
+for $x\in(0,1)$. Furthermore, $\mathrm{I\hskip-1.00006ptI}_{\alpha,m}(x)=0$ if $\alpha\in(0,2)$ because $K_{\alpha}=\{1\}$. Therefore, it follows from (2.31) that for $\alpha\in(0,2)$,
 $\mathrm{I}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{I}_{\alpha,m}^{(n)}(x).$ (2.32)
 Because $\mathrm{I}_{\alpha,m}^{(n)}(x)$ is a polynomial in $\alpha$ thanks to Lemma 2.3, (2.32) holds true for arbitrary $\alpha>0$, which in turn implies that
-$\mathrm{II}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{II}_{\alpha,m}^{(n)}(x),\qquad\alpha>0.$ (2.33)
+$\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x),\qquad\alpha>0.$ (2.33)
 In what follows, the integer $n\in\mathbb{N}$ is taken to be sufficiently large so that $\alpha n\geq m+1$.
 Step 2.
-We give a uniform estimate of $\mathrm{II}_{\alpha,m}^{(n)}(x)$.
+We give a uniform estimate of $\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$.
 Suppose that $\omega\in K_{\alpha}\setminus\{1\}$ and $x\in(0,1/2]$.
@@
 by using (2.18), (2.20), and Lemma 2.10. Thus, we have
-$\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq(m+1)(\#(K_{\alpha})-1)C_{2,5}\leq C_{2,6},$
+$\sup_{x\in(0,1/2]}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq(m+1)(\#(K_{\alpha})-1)C_{2,5}\leq C_{2,6},$
 where $C_{2,6}>0$ depends on only $m\in\mathbb{N}$ and a compact set $K$ in which $\alpha$ takes values. By (2.33), we obtain that
-$\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq C_{2,6}.$ (2.34)
-Step 3. We next give a uniform estimate of $\mathrm{III}_{\alpha,m}^{(n)}(x)$. We write
-$\mathrm{III}_{\alpha,m}^{(n)}(x)=\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)+(-1)^{m}\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(1-x)$
+$\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1/2]}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,6}.$ (2.34)
+Step 3. We next give a uniform estimate of $\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$. We write
+$\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)=\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)+(-1)^{m}\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(1-x)$
 with
-$\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)=\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
+$\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)=\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
 It follows from (2.20) that
@@
 Thus, we have
-$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,8}\sum_{k=0}^{m}x^{\alpha}(1-x)^{\alpha(n+1)}|\sin\alpha\pi|$
-×$\hskip 28.45274pt\times\{(nx)^{m-k}+1\}\int_{0}^{1}\frac{t^{\alpha-1}n^{k}t^{k}(1-t)^{\alpha n-k}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
+$|\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,8}\sum_{k=0}^{m}x^{\alpha}(1-x)^{\alpha(n+1)}|\sin\alpha\pi|$
+$\quad\times\{(nx)^{m-k}+1\}\int_{0}^{1}\frac{t^{\alpha-1}n^{k}t^{k}(1-t)^{\alpha n-k}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
 By noting
@@
 and Lemma 2.7, we have
-$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,10}\sum_{k=0}^{m}(1-x)^{\alpha n}\{(nx)^{m-k}+1\}$
-$\leq C_{2,10}\sum_{k=0}^{m}\{(1-x)^{\alpha n}(nx)^{m-k}+(1-x)^{\alpha n}\}$
+$|\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,10}\sum_{k=0}^{m}(1-x)^{\alpha n}\{(nx)^{m-k}+1\}$
+$=C_{2,10}\sum_{k=0}^{m}\{(1-x)^{\alpha n}(nx)^{m-k}+(1-x)^{\alpha n}\}$
 $\leq C_{2,10}\sum_{k=0}^{m}\left\{\left(\frac{m-k}{e\alpha n}\right)^{m-k}n^{m-k}+1\right\}\leq C_{2,11},$
 from which we conclude that
-$\sup_{x\in(0,1)}|\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,12},$ (2.35)
+$\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,12},$ (2.35)
 where $C_{2,12}>0$ depends on only $m\in\mathbb{N}$ and $K$ in which $\alpha$ takes values.
 By combining (2.31) with (2.34) and (2.35), we obtain that
-$\sup_{x\in(0,1)}|(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)-\mathrm{I}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,13}.$
+$\sup_{x\in(0,1)}|(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)-\mathrm{I}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)-\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,13}.$
 This implies (2.11) once we divide both sides of the above inequality by $(\alpha n)^{m}$ and take the supremum over $\alpha\in K$. ∎
-We also have the following by applying Theorem 2.11 and (2.23).
+We also have the following by applying Theorems 2.9 and 2.11, and (2.23).
 Corollary 2.12.
 Let $K$ be a compact subset of $(0,\infty)$. As $n\to\infty$,
-$\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}\right|=\begin{cases}\displaystyle O\left(\frac{1}{n}\right)&\text{if }m=1,2,\vspace{2mm}\\ \displaystyle O\left(\frac{1}{n^{2}}\right)&\text{if }m=3,4.\end{cases}$
-2.5. Characteristic function of the fractional binomial distribution
+$\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\mathbb{E}\left[\left(\frac{S_{\alpha,x}^{(n)}}{n}-x\right)^{m}\right]\right|=\begin{cases}\displaystyle O\left(\frac{1}{n}\right)&\text{if }m=1,2,\\ \displaystyle O\left(\frac{1}{n^{2}}\right)&\text{if }m=3,4.\end{cases}$
+2.5 Characteristic function of the fractional binomial distribution
 In this section, we obtain a quantitative estimate of the characteristic function of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, that is,
@@
 $=\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 14.22636pt\times\Bigg{\{}\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\Bigg{\}}\,\mathrm{d}t$ (2.37)
+$\quad\times\Bigg\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\Bigg\}\,\mathrm{d}t$ (2.37)
 for $\lambda\in D_{\alpha}:=\{re^{i\theta}\mid r>0,\,|\theta|<\theta_{\alpha}/\alpha\}$.
@@
 Then, (2.36) holds for all $\lambda\in(0,1]$ by Proposition 1.1.
-Because both sides of (2.36) are holomorphic on $\mathbb{C}$, (2.36) holds for all $\lambda\in\mathbb{C}$ by analytic continuation.
+Because both sides of (2.36) are holomorphic on $\mathbb{C}$, (2.36) holds for all $\lambda\in\mathbb{C}$ by the identity theorem for holomorphic functions.
 We next assume $\alpha\notin\mathbb{N}$.
@@
 where $\arg(e^{\mp i\alpha\pi})$ is taken from $(0,2\pi)$.
-Because $\arg(e^{-i\alpha\pi})=(\overline{\alpha}-\alpha)\pi$ and $\arg(e^{i\alpha\pi})=(\alpha-\underline{\alpha})\pi$, (2.38) is rewritten as $-\eta_{\alpha}<\theta<\eta_{\alpha}$, where
-$\eta_{\alpha}=-\big{[}\{2-(\alpha-\underline{\alpha})\}\wedge\{2-(\overline{\alpha}-\alpha)\}\wedge(\overline{\alpha}-\alpha)\wedge(\alpha-\underline{\alpha})\big{]}\frac{\pi}{\alpha}.$
-Noting that $\eta_{\alpha}=\theta_{\alpha}/\alpha$ because $\overline{\alpha}-\underline{\alpha}=2$, both sides of (2.37) turn out to be holmorphic in the domain $D_{\alpha}$.
-Thus, (2.37) also holds for all $\lambda\in D_{\alpha}$ by analytic continuation.
+Because $\arg(e^{-i\alpha\pi})=(\overline{\alpha}-\alpha)\pi$ and $\arg(e^{i\alpha\pi})=(\alpha-\underline{\alpha})\pi$, (2.38) is rewritten as $\eta_{\alpha}<\theta<\eta_{\alpha}$, where
+$\eta_{\alpha}=\big[\{2-(\alpha-\underline{\alpha})\}\wedge\{2-(\overline{\alpha}-\alpha)\}\wedge(\overline{\alpha}-\alpha)\wedge(\alpha-\underline{\alpha})\big]\frac{\pi}{\alpha}.$
+By noting that $\eta_{\alpha}=\theta_{\alpha}/\alpha$ because $\overline{\alpha}-\underline{\alpha}=2$, both sides of (2.37) turn out to be holomorphic in the domain $D_{\alpha}$.
+Thus, (2.37) also holds for all $\lambda\in D_{\alpha}$ by the identity theorem.
 ∎
@@
 $=(a-b\cos\theta)^{2}+4b^{2}\sin^{2}\frac{\theta}{2}\cos^{2}\frac{\theta}{2}$
-$\geq 2b^{2}\sin^{2}\frac{\theta}{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.\qed$
+$\geq 2b^{2}\sin^{2}\frac{\theta}{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.$
 A quantitative estimate of $\varphi_{\alpha,x}^{(n)}$ is now stated as follows.
@@
 $\varphi_{\alpha,x}^{(n)}(\xi)$ $=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega e^{i\xi/\alpha}\right)^{\alpha n}$
-$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\Bigg{[}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
-$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)}}{\{(1-x)^{\alpha}e^{-i(\xi+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{]}\,\mathrm{d}t$ (2.39)
+$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
+$\quad\times\Bigg[\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
+$\qquad+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)}}{\{(1-x)^{\alpha}e^{-i(\xi+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg]\,\mathrm{d}t$ (2.39)
 $=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+O(e^{-\delta n})$ (2.40)
-as $n\to\infty$ for some $\delta>0$ that is chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\xi\in(-\theta_{\alpha},\theta_{\alpha})$.
-Here, if $\alpha\in\mathbb{N}$, then the last term of (2.39) is regarded as 00.
+as $n\to\infty$ for some $\delta>0$ that is chosen uniformly on any compact subset of $x\in(0,1)$ and any compact subset of $\xi\in(-\theta_{\alpha},\theta_{\alpha})$.
+Here, if $\alpha\in\mathbb{N}$, then the last term of (2.39) is regarded as 0.
 Proof.
@@
 $=\frac{1}{Z_{\alpha,x}^{(n)}}(1-x)^{\alpha n}\Bigg[\sum_{\omega\in K_{\alpha}}(1+\lambda\omega)^{\alpha n}-\frac{\alpha\lambda^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\left\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\right\}\,\mathrm{d}t\Bigg{]}$
+$\quad\times\left\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\right\}\,\mathrm{d}t\Bigg]$
 $=\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+xe^{i\xi/\alpha}\right)^{\alpha n}+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega e^{i\xi/\alpha}\right)^{\alpha n}$
-$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n-1)}e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\Bigg{[}\frac{(1-x)^{2\alpha}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
-$\hskip 28.45274pt+\frac{x^{\alpha n}(1-x)^{-\alpha(n-2)}e^{i\xi n}}{\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}}\Bigg{]}\,\mathrm{d}t,$
+$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n-1)}e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
+$\quad\times\Bigg[\frac{(1-x)^{2\alpha}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
+$\qquad+\frac{x^{\alpha n}(1-x)^{-\alpha(n-2)}e^{i\xi n}}{\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}}\Bigg]\,\mathrm{d}t,$
 which implies (2.39). Next, we prove (2.40). We have
@@
 $|(1-x)^{\alpha}e^{-i(\xi\pm\alpha\pi)}-(tx)^{\alpha}|\geq(1-x)^{\alpha}\left|\sin\frac{\xi\pm\alpha\pi}{2}\right|.$
-Note that $|\xi|<\theta_{\alpha}$ implies $\sin(\xi\pm\alpha\pi)/2\neq 0$. From these estimates and (2.13), we obtain (2.40). The case for $\alpha\in\mathbb{N}$ is simpler to prove. ∎
-2.6. The fractional multinomial distribution
+Note that $|\xi|<\theta_{\alpha}$ implies $\sin(\xi\pm\alpha\pi)/2\neq 0$. From these estimates and Proposition 2.6 (1), we obtain (2.40). The case for $\alpha\in\mathbb{N}$ is simpler to prove. ∎
+2.6 The fractional multinomial distribution
 Let $d\geq 2$ be an integer. For $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\mathbb{R}^{d}$, we write
 $|\bm{x}|:=x_{1}+x_{2}+\cdots+x_{d},\qquad\|\bm{x}\|:=(x_{1}^{2}+x_{2}^{2}+\cdots+x_{d}^{2})^{1/2}.$
-We define the $(d-1)$-simplex in $\mathbb{R}^{d}$ by
+We define the (d1)(d-1)-simplex in $\mathbb{R}^{d}$ by
 $\Delta_{d-1}:=\left\{\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\mathbb{R}^{d}\,|\,x_{i}\geq 0,i=1,2,\dots,d,\,|\bm{x}|=1\right\}.$
@@
 $J_{d}^{(n)}=\{\bm{j}=(j_{1},j_{2},\dots,j_{d})\mid j_{k}\in\mathbb{N}_{0},\,k=1,2,\dots,d,\,|\bm{j}|=n\}.$
-According to the preliminary version [FR11, Lemma 1] of the paper [FR13], we can generalize Proposition 1.1 to the multivariate cases, which is stated as follows.
-Theorem 2.16 (Multivariate neo-classical inequality; see also [LCL07, Exercise 3.9]).
+According to the preliminary version [8, Lemma 1] of the paper [9], we can generalize the neo-classical inequality presented in [12, Theorem 1.2] to the multivariate cases, which is stated as follows.
+Theorem 2.16 (Multivariate neo-classical inequality; see also [23, Exercise 3.9]).
 Let $\alpha\in(0,2)$, $n\in\mathbb{N}$ and $x_{1},x_{2},\dots,x_{d}\geq 0$. Then, we have
 $\alpha^{d}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
-$\begin{cases}\leq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(0,1],\\ \geq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(1,2).\\ \end{cases}$ (2.41)
+[…]
 The equality holds if and only if $\alpha=1$ or $x_{1}=x_{2}=\cdots=x_{d}=0$.
@@
 $=\alpha\sum_{j_{d+1}=0}^{n}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha(n-j_{d+1})+1)\Gamma(\alpha j_{d+1}+1)}x_{d+1}^{\alpha j_{d+1}}$
-×$\hskip 28.45274pt\times\alpha^{d}\sum_{(j_{1},j_{2},\dots,j_{d})\in J_{d}^{(n-j_{d+1})}}\frac{\Gamma(\alpha(n-j_{d+1})+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
+$\quad\times\alpha^{d}\sum_{(j_{1},j_{2},\dots,j_{d})\in J_{d}^{(n-j_{d+1})}}\frac{\Gamma(\alpha(n-j_{d+1})+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
 $\leq\alpha\sum_{j_{d+1}=0}^{n}\binom{\alpha n}{\alpha j_{d+1}}(x_{1}+x_{2}+\cdots+x_{d})^{\alpha(n-j_{d+1})}x_{d+1}^{\alpha j_{d+1}}$
@@
 The proof for $\alpha\in(1,2)$ is similar to that for $\alpha\in(0,1]$. ∎
-Needless to say, if $\alpha=1$, the (2.41) reads as the usual multinomial theorem.
+Needless to say, if $\alpha=1$, the (2.41) reads as usual multinomial theorem.
 As with Definition 1.2, this multivariate neo-classical inequality allows us to define the fractional analogue of the multinomial distribution.
@@
 Let $\alpha>0$, $n\in\mathbb{N}$ and $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\Delta_{d-1}$.
-Then, a probability measure $\mu_{\alpha,\bm{x}}^{(n)}$ on $J_{d}^{(n)}$ is called the $\alpha$-fractional multinomial distribution if
-$\mu_{\alpha,\bm{x}}^{(n)}(\{\bm{j}\})=\frac{\alpha^{d}}{Z_{\alpha,\bm{x}}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$ (2.42)
-for $\bm{j}\in J_{d}^{(n)}$, where $Z_{\alpha,\bm{x}}^{(n)}$ is the normalized constant given by
+Then, a probability measure $\mu_{\alpha,\bm{x}}^{(n)}$ is called the $\alpha$-fractional multinomial distribution if
+$\mu_{\alpha,\bm{x}}^{(n)}(\mathrm{d}\bm{z}):=\frac{\alpha^{d}}{Z_{\alpha,\bm{x}}^{(n)}}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}\delta_{\bm{j}}(\mathrm{d}\bm{z}),$ (2.42)
+where $Z_{\alpha,\bm{x}}^{(n)}$ is the normalized constant given by
 $Z_{\alpha,\bm{x}}^{(n)}=\alpha^{d}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}\,(>0).$
@@
 for $d>2$.
-Obtaining such a formula explicitly might require a multivariate version of the fractional Taylor series, which although seeming rather difficult might nevertheless be worth investigating further.
-3. Limit theorems for the fractional binomial distribution
-3.1. The weak law of large numbers
-First, we prove the weak law of large numbers for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
+Obtaining such a formula explicitly might require a multivariate version of the fractional Taylor series, which, although it appears rather difficult, may nevertheless be worth further investigation.
+3 Limit theorems for the fractional binomial distribution
+3.1 The weak law of large numbers
+First, we prove the weak law of large numbers for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
 Theorem 3.1 (Weak law of large numbers for $\mu_{\alpha,x}^{(n)}$).
@@
 $=\frac{1}{\varepsilon^{2}}\sum_{j=0}^{n}\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}$
-$\leq\frac{1}{\varepsilon^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|=O\left(\frac{1}{n}\right)$
+$\leq\frac{1}{\varepsilon^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|$
+$=O\left(\frac{1}{n}\right)$
 as $n\to\infty$, which implies (3.1). ∎
-3.2. The central limit theorem
-We show the central limit theorem for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, which is stated as follows.
+3.2 The central limit theorem
+We prove the central limit theorem for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, which is stated as follows.
 Theorem 3.2 (Central limit theorem for $\mu_{\alpha,x}^{(n)}$).
-Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Let $\widetilde{\mu}_{\alpha,x}^{(n)}$ denote the probability law of the random variable
+Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Let $\widetilde{\mu}_{\alpha,x}^{(n)}$ denote the law of the random variable
 $\widetilde{S}_{\alpha,x}^{(n)}:=\frac{S_{\alpha,x}^{(n)}-\mathbb{E}[S_{\alpha,x}^{(n)}]}{\sqrt{\mathrm{Var}(S_{\alpha,x}^{(n)})}}.$
@@
 $=\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\Bigg[\frac{1}{Z_{\alpha,x}^{(n)}}\left(1-x+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
-$\hskip 28.45274pt+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
-$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha\sin\alpha\pi}{\pi}\exp\left(\frac{i\xi}{\sqrt{v_{n}}}\right)\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
-×$\hskip 28.45274pt\times\Bigg{\{}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}+\alpha\pi)}\}}$
-$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)/\sqrt{v_{n}}}}{\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{\}}\,\mathrm{d}t\Bigg{]}.$ (3.2)
+$\quad+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
+$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha\sin\alpha\pi}{\pi}\exp\left(\frac{i\xi}{\sqrt{v_{n}}}\right)\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
+$\quad\times\Bigg\{\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}+\alpha\pi)}\}}$
+$\qquad+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)/\sqrt{v_{n}}}}{\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg\}\,\mathrm{d}t\Bigg].$ (3.2)
 From Corollary 2.8, we have
@@
 $\exp\left(-\frac{i\xi m_{n}}{\sqrt{v_{n}}}\right)\left(1-x+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
-=$=\left\{(1-x)\exp\left(-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}\right)+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)\right)\right\}^{\alpha n}$
+$=\left\{(1-x)\exp\left(-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}\right)+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)\right)\right\}^{\alpha n}$
 $=\Bigg[(1-x)\left\{1-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}-\frac{\xi^{2}m_{n}^{2}}{2(\alpha n\sqrt{v_{n}})^{2}}+O(n^{-3/2})\right\}$
-$\hskip 28.45274pt+x\left\{1+\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)-\frac{\xi^{2}}{2(\alpha\sqrt{v_{n}})^{2}}\left(1-\frac{m_{n}}{n}\right)^{2}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
+$\quad+x\left\{1+\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)-\frac{\xi^{2}}{2(\alpha\sqrt{v_{n}})^{2}}\left(1-\frac{m_{n}}{n}\right)^{2}+O(n^{-3/2})\right\}\Bigg]^{\alpha n}$
 $=\Bigg[(1-x)\left\{1-\frac{i\xi x}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}x^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}$
-$\hskip 28.45274pt+x\left\{1+\frac{i\xi(1-x)}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}(1-x)^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
-=$=\left\{1-\frac{\xi^{2}}{2\alpha n}+O(n^{-3/2})\right\}^{\alpha n}\to\exp\left(-\frac{\xi^{2}}{2}\right)$ (3.3)
+$\quad+x\left\{1+\frac{i\xi(1-x)}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}(1-x)^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}\Bigg]^{\alpha n}$
+$=\left\{1-\frac{\xi^{2}}{2\alpha n}+O(n^{-3/2})\right\}^{\alpha n}\to\exp\left(-\frac{\xi^{2}}{2}\right)$ (3.3)
 as $n\to\infty$. For each $\omega\in K_{\alpha}\setminus\{1\}$, we also have
@@
 $|(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}\pm\alpha\pi)}-(tx)^{\alpha}|\geq(1-x)^{\alpha}\left|\sin\frac{1}{2}\left(\frac{\xi}{\sqrt{v_{n}}}\pm\alpha\pi\right)\right|\to(1-x)^{\alpha}\left|\sin\frac{\alpha\pi}{2}\right|>0$
-as $n\to\infty$, (2.13) implies that the third term of the right-hand side of (3.2) converges to zero as $n\to\infty$. Therefore, (3.2), (3.3), and (3.4) lead to
+as $n\to\infty$, Proposition 2.6 (1) implies that the third term of the right-hand side of (3.2) converges to zero as $n\to\infty$. Therefore, (3.2), (3.3), and (3.4) lead to
 $\lim_{n\to\infty}\widetilde{\varphi}_{\alpha,x}^{(n)}(\xi)=\exp\left(-\frac{\xi^{2}}{2}\right),\qquad\xi\in\mathbb{R}.$
 Because the right-hand side of the above is nothing but the characteristic function of the normal distribution $N(0,1)$, Lévy’s continuity theorem (cf.
-[Kle08, Theorem 15.23]) concludes the desired weak convergence.
+[17, Theorem 15.23]) concludes the desired weak convergence.
 ∎
-3.3. The law of small numbers
+By combining Corollary 2.8 and Theorem 3.2, we also have the following.
+Corollary 3.3.
+In the notation of Theorem 3.2, the law of $\frac{S_{\alpha,x}^{(n)}-nx}{\sqrt{nx(1-x)/\alpha}}$ converges weakly to $N(0,1)$ as $n\to\infty$.
+3.3 The law of small numbers
 Recall that the Mittag-Leffler function of parameter $\alpha>0$ is defined by
@@
 Definition 3.4 (Fractional Poisson distribution).
-Let $\alpha>0$ and $\lambda>0$. The $\alpha$-fractional Poisson distribution with parameter $\lambda$ is a probability measure $\nu_{\alpha,\lambda}$ on $\mathbb{N}_{0}$ defined as
-$\nu_{\alpha,\lambda}(\{j\})=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)},\qquad j\in\mathbb{N}_{0}.$
+Let $\alpha>0$ and $\lambda>0$. The $\alpha$-fractional Poisson distribution with parameter $\lambda$ is a probability measure $\nu_{\alpha,\lambda}$ defined as
+$\nu_{\alpha,\lambda}(\mathrm{d}z):=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\sum_{j=0}^{\infty}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}\delta_{j}(\mathrm{d}z).$
 The $1$-fractional Poisson distribution coincides with the Poisson distribution.
@@
 We note that $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions defined via the generalized Mittag-Leffler function $E_{\alpha,\beta}(z)$ given by (3.9) below.
-See [Her16, CO17] for some properties as well as its definition.
-Here, we show that $\nu_{\alpha,\lambda}$ is captured through the law of small numbers for the fractional binomial distribution.
+See [13, 6] for some properties as well as its definition.
+We also refer to [19] for the study of non-Markovian fractional Poisson processes based on Riemann–Liouville type fractional calculus, and [21] for some generalizations of Bernoulli processes having long-range dependence and their convergences to fractional Poisson processes, both of which are different concepts from ours.
+We also mention that the Mittag-Leffler function has already been used to introduce some probability distributions in different situations.
+See [26, 27] for the notion of the Mittag-Leffler distribution, which is, however, far from the Poisson distribution.
+Here, we prove that $\nu_{\alpha,\lambda}$ is captured through the law of small numbers for the fractional binomial distribution.
 Theorem 3.5 (Law of small numbers for $\mu_{\alpha,x}^{(n)}$).
-Let $\alpha>0$, $\lambda>0$, and $x\in(0,1)$.
-Suppose that two sequences $\{\alpha_{n}\}_{n=1}^{\infty}\subset(0,\infty)$ and $\{x_{n}\}_{n=1}^{\infty}\subset(0,1)$ satisfy that $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ as $n\to\infty$.
+Let $\alpha>0$ and $\lambda>0$.
+Suppose that two sequences $\{\alpha_{n}\}_{n=1}^{\infty}\subset(0,\infty)$ and $\{x_{n}\}_{n=1}^{\infty}\subset(0,1)$ satisfy that $\alpha_{n}\to\alpha$ and $nx_{n}\to\lambda/\alpha$ as $n\to\infty$.
 Then, $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to $\nu_{\alpha,\lambda}$ as $n\to\infty$.
@@
 $b_{\alpha_{n},x_{n}}^{(n)}(j)=\alpha_{n}\binom{\alpha_{n}n}{\alpha_{n}j}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}\bm{1}_{[0,n]}(j),\qquad j\in\mathbb{N}_{0}.$
-Because $\alpha_{n}\to\alpha>0$ and $\alpha_{n}nx_{n}\to\lambda>0$, there exist four constants $C_{3,1},C_{3,2},C_{3,3},C_{3,4}>0$ such that
+Because $\alpha_{n}\to\alpha>0$ and $nx_{n}\to\lambda/\alpha>0$, there exist four constants $C_{3,1},C_{3,2},C_{3,3},C_{3,4}>0$ such that
 $C_{3,1}\leq\alpha_{n}\leq C_{3,2},\qquad C_{3,3}\leq nx_{n}\leq C_{3,4},\qquad n\in\mathbb{N}.$ (3.5)
@@
 On the other hand, if $n>j$, then (3.5) and (3.6) allow us to obtain
-(0(0( 0 $\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$
-$=\frac{\alpha_{n}\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}(n-j)+1)\Gamma(\alpha_{n}j+1)}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}$
+$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\frac{\alpha_{n}\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}(n-j)+1)\Gamma(\alpha_{n}j+1)}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}$
 $\leq\frac{C_{3,2}C_{3,6}(\alpha_{n}n)^{\alpha_{n}n+1/2}e^{-\alpha_{n}n}}{C_{3,5}\{\alpha_{n}(n-j)\}^{\alpha_{n}(n-j)+1/2}e^{-\alpha_{n}(n-j)}\Gamma(\alpha_{n}j+1)}\left(\frac{C_{3,4}}{n}\right)^{\alpha_{n}j}$
@@
 Bringing everything together, we have
-$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ =$=\begin{cases}\displaystyle C_{3,2}e^{-C_{3,1}C_{3,3}}&\text{if }j=0,\vspace{2mm}\\ \displaystyle\frac{C_{3,2}C_{3,6}(C_{3,2}C_{3,4}\vee 1)^{C_{3,2}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}(j+1)^{1/2}&\text{if }j=1,2,\dots,n-1,\vspace{2mm}\\ \displaystyle C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}&\text{if }j=n,\\ 0&\text{if }j>n\end{cases}$
+$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\begin{cases}\displaystyle C_{3,2}e^{-C_{3,1}C_{3,3}}&\text{if }j=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{C_{3,2}C_{3,6}(C_{3,2}C_{3,4}\vee 1)^{C_{3,2}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}(j+1)^{1/2}&\text{if }j=1,2,\dots,n-1,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}&\text{if }j=n,\\ 0&\text{if }j>n\end{cases}$
 $\leq C_{3,7}\left\{\frac{C_{3,8}^{C_{3,2}j}}{\Gamma(C_{3,1}j+1)}\vee\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}\right\}$ (3.7)
@@
 as $n\to\infty$ for all $j\in\mathbb{N}_{0}$. ∎
-3.4. Some properties of the fractional Poisson distribution
+3.4 Some properties of the fractional Poisson distribution
 First, we seek the mean and variance of the fractional Poisson distribution $\nu_{\alpha,\lambda}$ introduced in Definition 3.4.
 Lemma 3.6.
-Let $S_{\alpha,\lambda}$ be a random variable whose probability law is $\nu_{\alpha,\lambda}$. Then, we have
+Let $S_{\alpha,\lambda}$ be a random variable whose law is $\nu_{\alpha,\lambda}$. Then, we have
 $\mathbb{E}[S_{\alpha,\lambda}]$ $=\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})},$
 $\mathrm{Var}(S_{\alpha,\lambda})$ $=\frac{\lambda^{\alpha}}{\alpha^{2}E_{\alpha}(\lambda^{\alpha})^{2}}\Big[\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
-$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]},$
+$\quad+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big\{E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big\}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big],$
 where in general $E_{\alpha,\beta}(z)$ is the Mittag-Leffler function of two parameters $\alpha>0$ and $\beta\in\mathbb{C}$ defined by
@@
 and
-$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\varphi_{\alpha,\lambda}(\xi)$ $=i\frac{\lambda^{\alpha}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}ie^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
-$\hskip 28.45274pt+i\lambda^{\alpha}e^{2i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}$
-$=-\frac{\lambda^{\alpha}e^{i\xi}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
-$\hskip 28.45274pt+\lambda^{\alpha}e^{i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}.$ (3.13)
+$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\varphi_{\alpha,\lambda}(\xi)$
+$=i\frac{\lambda^{\alpha}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg\{ie^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})+i\lambda^{\alpha}e^{2i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg\}$
+$=-\frac{\lambda^{\alpha}e^{i\xi}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg\{E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})+\lambda^{\alpha}e^{i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg\}.$ (3.13)
 By substituting $\xi=0$ for (3.12) and (3.13), we obtain
@@
 $\frac{\mathrm{d}}{\mathrm{d}\lambda}M_{\alpha,\lambda}^{(m)}$ $=-m\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m-1}\nu_{\alpha,\lambda}(\{j\})\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$
-$\hskip 28.45274pt+\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
+$\quad+\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
 $=-mM_{\alpha,\lambda}^{(m-1)}\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)+\frac{\alpha}{\lambda}M_{\alpha,\lambda}^{(m+1)}.$
 Because it holds that
-$\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$
-$=\frac{\lambda^{\alpha-1}}{\alpha E_{\alpha}(\lambda^{\alpha})^{2}}\Big{[}\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
-$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]}$
+$\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$ $=\frac{\lambda^{\alpha-1}}{\alpha E_{\alpha}(\lambda^{\alpha})^{2}}\Big[\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
+$\quad+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big\{E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big\}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big]$
 $=\frac{\alpha}{\lambda}\mathrm{Var}(S_{\alpha,\lambda})=\frac{\alpha}{\lambda}M_{\alpha,\lambda}^{(2)},$
 we conclude (3.15). ∎
-4. The fractional Bernstein operator and limit theorems for its iterates
-4.1. The fractional Bernstein operator and its approximating property
+4 The fractional Bernstein operator and limit theorems for its iterates
+4.1 The fractional Bernstein operator and its approximating property
 Let $\alpha>0$.
-In this section, we show that the fractional Bernstein operator $B_{\alpha,n}$ introduced in Definition 1.8 uniformly approximates all continuous functions on $[0,1]$, similarly to Proposition 1.5.
+In this section, we prove that the fractional Bernstein operator $B_{\alpha,n}$ introduced in Definition 1.3 uniformly approximates all continuous functions on $[0,1]$, similarly to the classical case.
 Theorem 4.1.
@@
 Proof.
-Let $\alpha>0$, $n\in\mathbb{N}$, $x\in[0,1]$, and $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is $\mu_{\alpha,x}^{(n)}$. Then, $B_{\alpha,n}f$, $f\in C([0,1])$ can be expressed as
+Let $\alpha>0$, $n\in\mathbb{N}$, $x\in[0,1]$, and $S_{\alpha,x}^{(n)}$ be a random variable whose law is $\mu_{\alpha,x}^{(n)}$. Then, $B_{\alpha,n}f$, $f\in C([0,1])$ can be expressed as
 $B_{\alpha,n}f(x)=\mathbb{E}\left[f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)\right],\qquad x\in[0,1].$ (4.1)
@@
 $\leq\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|<\rho\right]$
-$\hskip 28.45274pt+\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right]$
+$\quad+\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right]$
 $\leq\varepsilon+2\|f\|_{\infty}\mathbb{P}\left(\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right).$
@@
 as $n\to\infty$, which concludes the desired uniform convergence because $\varepsilon>0$ is chosen arbitrarily. ∎
-4.2. The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
+4.2 The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
 It is natural to consider the behaviors of the $k$-fold iterates $B_{\alpha,n}^{k}$ of the $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ as $k\to\infty$.
-In this section, we prove the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$ based on an argument similar to that in the proof of [Rus04, Theorem 1].
+In this section, we prove the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$ based on an argument similar to that in the proof of [28, Theorem 1].
 Theorem 4.2.
@@
 Then, the contraction principle allows us to find the unique function $f_{\alpha,n}\in C([0,1])$ satisfying $B_{\alpha,n}f_{\alpha,n}=f_{\alpha,n}$, $f_{\alpha,n}(0)=f(0)$, $f_{\alpha,n}(1)=f(1)$, and (4.2).
-The explicit representation of $f_{1,n}$ was obtained in Proposition 1.6.
+The explicit representation of $f_{1,n}$ is obvious from [16, Theorem 1].
 ∎
-4.3. Limit theorems for iterates of the fractional Bernstein operator
+4.3 Limit theorems for iterates of the fractional Bernstein operator
 Let $\alpha>0$ be fixed.
@@
 for $f\in C^{2}([0,1])$ and $x\in[0,1]$. The following property is well known.
-Lemma 4.3 (cf. [EK86, Theorem 8.2.8] and [AC94, Theorem 6.2.6]).
+Lemma 4.3 (cf. [7, Theorem 8.2.8] and [2, Theorem 6.2.6]).
 The linear operator $(\mathcal{L}_{\alpha},C^{2}([0,1]))$ is closable on $C([0,1])$ and its closure $(\mathcal{L}_{\alpha},\mathrm{Dom}(\mathcal{L}_{\alpha}))$ generates the Feller semigroup $(T_{\alpha}(t))_{t\geq 0}$ on $C([0,1])$.
-We prove that the discrete semigroup defined by the iterates of $B_{\alpha,n}$ converges to $(T_{\alpha}(t))_{t\geq 0}$ as follows.
+We prove that the appropriately scaled operators $(B_{\alpha,n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, defined by the iterates of $B_{\alpha,n}$ converges to $(T_{\alpha}(t))_{t\geq 0}$ as follows.
 Theorem 4.4.
@@
 $=nf^{\prime}(x)\mathbb{E}\left[\frac{1}{n}S_{\alpha,x}^{(n)}-x\right]$
-$\hskip 28.45274pt+n\mathbb{E}\Bigg{[}\int_{0}^{1}(1-t)\left\{f^{\prime\prime}\left(x+t\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)\right)-f^{\prime\prime}(x)\right\}\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\,\mathrm{d}t\Bigg{]}$
-$\hskip 28.45274pt+\frac{n}{2Z_{\alpha,x}^{(n)}}\left\{\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}-\frac{1}{\alpha n}x(1-x)\right\}f^{\prime\prime}(x)$
-$\hskip 28.45274pt+\frac{1}{2\alpha}\left(\frac{1}{Z_{\alpha,x}^{(n)}}-1\right)x(1-x)f^{\prime\prime}(x)$
-$=:\mathrm{I}_{\alpha}^{(n)}(x)+\mathrm{II}_{\alpha}^{(n)}(x)+\mathrm{III}_{\alpha}^{(n)}(x)+\mathrm{IV}_{\alpha}^{(n)}(x).$ (4.5)
+$\quad+n\mathbb{E}\Bigg[\int_{0}^{1}(1-t)\left\{f^{\prime\prime}\left(x+t\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)\right)-f^{\prime\prime}(x)\right\}\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\,\mathrm{d}t\Bigg]$
+$\quad+\frac{n}{2Z_{\alpha,x}^{(n)}}\left\{\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}-\frac{1}{\alpha n}x(1-x)\right\}f^{\prime\prime}(x)$
+$\quad+\frac{1}{2\alpha}\left(\frac{1}{Z_{\alpha,x}^{(n)}}-1\right)x(1-x)f^{\prime\prime}(x)$
+$=:\mathrm{I}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptV}_{\alpha}^{(n)}(x).$ (4.5)
 Let $\varepsilon>0$. By assumption, there exists some $r>0$ such that
@@
 $\left|\mathrm{I}_{\alpha}^{(n)}(x)\right|$ $\leq n\max_{x\in[0,r]\cup[1-r,1]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right||f^{\prime}(x)|$
-$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\max_{x\in[r,1-r]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|$
+$\quad+n\|f^{\prime}\|_{\infty}\max_{x\in[r,1-r]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|$
 $\leq n\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|\varepsilon$
-$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)O(e^{-\delta n})$
+$\quad+n\|f^{\prime}\|_{\infty}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)O(e^{-\delta n})$
 $=\varepsilon O(1)+nO(e^{-\delta n})=\varepsilon O(1)+o(1)$ (4.6)
@@
 Hence, Theorem 2.9 and Corollary 2.12 allow us to obtain that
-$\left|\mathrm{II}_{\alpha}^{(n)}(x)\right|$
+$\left|\mathrm{I\hskip-1.00006ptI}_{\alpha}^{(n)}(x)\right|$
 $\leq\frac{n\varepsilon}{2}\max_{x\in[0,1]}\mathbb{E}\left[\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\right]+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\max_{x\in[0,1]}\mathbb{E}\left[\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{4}\right]$
 $\leq\frac{n\varepsilon}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|$
-$\hskip 14.22636pt+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{4}\right|$
+$\quad+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{4}\right|$
 $=n\varepsilon O\left(\frac{1}{n}\right)+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}O\left(\frac{1}{n^{2}}\right)$
@@
 as $n\to\infty$. Furthermore, by applying Theorems 2.9 and 2.11, we have
-$\left|\mathrm{III}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{n}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\|f^{\prime\prime}\|_{\infty}O\left(\frac{1}{n^{2}}\right)=O\left(\frac{1}{n}\right)$ (4.8)
-as $n\to\infty$. On the other hand, for $\varepsilon>0$, there exists some sufficiently small $u=u(\varepsilon)>0$ satisfying $x(1-x)<\varepsilon$ for $x\in[0,u]\cup[1-u,1]$. Then, we have
-$\left|\mathrm{IV}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\max_{x\in[0,u]\cup[1-u,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
-$\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\sup_{\alpha\in K}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
+$\left|\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{n}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\|f^{\prime\prime}\|_{\infty}O\left(\frac{1}{n^{2}}\right)=O\left(\frac{1}{n}\right)$ (4.8)
+as $n\to\infty$. On the other hand, for $\varepsilon>0$, $x(1-x)<\varepsilon$ for $x\in[0,\varepsilon]\cup[1-\varepsilon,1]$. Then, we have
+$\left|\mathrm{I\hskip-1.00006ptV}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\max_{x\in[0,\varepsilon]\cup[1-\varepsilon,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\max_{x\in[\varepsilon,1-\varepsilon]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
+$\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\sup_{\alpha\in K}\max_{x\in[\varepsilon,1-\varepsilon]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
 $=\varepsilon O(1)+o(1)$ (4.9)
-as $n\to\infty$, by virtue of (2.13) and Theorem 2.9. Therefore, by combining (4.5) with (4.3), (4.7), (4.8), and (4.9), we obtain
+as $n\to\infty$, by virtue of Proposition 2.6 (1) and Theorem 2.9. Therefore, by combining (4.5) with (4.3), (4.7), (4.8), and (4.9), we obtain
 $\sup_{\alpha\in K}\|n(B_{\alpha,n}-I)f-\mathcal{L}_{\alpha}f\|_{\infty}=\varepsilon O(1)+o(1)$
@@
 The function $g_{\delta}$ (resp.
-$\widehat{g}_{\delta}$) is non-negative and continuous on $[0,\varepsilon]$ (resp.
-$[1-\varepsilon,1]$) with $g_{\delta}(\varepsilon)=0$ (resp.
+$\widehat{g}_{\delta}$) is non-negative and continuous on $0,\varepsilon]$ (resp.
+$1-\varepsilon,1]$) with $g_{\delta}(\varepsilon)=0$ (resp.
 $\widehat{g}_{\delta}(1-\varepsilon)=0$).
@@
 Then, both $G$ and $\widehat{G}$ are continuous on $(0,\varepsilon]$ and satisfy that
-$G(0+)$ =+=$\infty$,= + ∞ , $\qquad G(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(0)|,$
-$\widehat{G}(0+)$ =+=$\infty$,= + ∞ , $\qquad\widehat{G}(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(1)|$
+$G(0+)$ $=+\infty,$ $\qquad G(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(0)|,$
+$\widehat{G}(0+)$ $=+\infty,$ $\qquad\widehat{G}(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(1)|$
 due to (4.11).
-Hence, the mean value theorem implies the existences of $\delta_{0},\delta_{1}\in(0,\varepsilon)$ such that $G(\delta_{0})=|f^{\prime}(0)|$ and $\widehat{G}(\delta_{1})=|f^{\prime}(1)|$.
+Hence, the mean value theorem implies the existence of $\delta_{0},\delta_{1}\in(0,\varepsilon)$ such that $G(\delta_{0})=|f^{\prime}(0)|$ and $\widehat{G}(\delta_{1})=|f^{\prime}(1)|$.
 We now define a family of continuous functions $(f_{\varepsilon})_{\varepsilon>0}\subset C([0,1])$ by
-$f_{\varepsilon}(x):=\begin{cases}f(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}\int_{y}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in[0,\varepsilon),\\ f(x)&\text{if }x\in[\varepsilon,1-\varepsilon],\\ f(x)+\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\int_{1-\varepsilon}^{y}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
+[…]
 Then, we have
-$f_{\varepsilon}^{\prime}(x):=\begin{cases}f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s&\text{if }x\in[0,\varepsilon),\\ f^{\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
+[…]
 In particular, we see that
@@
 We also have
-$f_{\varepsilon}^{\prime\prime}(x):=\begin{cases}f^{\prime\prime}(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle g_{\delta_{0}}(x)&\text{if }x\in[0,\varepsilon),\\ f^{\prime\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\widehat{g}_{\delta_{1}}(x)&\text{if }x\in(1-\varepsilon,1]\\ \end{cases}$
+[…]
 and
@@
 Proof of Theorem 4.4.
-It follows from Lemma 4.6 and [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$.
+It follows from Lemma 4.6 and [7, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$.
 Then, by combining this fact with Lemma 4.5, we can apply Trotter’s approximation theorem (cf.
-[Tro58]) to conclude (4.3).
+[29, Theorem 5.2]) to conclude (4.3).
 ∎
@@
 Moreover, we note that the diffusion coefficient $(2\alpha)^{-1}x(1-x)$ of $\mathcal{L}_{\alpha}$ is continuous and bounded on $[0,1]$.
-Then, we can apply [Kal21, Theorem 32.11] to conclude that
+Then, we can apply [14, Theorem 32.11] to conclude that
 $\widetilde{T}_{\alpha}(t)f(x):=\mathbb{E}[f(X_{\alpha,x}(t))],\qquad f\in C([0,1]),\,x\in[0,1],\,t\geq 0$
@@
 Because a Feller semigroup is uniquely determined by its generator (cf.
-[Kal21, Lemma 17.5]), we have $T_{\alpha}(t)=\widetilde{T}_{\alpha}(t)$ for all $t\geq 0$.
+[14, Lemma 17.5]), we have $T_{\alpha}(t)=\widetilde{T}_{\alpha}(t)$ for all $t\geq 0$.
 ∎
-4.4. The fractional Bernstein operator with $\alpha=\alpha(x)$ being a function of $x$
+4.4 The fractional Bernstein operator with $\alpha=\alpha(x)$ being a function of $x$
 In this section, we consider the fractional Bernstein operator whose parameter $\alpha>0$ is replaced with a function $\alpha=\alpha(x)\colon[0,1]\to(0,\infty)$.
@@
 Let $n\in\mathbb{N}$ and $x\in[0,1]$. Then, the $\alpha(x)$-fractional binomial distribution
-$\mu_{\alpha(x),x}^{(n)}(\{j\})=\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)},\qquad x\in[0,1],\,j=0,1,2,\dots,n$
+$\mu_{\alpha(x),x}^{(n)}(\mathrm{d}z)=\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)}\delta_{j}(\mathrm{d}z)$
 with (4.15) turns out to have almost the same properties as those of $\mu_{\alpha,x}^{(n)}$ discussed in Section 2.
@@
 We now assume that $\alpha(x)$ is continuous in $x$. Let us consider the $\alpha(x)$-fractional Bernstein operator $B_{\alpha(\cdot),n}$ given by
-$B_{\alpha(\cdot),n}f(x)$ $=\sum_{j=0}^{n}\mu_{\alpha(x),x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
+$B_{\alpha(\cdot),n}f(x)$ $=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{\alpha(x),x}^{(n)}(\mathrm{d}z)$
 $=\sum_{j=0}^{n}\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)}f\left(\frac{j}{n}\right)$
@@
 for $f\in C^{2}([0,1])$ and $x\in[0,1]$.
-Then, the operator $(\mathcal{L}_{\alpha(\cdot)},C^{2}([0,1]))$ is also closable on $C([0,1])$, and its closure $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$ generates the Feller semigroup $(T_{\alpha(\cdot)}(t))_{t\geq 0}$ on $C([0,1])$, which is no longer the Wright–Fisher diffusion semigroup in general (see [AC94, Theorems 1.6.11 and 6.2.6]).
-By following the proof of Theorem 4.4, we also obtain that the discrete semigroup defined by the iterates of $B_{\alpha(\cdot),n}$ converges to $(T_{\alpha(\cdot)}(t))_{t\geq 0}$.
+Then, the operator $(\mathcal{L}_{\alpha(\cdot)},C^{2}([0,1]))$ is also closable on $C([0,1])$, and its closure $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$ generates the Feller semigroup $(T_{\alpha(\cdot)}(t))_{t\geq 0}$ on $C([0,1])$, which is no longer the Wright–Fisher diffusion semigroup in general (see [2, Theorems 1.6.11 and 6.2.6]).
+By following the proof of Theorem 4.4, we also prove that the appropriately scaled operators $(B_{\alpha(\cdot),n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, defined by the iterates of $B_{\alpha(\cdot),n}$ converges to $(T_{\alpha(\cdot)}(t))_{t\geq 0}$.
 Theorem 4.8.
@@
 as $\varepsilon\searrow 0$, which implies that $\mathcal{D}_{0}$ is also a core for $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$.
-Therefore, it follows from [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha(\cdot)}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$.
+Therefore, it follows from [7, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha(\cdot)}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$.
 Then, Trotter’s approximation theorem allows us to obtain (4.16).
@@
 Proposition 4.9.
+We assume that $\alpha=\alpha(x)\colon[0,1]\to(0,\infty)$ is continuous and $\alpha(x)^{-1}=1/\alpha(x)$ is Lipschitz continuous on $[0,1]$.
 For $x\in[0,1]$, let $(X_{\alpha(\cdot),x}(t))_{t\geq 0}$ be the one-dimensional diffusion process that solves the stochastic differential equation
@@
 $T_{\alpha(\cdot)}(t)f(x)=\mathbb{E}\big[f\big(T_{\alpha(\cdot),x}(t)\big)\big],\qquad f\in C([0,1]),\,x\in[0,1],\,t\geq 0.$
-Here, we note that (4.17) has the unique strong solution by applying the Yamada–Watanabe theorem (cf.
-[YW71, Theorem 1]).
-We omit the proof of Proposition 4.9 because it is similar to that of Proposition 4.7.
-Acknowledgements. The first-named author was supported by JSPS KAKENHI Grant Number 19H00643. The second-named author was supported by JSPS KAKENHI Grant Number 23K12986.
+Proof.
+We omit the details of the proof because it is essentially same as that of Proposition 4.7.
+On the other hand, we should note that (4.17) has the unique strong solution.
+Indeed, the existence of the weak solution to (4.17) holds true since the diffusion coefficient
+$\sigma(x):=\sqrt{\alpha(x)^{-1}x(1-x)},\qquad x\in[0,1],$
+of (4.17) is bounded and continuous on $[0,1]$. Moreover, it follows from the Lipschitz continuity of $1/\alpha(x)$ that
+$|\sigma(x)-\sigma(y)|$ $=\left|\sqrt{\alpha(x)^{-1}x(1-x)}-\sqrt{\alpha(y)^{-1}y(1-y)}\right|$
+$\leq\sqrt{\left|\alpha(x)^{-1}x(1-x)-\alpha(y)^{-1}y(1-y)\right|}$
+$\leq\sqrt{\left|\alpha(x)^{-1}-\alpha(y)^{-1}\right|x(1-x)+\alpha(y)^{-1}|x(1-x)-y(1-y)|}$
+$\leq\sqrt{C_{4,3}|x-y|+3\alpha_{L}^{-1}|x-y|}\leq C_{4,4}|x-y|^{1/2},\qquad x,y\in[0,1],$
+which implies the pathwise uniqueness of (4.17). Therefore, the Yamada–Watanabe theorem (cf. [30, Theorem 1]) leads to the existence of the unique strong solution to (4.17). ∎
+Acknowledgements.
+The first-named author was supported by JSPS KAKENHI Grant Numbers 19H00643 and 25K07056 and The Kyoto University Foundation.
+The second-named author was supported by JSPS KAKENHI Grant Numbers 23K12986 and 23K03155.
+Data availability No datasets were generated or analyzed during the current study.
+Declarations
+Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.